### v1

# On the optimality of coin-betting for mean estimation

Abstract
Confidence sequences are sequences of confidence sets that adapt to incoming data while maintaining validity. Recent advances have introduced an algorithmic formulation for constructing some of the tightest confidence sequences for bounded real random variables. These approaches use a coin-betting framework, where a player sequentially bets on differences between potential mean values and observed data. This letter establishes that such coin-betting formulation is optimal among all possible algorithmic frameworks for constructing confidence sequences that build on e-variables and sequential hypothesis testing.
keywords:
mean estimation , confidence sequences , coin-betting , e-variables , sequential hypothesis testing
journal: Statistics & Probability Letters
1 Introduction
Estimating the mean of a random variable from empirical observations is a classical problem in statistics. To account for uncertainty, a widely used approach consists in constructing a confidence set, known to contain the true mean with high probability, rather than relying solely on a point estimate. When the data are observed sequentially, one might want to update this set as new data-points become available. However, such procedure can compromise the validity of the statistical guarantee, if this was designed for a fixed sample size. To address this issue, Darling and Robbins (1967) introduced the concept of confidence sequence, a data-adaptive sequence of confidence sets whose intersection contains the desired mean with high probability.
Orabona and Jun (2023) and Waudby-Smith and Ramdas (2023) have recently suggested algorithmic approaches that yield some of the tightest confidence sequences for the mean of a bounded real random variable. Both papers propose setting up a series of sequential coin-betting games, one per each mean candidate value $\mu$, where a player sequentially bets on the difference between $\mu$ and the upcoming observation. If $\mu$ matches the true mean, the game is fair, making substantial gains unlikely. A confidence sequence is then obtained by excluding those values $\mu$ that allowed the player to accumulate significant wealth.
This coin-betting approach to mean estimation is a particular instance of a broader algorithmic framework for constructing confidence sequences through sequential hypothesis testing, which can be framed in terms of betting games where at each round the player has to select an e-variable (Shafer, 2021; Ramdas et al., 2022, 2023). E-variables, namely non-negative random variables whose expectation is bounded by one under the tested hypothesis (Grünwald et al., 2024), have recently emerged as a powerful and increasingly popular tool for anytime-valid hypothesis testing. By serving as building blocks for constructing non-negative super-martingales, which can be seen as representing the wealth of a player in a betting game, e-variables naturally lend themselves to game-theoretic interpretations (Shafer and Vovk, 2019; Ramdas and Wang, 2024).
The purpose of this letter is to illustrate that, when estimating the mean of a bounded real random variable, no instantiation of an e-variable-based algorithmic framework can yield strictly better results than the coin-betting approach, in terms of tightness of the resulting confidence sequence. In a sense to be clarified later, coin-betting is optimal, as it represents the “simplest” formulation among those that cannot be strictly outperformed by any other such algorithmic approach. The main novelty of this work is the introduction of a notion of optimality at the level of sets of e-variables. This perspective differs from the existing literature, which is primarily focused on the optimality of an individual e-variable (e.g., log-optimality in Koolen and Grünwald 2022; Grünwald et al. 2024; Larsson et al. 2024) or of a single wealth super-martingale (e.g., admissibility in Ramdas et al. 2022). The author hopes that this new viewpoint will inspire further research.
Notation
Endow the interval $[0,1]$ with the standard topology, and let $\mathcal{P}$ be the set of all Borel probability measures on $[0,1]$. For $x$ in $[0,1]$, $\delta_{x}$ denotes the Dirac unit mass on $x$. For $P\in\mathcal{P}$ and a Borel measurable function $f$ on $[0,1]$, $\langle P,f(X)\rangle$ (or more compactly $\langle P,f\rangle$) denotes the expectation of $f$ under $X\sim P$. Sequences are denoted as $(s_{t})_{t\geq T_{0}}$, with $t$ an integer index and $T_{0}$ its smallest value (typically 00 or $1$). For high probability statements, $\mathbb{P}$ expresses probability with respect to all the randomness involved. For instance, if $(X_{t})_{t\geq 1}$ is a sequence of i.i.d. draws from $P\in\mathcal{P}$, we write $\mathbb{P}(X_{t}\geq 1/2\,,\,\forall t\geq 1)$, with obvious meaning.
2 Algorithmic mean estimation
Let $(X_{t})_{t\geq 1}$ be a sequence of independent random variables drawn from an unknown fixed distribution $P^{\star}\in\mathcal{P}$, whose mean $\mu^{\star}\in(0,1)$ has to be estimated. Let $\mathcal{F}=(\mathcal{F}_{t})_{t\geq 0}$ represent the natural filtration generated by this sequence, where $\mathcal{F}_{t}=\sigma(X_{1},\dots,X_{t})$ captures all information available up to time $t$. Fix a confidence level parameter $\delta\in(0,1)$. A confidence sequence $(S_{t})_{t\geq 1}$ is a sequence of random sets such that the sequence of events ($(\{\mu^{\star}\in S_{t}\})_{t\geq 1}$ is adapted with respect to the filtration $\mathcal{F}$ (i.e., for all $t\geq 1$, $\{\mu^{\star}\in S_{t}\}$ is $\mathcal{F}_{t}$-measurable) and satisfies the property
$\mathbb{P}\big{(}\mu^{\star}\in S_{t}\,,\;\forall t\geq 1\big{)}\geq 1-\delta\,.$
Intuitively, this means that $(S_{t})_{t\geq 1}$ provides a set of plausible values for $\mu^{\star}$ at each time step $t$, while ensuring that the true mean remains in these sets indefinitely with high probability.
2.1 Coin-betting confidence sequence
Recently, Orabona and Jun (2023) and Waudby-Smith and Ramdas (2023) obtained some of the tightest known confidence sequences for bounded random variables via an algorithmic approach. The idea behind both papers is to set up a sequential coin-betting game for each candidate value $\mu\in(0,1)$ for the mean, and to exclude from the confidence sets those values $\mu$ that allow the player to earn unlikely high rewards.
Definition 1 (Coin-betting game)
Fix $\mu\in(0,1)$ and let $I_{\mu}=[(\mu-1)^{-1},\mu^{-1}]$. Consider the following sequential procedure. At each round $t\geq 1$, a player
  • 1.
    picks [^1] We implicitly assume a measurable selection of $\beta_{t}$, namely the mapping $(x_{1},\dots x_{t-1})\mapsto\beta_{t}$ is Borel measurable. $\beta_{t}\in I_{\mu}$, based solely on the past observations $x_{1},\dots,x_{t-1}$;
  • 2.
    observes a new data-point $x_{t}\in[0,1]$;
  • 3.
    gets a reward $\log\big{(}1+\beta_{t}(x_{t}-\mu)\big{)}$.
Now, for each $\mu\in(0,1)$, consider a coin-betting game where at each round the player observes a data-point from the sequence $(X_{t})_{t\geq 1}$ drawn from $P^{\star}$. Let $R_{n}(\mu)=\sum_{t=1}^{n}\log(1+\beta_{t}(X_{t}-\mu))$ denote the cumulative reward at round $n$. Then, via standard martingale concentration arguments (which will become clear later) one can show that the sets $S_{n}=\big{\{}\mu\in(0,1)\,:\,R_{n}(\mu)\leq\log\tfrac{1}{\delta}\big{\}}$ define a valid confidence sequence for $\mu^{\star}$.
2.2 Sequential testing game
The coin-betting approach is a particular instance of constructing confidence sequences via a sequential hypothesis testing procedure based on e-variables. We call hypothesis an arbitrary subset $\mathcal{H}$ of $\mathcal{P}$. An e-variable (with respect to $\mathcal{H}$) is any non-negative Borel measurable function $E:[0,1]\to[0,+\infty)$, such that
$\langle P,E\rangle\leq 1\,,\qquad\forall P\in\mathcal{H}\,.$
We denote as $\mathcal{E}_{\mathcal{H}}$ the set of all the e-variables with respect to $\mathcal{H}$, and we call e-class any subset of $\mathcal{E}_{\mathcal{H}}$. We remark that $\mathcal{E}_{\mathcal{H}}$ is never empty, since the constant function one is always an e-variable, independently of $\mathcal{H}$.
Definition 2 (Testing game)
Fix a hypothesis $\mathcal{H}\subseteq\mathcal{P}$ and an e-class $\mathcal{E}\subseteq\mathcal{E}_{\mathcal{H}}$. We call $\mathcal{E}$-restricted testing game (on $\mathcal{H}$) the following sequential procedure. At each round $t\geq 1$, a player
  • 1.
    picks [^2] Again, we assume that $E_{t}$ is picked in a measurable fashion, namely the mapping $(x_{1},\dots x_{t})\mapsto E_{t}(x_{t})$ is Borel measurable. an e-variable $E_{t}\in\mathcal{E}$ based solely on the past observations $x_{1},\dots,x_{t-1}$;
  • 2.
    observes a new data-point $x_{t}\in[0,1]$;
  • 3.
    gets a reward $\log E_{t}(x_{t})$.
If $\mathcal{E}=\mathcal{E}_{\mathcal{H}}$, we speak of unrestricted testing game.
The above game is at the core of e-variable testing (see, e.g., Ramdas and Wang 2024 for a recent overview), where one designs a test that rejects the hypothesis $\mathcal{H}$ whenever the total reward earned by the player gets excessively high. This procedure is justified by the fact that, if the data-points observed during the game were independently drawn from $P\in\mathcal{H}$, then the cumulative reward would be unlikely to grow very large.
Lemma 1
Let $\mathcal{H}\subseteq\mathcal{P}$ and consider a sequence $(X_{t}^{\prime})_{t\geq 1}$ of independent draws from some $P\in\mathcal{H}$. Fix $\delta\in(0,1)$ and $\mathcal{E}\subseteq\mathcal{E}_{\mathcal{H}}$. Consider an $\mathcal{E}$-restricted testing game, where the observations are the sequence $(X_{t}^{\prime})_{t\geq 1}$. Let $R_{n}=\sum_{t=1}^{n}\log E_{t}(X_{t}^{\prime})$ represent the player’s cumulative reward at round $n$. We have
$\mathbb{P}\big{(}R_{n}\leq\log\tfrac{1}{\delta}\,,\;\forall n\geq 1\big{)}\geq 1-\delta\,.$
Proof 1
The result follows directly from Ville’s inequality, since $M_{n}=\prod_{t=1}^{n}E_{t}(X_{t}^{\prime})$ defines a non-negative super-martingale with respect to the natural filtration of the process $(X_{t}^{\prime})_{t\geq 1}$, with $M_{0}\equiv 1$$\qed$
2.3 Mean estimation via e-variables
We can leverage the testing game of Definition 2 to obtain a confidence sequence for the mean $\mu^{\star}\in(0,1)$ of $P^{\star}$. For $\mu\in(0,1)$, define the hypothesis
$\mathcal{H}_{\mu}=\{P\in\mathcal{P}\,:\,\langle P,X\rangle=\mu\}\,,$ (1)
which contains all probability measures on $[0,1]$ with mean $\mu$. For each $\mu$, we fix an e-class $\mathcal{E}_{\mu}\subseteq\mathcal{E}_{\mathcal{H}_{\mu}}$ and we consider an $\mathcal{E}_{\mu}$-restricted testing game on $\mathcal{H}_{\mu}$, where the player observes $(X_{t})_{t\geq 1}$, the sequence of draws from $P^{\star}$. For each one of these games, denote as $R_{n}(\mu)$ the player’s cumulative reward at round $n$.
Lemma 2
The sequence $(S_{n})_{n\geq 1}$, defined as
$S_{n}=\big{\{}\mu\in(0,1)\;:\;R_{n}(\mu)\leq\log\tfrac{1}{\delta}\big{\}}\,,$
is a confidence sequence for the mean $\mu^{\star}$ of the data-generating probability measure $P^{\star}$.
Proof 2
For each $n\geq 1$, we have that $\{\mu^{\star}\in S_{n}\}=\{R_{n}(\mu^{\star})\leq\log\frac{1}{\delta}\}$, which is a $\mathcal{F}_{n}$-measurable event. Moreover,
$\mathbb{P}\big{(}\mu^{\star}\in S_{n}\,,\;\forall n\geq 1\big{)}=\mathbb{P}\big{(}R_{n}(\mu^{\star})\leq\log\tfrac{1}{\delta}\,,\;\forall n\geq 1\big{)}\geq 1-\delta$
by Lemma 1, as the observations are drawn from $P^{\star}\in\mathcal{H}_{\mu^{\star}}$. $\qed$
We remark that the approach to derive confidence sequences via the coin-betting game, proposed by Orabona and Jun (2023) and Waudby-Smith and Ramdas (2023), is a specific instance of the framework that we have just described. Indeed, letting $E_{t}:x\mapsto 1+\beta_{t}(x-\mu)$, it is straightforward to verify that $E_{t}$ is non-negative on $[0,1]$, due to the restriction $\beta_{t}\in I_{\mu}=[(\mu-1)^{-1},\mu^{-1}]$ in Definition 1. Moreover, for any $P\in\mathcal{H}_{\mu}$, we have $\langle P,E_{t}\rangle=1$, which implies that $E_{t}$ is an e-variable for $\mathcal{H}_{\mu}$. Finally, the reward in the coin-betting game is precisely equal to $\log E_{t}(x_{t})$. Hence, for the hypothesis $\mathcal{H}_{\mu}$, the coin-betting game matches exactly the testing game of Definition 2, restricted to the coin-betting e-class
$\mathcal{E}^{\mathrm{cb}}_{\mu}=\big{\{}E_{\beta}:x\mapsto 1+\beta(x-\mu)\,,\;\;\beta\in I_{\mu}\big{\}}\,.$ (2)
Ideally, one aims to construct tight confidence sets. However, achieving this depends on the strategy employed in the testing games. For example, stubbornly playing $E_{t}\equiv 1$ at all rounds in every game would produce valid confidence sets ($S_{n}=(0,1)$, for every $n$) of no practical interest. Indeed, a player’s strategy is most effective when it can rapidly increase the cumulative reward, whenever possible, as this allows to fast exclude redundant points from the confidence sets. In short, the highest the rewards, the tightest the confidence sequence. With this in mind, it is clear that carelessly restricting the player’s choice to a subset $\mathcal{E}_{\mu}$ of $\mathcal{E}_{\mathcal{H}_{\mu}}$ could be highly detrimental, as it might force the player to adopt poor strategies (e.g., consider the trivial case $\mathcal{E}_{\mu}=\{1\}$). This point naturally raises the question: “Does restricting the player to the coin-betting e-class (2) loosen the confidence sequence?”. Interestingly, the answer turns out to be negative, as we will demonstrate throughout the remainder of this letter.
3 Majorising e-classes and optimal e-class
We introduce a partial ordering on the set of real functions on $[0,1]$. Given two functions $f,f^{\prime}:[0,1]\to\mathbb{R}$, we say that $f$ majorises $f^{\prime}$, and write $f\succeq f^{\prime}$, if $f(x)\geq f^{\prime}(x)$ for all $x\in[0,1]$. If $f\succeq f^{\prime}$ and there exists a point $x\in[0,1]$ such that $f(x)>f^{\prime}(x)$, we say that $f$ is a strict majoriser of $f^{\prime}$, and we write $f\succ f^{\prime}$.
Definition 3
$E\in\mathcal{E}_{\mathcal{H}}$ is called a maximal e-variable when there is no $E^{\prime}\in\mathcal{E}_{\mathcal{H}}$ such that $E^{\prime}\succ E$.
Next, we introduce a way to compare different e-classes.
Definition 4
Given two e-classes $\mathcal{E}$ and $\mathcal{E}^{\prime}$, we say that $\mathcal{E}$ majorises $\mathcal{E}^{\prime}$ if, for any $E^{\prime}\in\mathcal{E}$, there is an e-variable $E\in\mathcal{E}$ such that $E\succeq E^{\prime}$. We call majorising e-class any e-class that majorises $\mathcal{E}_{\mathcal{H}}$.
Lemma 3
Every majorising e-class contains all the maximal e-variables.
Proof 3
Let $E\in\mathcal{E}_{\mathcal{H}}$ be maximal and $\mathcal{E}$ a majorising e-class. There must be $E^{\prime}\in\mathcal{E}$ such that $E^{\prime}\succeq E$, but since $E$ is maximal it has to be that $E=E^{\prime}$. Hence, $E\in\mathcal{E}$.
The significance of the notion of majorising e-class for our problem is straightforward: if $\mathcal{E}$ majorises $\mathcal{E}^{\prime}$, then any strategy in an $\mathcal{E}^{\prime}$-restricted game can be matched or outperformed (in terms of rewards) by a corresponding strategy in the $\mathcal{E}$-restricted game, regardless of the sequence of observations. This allows us to compare how the restriction to different e-classes affects the testing game of Definition 2.
Definition 5
If a majorising e-class is contained in every other majorising e-class, it is called optimal.
For any $\mathcal{H}$, a majorising e-class always exists, as $\mathcal{E}_{\mathcal{H}}$ itself is a majorising e-class. However, an optimal e-class may not exist. Next, we state a sufficient and necessary condition for the existence of optimal e-classes.
Lemma 4
An optimal e-class exists if, and only if, the set of all maximal e-variables is a majorising e-class. If an optimal e-class exists, it is unique, it corresponds to the set of all maximal e-variables, and it is the only majorising e-class whose elements are all maximal.
Proof 4
Denote as $\hat{\mathcal{E}}$ the set of all the maximal e-variables. Now, assume that there exists an optimal e-class $\mathcal{E}$. First, we show that all its elements must be maximal. For $E\in\mathcal{E}$, consider any element $E^{\prime}\in\mathcal{E}_{\mathcal{H}}$ such that $E^{\prime}\succeq E$. We can construct an e-class $\mathcal{E}^{\prime}$ replacing $E$ with $E^{\prime}$ in $\mathcal{E}$, namely =($\mathcal{E}^{\prime}=(\mathcal{E}\setminus\{E\})\cup\{E^{\prime}\}$. Since $E^{\prime}\succeq E$, it is clear that $\mathcal{E}^{\prime}$ majorises $\mathcal{E}$. It follows that $\mathcal{E}^{\prime}$ is a majorising e-class, as $\mathcal{E}$ is. Since $\mathcal{E}$ is optimal, $\mathcal{E}\subseteq\mathcal{E}^{\prime}$, which implies $E=E^{\prime}$. In particular, $E$ does not have any strict majoriser, and so it is maximal. Hence, $\mathcal{E}\subseteq\hat{\mathcal{E}}$. As $\mathcal{E}\supseteq\hat{\mathcal{E}}$ by Lemma 3, we conclude that $\mathcal{E}=\hat{\mathcal{E}}$, and so $\hat{\mathcal{E}}$ is a majorising e-class.
Conversely, assume that $\hat{\mathcal{E}}$ is a majorising e-class. Let $\mathcal{E}$ be any other majorising e-class. By Lemma 3, $\hat{\mathcal{E}}\subseteq\mathcal{E}$. So, $\hat{\mathcal{E}}$ is contained in all the majorising e-classes, and hence it is optimal.
Now, the final statements are a trivial consequence of what shown above and Lemma 3$\qed$
Let us emphasise once more that, from our discussion thus far, it is clear that restricting the testing game of Definition 2 to a majorising e-class does not hinder the performance of the player, as for any unrestricted strategy $(E_{t})_{t\geq 1}$ they can always pick a restricted strategy $(E_{t}^{\prime})_{t\geq 1}$, whose cumulative rewards inevitably match or outperform those of $(E_{t})_{t\geq 1}$, regardless of the sequence of observations. From a practical perspective, identifying the optimal e-class, when it exists, greatly simplifies the design of an effective strategy by narrowing the player’s choice to the best possible e-variables. Specifically, if the optimal e-class exists and a player chooses an e-variable $E_{t}$ outside of it, they could always have picked an alternative $E_{t}^{\prime}\succ E_{t}$, within the optimal e-class, whose reward is never worse than that of $E_{t}$ and is strictly higher for at least one possible value that $x_{t}$ might take. Conversely, when a player selects an e-variable from the optimal e-class, no other choice can be guaranteed to be better before observing $x_{t}$, since the player’s pick is a maximal e-variable. As a straightforward consequence, the optimal approach to obtain confidence sequences consists in restricting the testing game to the optimal e-class, for each mean candidate value $\mu$. We will show next that this coincides precisely with the coin-betting formulation.
4 Optimality of the coin-betting e-class
For any $\mu\in(0,1)$, define the hypothesis $\mathcal{H}_{\mu}$ as in (1). We now show that the optimal e-class for $\mathcal{H}_{\mu}$ exists and coincides with the coin-betting e-class $\mathcal{E}^{\mathrm{cb}}_{\mu}$, defined in (2).
Lemma 5
Fix $\mu\in(0,1)$ and consider the function $F_{\mu}:[0,1]\to[1,+\infty)$ defined as
$F_{\mu}\;:\;x\mapsto\begin{cases}1+\frac{1}{\mu}(x-\mu)&\text{if $x\geq\mu$;}\\ 1+\frac{1}{\mu-1}(x-\mu)&\text{if $x<\mu$.}\end{cases}$
For any $x\in[0,1]$ there is $P_{x}\in\mathcal{H}_{\mu}$ such that $P_{x}(\{x\})=1/F_{\mu}(x)>0$. Moreover, $F_{\mu}\succeq E$ for every $E\in\mathcal{E}_{\mathcal{H}_{\mu}}$.
Proof 5
For $x\geq\mu$, let $P_{x}=\frac{\mu}{x}\delta_{x}+(1-\frac{\mu}{x})\delta_{0}$. Then, $P_{x}\in\mathcal{H}_{\mu}$ and $P_{x}(\{x\})=\mu/x=1/F_{\mu}(x)$. With a similar reasoning, if $x<\mu$ we can find a measure $P_{x}$ in $\mathcal{H}_{\mu}$, supported on $\{x,1\}$, with mass $1/F_{\mu}(x)$ on $x$.
Now, to check that $F$ majorises all the e-variables, fix $E\in\mathcal{E}_{\mathcal{H}_{\mu}}$ and $x\in[0,1]$. Consider a $P_{x}\in\mathcal{H}_{\mu}$ with mass $1/F_{\mu}(x)$ on $x$. Then, $1\geq\langle P_{x},E\rangle\geq P_{x}(x)E(x)=E(x)/F_{\mu}(x)$, and we conclude. $\qed$
Theorem 1
For any $\mu\in(0,1)$, the coin-betting e-class $\mathcal{E}^{\mathrm{cb}}_{\mu}$ is the optimal e-class for $\mathcal{H}_{\mu}$.
Proof 6
First, let us show that $\mathcal{E}^{\mathrm{cb}}_{\mu}$ is a majorising e-class. Fix an arbitrary $E\in\mathcal{E}_{\mathcal{H}_{\mu}}$. Define the sets
$B_{0}=\left\{\beta\in I_{\mu}\;:\;\inf_{x\in[0,\mu]}\big{(}E_{\beta}(x)-E(x)\big{)}\geq 0\right\}\qquad\text{and}\qquad B_{1}=\left\{\beta\in I_{\mu}\;:\;\inf_{x\in[\mu,1]}\big{(}E_{\beta}(x)-E(x)\big{)}\geq 0\right\}\,,$
where we recall that $I_{\mu}=[(1-\mu)^{-1},\mu^{-1}]$ and $E_{\beta}:x\mapsto 1+\beta(x-\mu)$. Both sets are closed and convex (as they are intersections of closed and convex sets). By Lemma 5, $E_{(\mu-1)^{-1}}\in B_{0}$ and $E_{\mu^{-1}}\in B_{1}$, so $B_{0}=[(\mu-1)^{-1},\beta_{0}]$ and $B_{1}=[\beta_{1},\mu]$, for some $\beta_{0}$ and $\beta_{1}$ in $I_{\mu}$. We will now show that $B_{0}\cap B_{1}\neq\varnothing$, or equivalently that $\beta_{0}\geq\beta_{1}$. Assume that this was not the case and $\beta_{0}<\beta_{1}$. Let $\beta^{\star}\in(\beta_{0},\beta_{1})$. Then, $\beta^{\star}\notin B_{0}$ and $\beta^{\star}\notin B_{1}$. In particular, there are $x_{0}\in[0,\mu]$ and $x_{1}\in[\mu,1]$ such that $E(x_{0})>E_{\beta^{\star}}(x_{0})$ and $E(x_{1})>E_{\beta^{\star}}(x_{1})$. Notice that since $\delta_{\mu}\in\mathcal{H}_{\mu}$, it must be that $E(\mu)\leq 1$, and so $x_{0}<\mu$ and $x_{1}>\mu$. Hence, $\mu\in(x_{0},x_{1})$ and so there is $\hat{P}\in\mathcal{H}_{\mu}$ with support $\{x_{0},x_{1}\}$. Notice that $\langle\hat{P},E_{\beta^{\star}}\rangle=1+\beta^{\star}(\langle\hat{P},X\rangle-\mu)=1$. But $E$ is strictly larger than $E_{\beta^{\star}}$ on $\mathrm{Supp}(\hat{P})$, and so $\langle\hat{P},E\rangle>1$, which is a contradiction since $E$ is an e-variable. We conclude that $\beta_{0}\geq\beta_{1}$ and so there exists $\hat{\beta}\in B_{0}\cap B_{1}$. By construction, $E_{\hat{\beta}}\in\mathcal{E}^{\mathrm{cb}}_{\mu}$ and $E_{\hat{\beta}}\succeq E$. Because the choice of $E$ was arbitrary, $\mathcal{E}^{\mathrm{cb}}_{\mu}$ is a majorising e-class.
Now that we have established that $\mathcal{E}^{\mathrm{cb}}_{\mu}$ is a majorising e-class, by Lemma 4 we only need to show that all its elements are maximal to conclude that it is optimal. Fix $E\in\mathcal{E}^{\mathrm{cb}}_{\mu}$, and consider an e-variable $\hat{E}\in\mathcal{E}_{\mathcal{H}_{\mu}}$ such that $\hat{E}\succeq E$. Fix any $x\in[0,1]$. By Lemma 5, there is $P\in\mathcal{H}_{\mu}$ such that $P(\{x\})>0$. Since $E\in\mathcal{E}^{\mathrm{cb}}_{\mu}$, we have $\langle P,E\rangle=1$, and so $0\leq P(\{x\})(\hat{E}(x)-E(x))\leq\langle P,\hat{E}-E\rangle=\langle P,\hat{E}\rangle-1\leq 0$. Since $P(\{x\})>0$, we get $\hat{E}(x)=E(x)$ and, $x$ being arbitrary, $\hat{E}=E$. Hence $E$ is maximal, as it has no strict majoriser. $\qed$
5 Perspectives
Theorem 1 gives a rigorous sense to the claim that the coin-betting formulation is optimal among the e-variable-based approaches to build confidence sequences for $\mu^{\star}$. To formalise this assertion, we introduced the notions of majorising and optimal e-classes, which, in the author’s view, are of independent interest in the context of sequential testing, beyond the scope of this letter. The main novelty of these concepts lies in defining “optimality” in terms of e-classes, rather than individual e-variables. This perspective contrasts with the notion of log-optimality of a single e-variable with respect to an alternative hypothesis, widely discussed in the literature (e.g., Koolen and Grünwald, 2022; Grünwald et al., 2024; Larsson et al., 2024).
Rather than log-optimality, what probably best matches our way of “ranking” e-variables is the idea of admissibility introduced by Ramdas et al. (2022). Roughly speaking, let us say that a wealth-process (namely, a non-negative adapted process starting at one that is a super-martingale under every $P\in\mathcal{H}$) is admissible if it is not dominated by any other wealth-process. This concept is tightly related to our notion of maximal e-variables, as it is not hard to see that any wealth-process $(M_{n})_{n\geq 1}$ on $\mathcal{H}_{\mu}$ is admissible if, and only if, it can be associated with a sequence $(E_{t})_{t\geq 1}$ of maximal e-variables such that $M_{n}=\prod_{t=1}^{n}E_{t}(X_{t}^{\prime})$, with $(X_{t}^{\prime})_{t\geq 1}$ a sequence generated by some $P\in\mathcal{H}_{\mu}$. Thus, in our setting, Theorem 1 provides a characterisation of the set of all admissible wealth-processes for $\mathcal{H}_{\mu}$. Let us emphasise that the tools from Ramdas et al. (2022) alone are insufficient to achieve this, as the necessary conditions for admissibility provided by their work rely on the existence of a “reference measure” for the hypothesis, [^3] A hypothesis $\mathcal{H}$ admits a “reference measure” if there exists a Borel measure $Q$ such that $P\ll Q$ for all $P\in\mathcal{H}$. a property that $\mathcal{H}_{\mu}$ does not satisfy. More precisely, their results are enough to prove that every e-variable in the coin-betting e-class is maximal, but not that $\mathcal{E}^{\mathrm{cb}}_{\mu}$ is a majorising e-class. This point is subtle but essential: establishing the optimality of coin-betting requires studying the properties of a set of e-variables (or equivalently, wealth-processes) rather than focusing on single e-variables or wealth-processes individually. Crucially, it is precisely the fact that $\mathcal{E}^{\mathrm{cb}}_{\mu}$ is a majorising e-class that ensures that nothing is lost by relying on the coin-betting approach.
To the author’s knowledge, this is the first work to rigorously examine the optimality of the coin-betting formulation, as previous discussion on the topic has relied solely on heuristic arguments. For instance, Waudby-Smith and Ramdas (2023) justify restricting to the coin-betting e-class by noticing that $\mathcal{E}^{\mathrm{cb}}_{\mu}$ is precisely the set ¯$\bar{\mathcal{E}}_{\mathcal{H}_{\mu}}$ of e-variables whose expectation equals one under every $P\in\mathcal{H}_{\mu}$. However, this property does not generally imply the optimality of the e-class, as it merely indicates that all its elements are maximal. As a simple counterexample, consider $\mathcal{H}^{\prime}_{\mu}=\{P\in\mathcal{P}\,:\,\langle P,X\rangle\leq\mu\}$. Then, ¯$\bar{\mathcal{E}}_{\mathcal{H}^{\prime}_{\mu}}=\{1\}$, while the optimal e-class exists and consists of all functions in the form $E_{\beta}:x\mapsto 1+\beta(x-\mu)$, with $\beta\in[0,\mu^{-1}]$.
To simplify the exposition, we have focused on random variables valued in $[0,1]$. However, all the results presented can be extended to any bounded closed interval. More broadly, the author anticipates that similar conclusions hold for more general settings, such as estimating the mean of bounded vector-valued random variables. Addressing this case requires more advanced analytic tools and lies beyond the scope of this letter.
As a final remark, this paper does not aim to delve into the design of effective coin-betting strategies for obtaining tight confidence sequences. For this, we refer the interested reader to the thorough analysis and discussion by Orabona and Jun (2023) and Waudby-Smith and Ramdas (2023).
Acknowledgements
The author would like to thank Gergely Neu, Peter Daniel Grünwald, Nishant Ajay Mehta, Hamish Edward Flynn, and Claudia Maria Chanu, for the insightful discussions that inspired this work. The author acknowledges the use of GPT-4o during the redaction of this letter, in order to polish the presentation. All AI-generated text was subsequently reviewed and edited by the author, who takes full responsibility for the content of this manuscript. This project was funded by the European Research Council (ERC), under the European Union’s Horizon 2020 research and innovation programme (grant agreement 950180).

### v2

@@
 Confidence sequences are sequences of confidence sets that adapt to incoming data while maintaining validity.
-Recent advances have introduced an algorithmic formulation for constructing some of the tightest confidence sequences for bounded real random variables.
+Recent advances have introduced an algorithmic formulation for constructing some of the tightest confidence sequences for the mean of bounded real random variables.
 These approaches use a coin-betting framework, where a player sequentially bets on differences between potential mean values and observed data.
-This letter establishes that such coin-betting formulation is optimal among all possible algorithmic frameworks for constructing confidence sequences that build on e-variables and sequential hypothesis testing.
-keywords:
-mean estimation , confidence sequences , coin-betting , e-variables , sequential hypothesis testing
-journal: Statistics & Probability Letters
+This work discusses the optimality of such coin-betting formulation among algorithmic frameworks building on e-variables methods to test and estimate the mean of bounded random variables.
 1 Introduction
@@
 A confidence sequence is then obtained by excluding those values $\mu$ that allowed the player to accumulate significant wealth.
-This coin-betting approach to mean estimation is a particular instance of a broader algorithmic framework for constructing confidence sequences through sequential hypothesis testing, which can be framed in terms of betting games where at each round the player has to select an e-variable (Shafer, 2021; Ramdas et al., 2022, 2023).
-E-variables, namely non-negative random variables whose expectation is bounded by one under the tested hypothesis (Grünwald et al., 2024), have recently emerged as a powerful and increasingly popular tool for anytime-valid hypothesis testing.
+This coin-betting approach to mean estimation is a particular instance of a broader algorithmic framework for constructing confidence sequences through sequential hypothesis testing, which can be framed in terms of betting games where at each round the player has to select an e-variable (Shafer, 2021; Ramdas et al., 2022a, 2023).
+E-variables, non-negative random variables whose expectation is bounded by one under the tested hypothesis (Grünwald et al., 2024), have recently emerged as a powerful and increasingly popular tool for anytime-valid hypothesis testing.
 By serving as building blocks for constructing non-negative super-martingales, which can be seen as representing the wealth of a player in a betting game, e-variables naturally lend themselves to game-theoretic interpretations (Shafer and Vovk, 2019; Ramdas and Wang, 2024).
-The purpose of this letter is to illustrate that, when estimating the mean of a bounded real random variable, no instantiation of an e-variable-based algorithmic framework can yield strictly better results than the coin-betting approach, in terms of tightness of the resulting confidence sequence.
-In a sense to be clarified later, coin-betting is optimal, as it represents the “simplest” formulation among those that cannot be strictly outperformed by any other such algorithmic approach.
-The main novelty of this work is the introduction of a notion of optimality at the level of sets of e-variables.
-This perspective differs from the existing literature, which is primarily focused on the optimality of an individual e-variable (e.g., log-optimality in Koolen and Grünwald 2022; Grünwald et al.
+The first part of this work focuses on round-wise testing-by-betting, a scenario where the player iteratively picks a single-round e-variable, whose choice can depend on the past observations.
+Our main goal is to illustrate that, when estimating the mean of a bounded real random variable, no instantiation of such round-wise testing-by-betting game can yield strictly better results than the coin-betting approach, in terms of tightness of the resulting confidence sequence.
+In a sense to be clarified later, coin-betting is optimal, as it represents the “simplest” formulation among those that cannot be strictly outperformed by any other such betting approach.
+One main novelty of this work is the introduction of a notion of optimality at the level of sets of e-variables.
+This perspective differs from much of the existing literature, which is primarily focused on the optimality of an individual e-variable or wealth process (e.g., log-optimality in Koolen and Grünwald 2022; Grünwald et al.
 2024; Larsson et al.
-2024) or of a single wealth super-martingale (e.g., admissibility in Ramdas et al.
-2022).
-The author hopes that this new viewpoint will inspire further research.
+2024, or admissibility in Ramdas et al.
+2022a).
+An alternative way to frame the first part of this work is as a characterisation of the family of single-round e-variables for testing the mean of a bounded random variable.
+In a similar spirit, since the appearance of the first version of this manuscript, two works have derived more general results for testing hypotheses defined by linear constraints (Clerico, 2024; Larsson et al., 2025), of which the present setting is a particular case.
+However, the approach we present here is direct and tailored to the simple setting considered, making it valuable for building intuition, while the more general results are significantly more abstract.
+We remark that testing via sequential betting with single-round e-variables is not the most general form of sequential testing with e-variables, which typically relies on multi-round e-variables and e-processes (Shafer, 2021; Koolen and Grünwald, 2022; Ramdas et al., 2022b; Ramdas and Wang, 2024).
+These allow for testing hypotheses over the entire data sequence, including assumptions about the dependence structure (e.g., i.i.d.
+or fixed conditional mean).
+In such cases, restricting the player to select a single-round e-variable at each step can be limiting (Koolen and Grünwald, 2022; Ramdas et al., 2022b).
+The second part of this work considers this broader setting.
+For technical convenience, we only focus on variables with finite discrete support, though we conjecture our conclusions hold more generally.
+We establish that when the sequence has fixed conditional mean, coin betting remains optimal even among testing methods based on multi-round e-variables and e-processes.
+However, we also show that this optimality result no longer holds under the more restrictive hypothesis of i.i.d.
+observations.
 Notation
-Endow the interval $[0,1]$ with the standard topology, and let $\mathcal{P}$ be the set of all Borel probability measures on $[0,1]$.
-For $x$ in $[0,1]$, $\delta_{x}$ denotes the Dirac unit mass on $x$.
-For $P\in\mathcal{P}$ and a Borel measurable function $f$ on $[0,1]$, $\langle P,f(X)\rangle$ (or more compactly $\langle P,f\rangle$) denotes the expectation of $f$ under $X\sim P$.
+We endow any Borel set $\mathcal{Z}\subseteq\mathbb{R}^{d}$ with the standard topology, and we denote as $\mathcal{P}_{\mathcal{Z}}$ the set of Borel probability measures on $\mathcal{Z}$.
+For $z\in\mathcal{Z}$, $\delta_{z}$ denotes the Dirac unit mass on $z$.
+For $P\in\mathcal{P}_{\mathcal{Z}}$ and a Borel measurable function $f$ on $\mathcal{Z}$, $\mathbb{E}_{P}[f(Z)]$ (or more compactly $\mathbb{E}_{P}[f]$) denotes the expectation of $f$ under $Z\sim P$.
+Given a sub-sigma-field $\mathcal{G}$, $\mathbb{E}_{P}[Z|\mathcal{G}]$ denotes the conditional expectation.
+Typically, $\mathcal{G}$ will be the sigma-field generated by some random variable $X$.
+In such case, we will write $\mathbb{E}_{P}[Z|X]$.
+We will also consider measures on product spaces.
+For $T\geq 1$, we write $\mathcal{P}_{\mathcal{Z}^{T}}$ the set of Borel probability measure on $\mathcal{Z}^{T}$ (endowed with the product topology).
+We will used $\otimes$ to denote the direct product of measures.
+For instance, given $P$ and $Q$ in $\mathcal{P}_{\mathcal{Z}}$, $P\otimes Q$ will be the element of $\mathcal{P}_{Z^{2}}$ that encode the law of $(Z_{1},Z_{2})$, where $Z_{1}\sim P$ and $Z_{2}\sim Q$ are independent.
+For any two given integers $s$ and $t$, with $s\leq t$, $[s:t]$ denotes the set of integers between $s$ and $t$ (included).
+For an integer $T$, given a vector $(x_{1},\dots,x_{T})$, we often represent it compactly as $x^{T}$ (upper indices).
+At times, we will also use the notation $x^{t:T}$ (with $t\in[1:T]$), to denote the vector $(x_{t},x_{t+1},\dots,x_{T})$.
 Sequences are denoted as $(s_{t})_{t\geq T_{0}}$, with $t$ an integer index and $T_{0}$ its smallest value (typically 00 or $1$).
+Sometimes we will also use the notation $s^{T_{0}:\infty}$ to denote $(s_{t})_{t\geq T_{0}}$, or simply $s^{\infty}$ if $T_{0}$ is clear from the context.
 For high probability statements, $\mathbb{P}$ expresses probability with respect to all the randomness involved.
 For instance, if $(X_{t})_{t\geq 1}$ is a sequence of i.i.d.
-draws from $P\in\mathcal{P}$, we write $\mathbb{P}(X_{t}\geq 1/2\,,\,\forall t\geq 1)$, with obvious meaning.
+draws from $P\in\mathcal{P}_{\mathcal{X}}$, we write $\mathbb{P}(X_{t}\geq 1/2\,,\,\forall t\geq 1)$, with obvious meaning.
 2 Algorithmic mean estimation
-Let $(X_{t})_{t\geq 1}$ be a sequence of independent random variables drawn from an unknown fixed distribution $P^{\star}\in\mathcal{P}$, whose mean $\mu^{\star}\in(0,1)$ has to be estimated.
+Fix a Borel set $\mathcal{X}\subseteq\mathbb{R}$.
+Let $(X_{t})_{t\geq 1}$ be a sequence of independent random variables drawn from an unknown fixed distribution $P^{\star}\in\mathcal{P}_{\mathcal{X}}$, whose mean $\mu^{\star}\in(0,1)$ has to be estimated.
 Let $\mathcal{F}=(\mathcal{F}_{t})_{t\geq 0}$ represent the natural filtration generated by this sequence, where $\mathcal{F}_{t}=\sigma(X_{1},\dots,X_{t})$ captures all information available up to time $t$.
@@
 Intuitively, this means that $(S_{t})_{t\geq 1}$ provides a set of plausible values for $\mu^{\star}$ at each time step $t$, while ensuring that the true mean remains in these sets indefinitely with high probability.
+We stress here that, for the time being, we assume that the sequence $(X_{t})_{t\geq 1}$ is i.i.d., and we shall not challenge this assumption, regardless of whether or not the observed data appear to exhibit such behaviour.
+In a way, this perspective aligns naturally with the goal of constructing confidence sequences, as simultaneously testing for the mean and the i.i.d.
+assumption might lead to rejecting everything and returning empty confidence sets because “the data do not look i.i.d.”.
+Under the i.i.d.
+model, mean estimation is well posed: we assume the existence of a fixed distribution $P^{\star}$ independently generating each observation, and our task is to estimate its mean.
+We remark that the independence assumption could be relaxed by only asking that each $X_{t}$ has fixed conditional mean $\mu^{\star}$ (to be estimated) given the past $\mathcal{F}_{t-1}$.
+The arguments we present next for the i.i.d.
+setting carry over with essentially no change to this less restrictive case.
+However, to keep the exposition clearer, we focus here solely on the independent case.
+The conditional setting will be addressed in the second part of this work, where we consider the more general scenario in which dependencies between observations are not assumed but are instead challenged by the statistical test.
 2.1 Coin-betting confidence sequence
@@
 The idea behind both papers is to set up a sequential coin-betting game for each candidate value $\mu\in(0,1)$ for the mean, and to exclude from the confidence sets those values $\mu$ that allow the player to earn unlikely high rewards.
-Definition 1 (Coin-betting game)
+Definition 1 (Coin-betting game).
 Fix $\mu\in(0,1)$ and let $I_{\mu}=[(\mu-1)^{-1},\mu^{-1}]$. Consider the following sequential procedure. At each round $t\geq 1$, a player
-  • 1.
   • picks [^2] We implicitly assume a measurable selection of $\beta_{t}$, namely the mapping $(x_{1},\dots x_{t-1})\mapsto\beta_{t}$ is Borel measurable.
 $\beta_{t}\in I_{\mu}$, based solely on the past observations $x_{1},\dots,x_{t-1}$;
-  • 2.
   • observes a new data-point $x_{t}\in[0,1]$;
-  • 3.
   • gets a reward $\log\big{(}1+\beta_{t}(x_{t}-\mu)\big{)}$.
@@
 The coin-betting approach is a particular instance of constructing confidence sequences via a sequential hypothesis testing procedure based on e-variables.
-We call hypothesis an arbitrary subset $\mathcal{H}$ of $\mathcal{P}$.
-An e-variable (with respect to $\mathcal{H}$) is any non-negative Borel measurable function $E:[0,1]\to[0,+\infty)$, such that
-$\langle P,E\rangle\leq 1\,,\qquad\forall P\in\mathcal{H}\,.$
+Here, we call hypothesis an arbitrary subset $\mathcal{H}$ of $\mathcal{P}_{[0,1]}$.
+An e-variable (for $\mathcal{H}$) is any non-negative Borel measurable function $E:[0,1]\to[0,+\infty)$, such that
+$\mathbb{E}_{P}[E]\leq 1\,,\qquad\forall P\in\mathcal{H}\,.$
 We denote as $\mathcal{E}_{\mathcal{H}}$ the set of all the e-variables with respect to $\mathcal{H}$, and we call e-class any subset of $\mathcal{E}_{\mathcal{H}}$.
 We remark that $\mathcal{E}_{\mathcal{H}}$ is never empty, since the constant function one is always an e-variable, independently of $\mathcal{H}$.
-Definition 2 (Testing game)
-Fix a hypothesis $\mathcal{H}\subseteq\mathcal{P}$ and an e-class $\mathcal{E}\subseteq\mathcal{E}_{\mathcal{H}}$.
-We call $\mathcal{E}$-restricted testing game (on $\mathcal{H}$) the following sequential procedure.
-At each round $t\geq 1$, a player
-  • 1.
-    picks [^2] Again, we assume that $E_{t}$ is picked in a measurable fashion, namely the mapping $(x_{1},\dots x_{t})\mapsto E_{t}(x_{t})$ is Borel measurable.
+Definition 2 (Testing-by-betting game).
+Fix a hypothesis $\mathcal{H}\subseteq\mathcal{P}_{[0,1]}$ and a non-empty e-class $\mathcal{E}\subseteq\mathcal{E}_{\mathcal{H}}$.
+We call $\mathcal{E}$-restricted testing-by-betting game (on $\mathcal{H}$) the following sequential procedure.
+At each round $t\geq 1$,
+  • the player picks [^3] Again, we assume that $E_{t}$ is picked in a measurable fashion, namely the mapping $(x_{1},\dots x_{t})\mapsto E_{t}(x_{t})$ is Borel measurable.
 an e-variable $E_{t}\in\mathcal{E}$ based solely on the past observations $x_{1},\dots,x_{t-1}$;
-  • 2.
-    observes a new data-point $x_{t}\in[0,1]$;
-  • 3.
-    gets a reward $\log E_{t}(x_{t})$.
-If $\mathcal{E}=\mathcal{E}_{\mathcal{H}}$, we speak of unrestricted testing game.
-The above game is at the core of e-variable testing (see, e.g., Ramdas and Wang 2024 for a recent overview), where one designs a test that rejects the hypothesis $\mathcal{H}$ whenever the total reward earned by the player gets excessively high.
+  • the player observes a new data-point $x_{t}\in[0,1]$;
+  • the player earns a reward $\log E_{t}(x_{t})$.
+If $\mathcal{E}=\mathcal{E}_{\mathcal{H}}$, we speak of unrestricted testing-by-betting game.
+The above game is an instance of e-variable testing (see Ramdas and Wang, 2024, for a recent overview), where one designs a test that rejects the hypothesis $\mathcal{H}$ whenever the total reward earned by the player gets excessively high.
 This procedure is justified by the fact that, if the data-points observed during the game were independently drawn from $P\in\mathcal{H}$, then the cumulative reward would be unlikely to grow very large.
-Lemma 1
-Let $\mathcal{H}\subseteq\mathcal{P}$ and consider a sequence $(X_{t}^{\prime})_{t\geq 1}$ of independent draws from some $P\in\mathcal{H}$.
+Lemma 1.
+Let $\mathcal{H}\subseteq\mathcal{P}_{[0,1]}$ and consider a sequence $(X_{t}^{\prime})_{t\geq 1}$ of independent draws from some $P\in\mathcal{H}$.
 Fix $\delta\in(0,1)$ and $\mathcal{E}\subseteq\mathcal{E}_{\mathcal{H}}$.
-Consider an $\mathcal{E}$-restricted testing game, where the observations are the sequence $(X_{t}^{\prime})_{t\geq 1}$.
+Consider an $\mathcal{E}$-restricted testing-by-betting game, where the observations are the sequence $(X_{t}^{\prime})_{t\geq 1}$.
 Let $R_{n}=\sum_{t=1}^{n}\log E_{t}(X_{t}^{\prime})$ represent the player’s cumulative reward at round $n$.
@@
 $\mathbb{P}\big{(}R_{n}\leq\log\tfrac{1}{\delta}\,,\;\forall n\geq 1\big{)}\geq 1-\delta\,.$
-Proof 1
-The result follows directly from Ville’s inequality, since $M_{n}=\prod_{t=1}^{n}E_{t}(X_{t}^{\prime})$ defines a non-negative super-martingale with respect to the natural filtration of the process $(X_{t}^{\prime})_{t\geq 1}$, with $M_{0}\equiv 1$$\qed$
-2.3 Mean estimation via e-variables
-We can leverage the testing game of Definition 2 to obtain a confidence sequence for the mean $\mu^{\star}\in(0,1)$ of $P^{\star}$. For $\mu\in(0,1)$, define the hypothesis
-$\mathcal{H}_{\mu}=\{P\in\mathcal{P}\,:\,\langle P,X\rangle=\mu\}\,,$ (1)
+Proof.
+The result follows directly from Ville’s inequality, since $M_{n}=\prod_{t=1}^{n}E_{t}(X_{t}^{\prime})$ defines a non-negative super-martingale with respect to the natural filtration of the process $(X_{t}^{\prime})_{t\geq 1}$, with $M_{0}\equiv 1$.
+∎
+2.3 Mean estimation via sequential betting
+We can leverage the testing-by-betting game of Definition 2 to obtain a confidence sequence for the mean $\mu^{\star}\in(0,1)$ of $P^{\star}$. For $\mu\in(0,1)$, define the hypothesis
+$\mathcal{H}_{\mu}=\{P\in\mathcal{P}_{[0,1]}\,:\,\mathbb{E}_{P}[X]=\mu\}\,,$ (1)
 which contains all probability measures on $[0,1]$ with mean $\mu$.
-For each $\mu$, we fix an e-class $\mathcal{E}_{\mu}\subseteq\mathcal{E}_{\mathcal{H}_{\mu}}$ and we consider an $\mathcal{E}_{\mu}$-restricted testing game on $\mathcal{H}_{\mu}$, where the player observes $(X_{t})_{t\geq 1}$, the sequence of draws from $P^{\star}$.
+For each $\mu$, we fix an e-class $\mathcal{E}_{\mu}\subseteq\mathcal{E}_{\mathcal{H}_{\mu}}$ and we consider an $\mathcal{E}_{\mu}$-restricted testing-by-betting game on $\mathcal{H}_{\mu}$, where the player observes $(X_{t})_{t\geq 1}$, the sequence of draws from $P^{\star}$.
 For each one of these games, denote as $R_{n}(\mu)$ the player’s cumulative reward at round $n$.
-Lemma 2
+Proposition 1.
 The sequence $(S_{n})_{n\geq 1}$, defined as
@@
 is a confidence sequence for the mean $\mu^{\star}$ of the data-generating probability measure $P^{\star}$.
-Proof 2
+Proof.
 For each $n\geq 1$, we have that $\{\mu^{\star}\in S_{n}\}=\{R_{n}(\mu^{\star})\leq\log\frac{1}{\delta}\}$, which is a $\mathcal{F}_{n}$-measurable event. Moreover,
 $\mathbb{P}\big{(}\mu^{\star}\in S_{n}\,,\;\forall n\geq 1\big{)}=\mathbb{P}\big{(}R_{n}(\mu^{\star})\leq\log\tfrac{1}{\delta}\,,\;\forall n\geq 1\big{)}\geq 1-\delta$
-by Lemma 1, as the observations are drawn from $P^{\star}\in\mathcal{H}_{\mu^{\star}}$. $\qed$
+by Lemma 1, as the observations are drawn from $P^{\star}\in\mathcal{H}_{\mu^{\star}}$. ∎
 We remark that the approach to derive confidence sequences via the coin-betting game, proposed by Orabona and Jun (2023) and Waudby-Smith and Ramdas (2023), is a specific instance of the framework that we have just described.
 Indeed, letting $E_{t}:x\mapsto 1+\beta_{t}(x-\mu)$, it is straightforward to verify that $E_{t}$ is non-negative on $[0,1]$, due to the restriction $\beta_{t}\in I_{\mu}=[(\mu-1)^{-1},\mu^{-1}]$ in Definition 1.
-Moreover, for any $P\in\mathcal{H}_{\mu}$, we have $\langle P,E_{t}\rangle=1$, which implies that $E_{t}$ is an e-variable for $\mathcal{H}_{\mu}$.
+Moreover, for any $P\in\mathcal{H}_{\mu}$, we have $\mathbb{E}_{P}[E_{t}]=1$, which implies that $E_{t}$ is an e-variable for $\mathcal{H}_{\mu}$.
 Finally, the reward in the coin-betting game is precisely equal to $\log E_{t}(x_{t})$.
@@
 This point naturally raises the question: “Does restricting the player to the coin-betting e-class (2) loosen the confidence sequence?”.
-Interestingly, the answer turns out to be negative, as we will demonstrate throughout the remainder of this letter.
+Interestingly, for the round-wise testing-by-betting framework that we are considering the answer turns out to be negative.
 3 Majorising e-classes and optimal e-class
-We introduce a partial ordering on the set of real functions on $[0,1]$.
-Given two functions $f,f^{\prime}:[0,1]\to\mathbb{R}$, we say that $f$ majorises $f^{\prime}$, and write $f\succeq f^{\prime}$, if $f(x)\geq f^{\prime}(x)$ for all $x\in[0,1]$.
-If $f\succeq f^{\prime}$ and there exists a point $x\in[0,1]$ such that $f(x)>f^{\prime}(x)$, we say that $f$ is a strict majoriser of $f^{\prime}$, and we write $f\succ f^{\prime}$.
-Definition 3
+We introduce a partial ordering on the set of real functions on a set $\mathcal{X}$.
+Given two functions $f,f^{\prime}:\mathcal{X}\to\mathbb{R}$, we say that $f$ majorises $f^{\prime}$, and write $f\succeq f^{\prime}$, if $f(x)\geq f^{\prime}(x)$ for all $x\in\mathcal{X}$.
+If $f\succeq f^{\prime}$ and there exists a point $x\in\mathcal{X}$ such that $f(x)>f^{\prime}(x)$, we say that $f$ is a strict majoriser of $f^{\prime}$, and we write $f\succ f^{\prime}$.
+Definition 3.
 $E\in\mathcal{E}_{\mathcal{H}}$ is called a maximal e-variable when there is no $E^{\prime}\in\mathcal{E}_{\mathcal{H}}$ such that $E^{\prime}\succ E$.
 Next, we introduce a way to compare different e-classes.
-Definition 4
-Given two e-classes $\mathcal{E}$ and $\mathcal{E}^{\prime}$, we say that $\mathcal{E}$ majorises $\mathcal{E}^{\prime}$ if, for any $E^{\prime}\in\mathcal{E}$, there is an e-variable $E\in\mathcal{E}$ such that $E\succeq E^{\prime}$.
+Definition 4.
+Given two e-classes $\mathcal{E}$ and $\mathcal{E}^{\prime}$, we say that $\mathcal{E}$ majorises $\mathcal{E}^{\prime}$ if, for any $E^{\prime}\in\mathcal{E}^{\prime}$, there is an e-variable $E\in\mathcal{E}$ such that $E\succeq E^{\prime}$.
 We call majorising e-class any e-class that majorises $\mathcal{E}_{\mathcal{H}}$.
-Lemma 3
+Lemma 2.
 Every majorising e-class contains all the maximal e-variables.
-Proof 3
+Proof.
 Let $E\in\mathcal{E}_{\mathcal{H}}$ be maximal and $\mathcal{E}$ a majorising e-class.
@@
 Hence, $E\in\mathcal{E}$.
+∎
 The significance of the notion of majorising e-class for our problem is straightforward: if $\mathcal{E}$ majorises $\mathcal{E}^{\prime}$, then any strategy in an $\mathcal{E}^{\prime}$-restricted game can be matched or outperformed (in terms of rewards) by a corresponding strategy in the $\mathcal{E}$-restricted game, regardless of the sequence of observations.
-This allows us to compare how the restriction to different e-classes affects the testing game of Definition 2.
-Definition 5
-If a majorising e-class is contained in every other majorising e-class, it is called optimal.
+This allows us to compare how the restriction to different e-classes affects the testing-by-betting game of Definition 2.
+Definition 5.
+If a majorising e-class is contained in every other majorising e-class, we call it optimal.
 For any $\mathcal{H}$, a majorising e-class always exists, as $\mathcal{E}_{\mathcal{H}}$ itself is a majorising e-class.
 However, an optimal e-class may not exist.
+[^4] We refer the reader to the follow-up work Clerico (2024) for an example of non-existence of the optimal e-class.
+Proposition 2 therein shows that =$\mathcal{H}=\{P\in\mathcal{P}_{[0,1]}\,:\,P(\{0\})\geq 1/2\}\cup\{U_{[0,1]}\}$ (with $U_{[0,1]}$ the uniform distribution on $[0,1]$) is a hypothesis for which the set of maximal e-variables is not a majorising e-class.
+Hence, $\mathcal{H}$ does not admit an optimal e-class.
 Next, we state a sufficient and necessary condition for the existence of optimal e-classes.
-Lemma 4
+Lemma 3.
 An optimal e-class exists if, and only if, the set of all maximal e-variables is a majorising e-class.
 If an optimal e-class exists, it is unique, it corresponds to the set of all maximal e-variables, and it is the only majorising e-class whose elements are all maximal.
-Proof 4
+Proof.
 Denote as $\hat{\mathcal{E}}$ the set of all the maximal e-variables.
@@
 So, $\hat{\mathcal{E}}$ is contained in all the majorising e-classes, and hence it is optimal.
-Now, the final statements are a trivial consequence of what shown above and Lemma 3$\qed$
-Let us emphasise once more that, from our discussion thus far, it is clear that restricting the testing game of Definition 2 to a majorising e-class does not hinder the performance of the player, as for any unrestricted strategy $(E_{t})_{t\geq 1}$ they can always pick a restricted strategy $(E_{t}^{\prime})_{t\geq 1}$, whose cumulative rewards inevitably match or outperform those of $(E_{t})_{t\geq 1}$, regardless of the sequence of observations.
+Now, the final statements are a trivial consequence of what was shown above and Lemma 2. ∎
+Let us emphasise once more that, from our discussion thus far, it is clear that restricting the testing-by-betting game of Definition 2 to a majorising e-class does not hinder the performance of the player, as for any unrestricted strategy $(E_{t})_{t\geq 1}$ they can always pick a restricted strategy $(E_{t}^{\prime})_{t\geq 1}$, whose cumulative rewards inevitably match or outperform those of $(E_{t})_{t\geq 1}$, regardless of the sequence of observations.
 From a practical perspective, identifying the optimal e-class, when it exists, greatly simplifies the design of an effective strategy by narrowing the player’s choice to the best possible e-variables.
@@
 Conversely, when a player selects an e-variable from the optimal e-class, no other choice can be guaranteed to be better before observing $x_{t}$, since the player’s pick is a maximal e-variable.
-As a straightforward consequence, the optimal approach to obtain confidence sequences consists in restricting the testing game to the optimal e-class, for each mean candidate value $\mu$.
-We will show next that this coincides precisely with the coin-betting formulation.
+As a straightforward consequence, within this round-wise testing-by-betting approach, the optimal approach to obtain confidence sequences consists in restricting game to the optimal e-class, for each mean candidate value $\mu$.
+We show next that this coincides precisely with the coin-betting formulation.
 4 Optimality of the coin-betting e-class
@@
 We now show that the optimal e-class for $\mathcal{H}_{\mu}$ exists and coincides with the coin-betting e-class $\mathcal{E}^{\mathrm{cb}}_{\mu}$, defined in (2).
-Lemma 5
+Lemma 4.
 Fix $\mu\in(0,1)$ and consider the function $F_{\mu}:[0,1]\to[1,+\infty)$ defined as
+[…]
+Figure 1: Pictorial representation of the main step in the proof of Theorem 1.
+$1+\beta_{1}(x-\mu)$ dominates $E$ for $x\in[0,\mu)$, while $1+\beta_{0}(x-\mu)$ dominates $E$ for $x\in(\mu,1]$.
+If there is $\beta^{\star}\in(\beta_{0},\beta_{1})$, then we can find $u_{0}\in[0,\mu)$ and $u_{1}\in(\mu,1]$ such that $E(u_{0})>1+\beta^{\star}(u_{0}-\mu)$ and $E(u_{1})>1+\beta^{\star}(u_{1}-\mu)$, represented by the points $U_{0}$ and $U_{1}$ being above the purple line $1+\beta^{\star}(x-\mu)$.
+The probability measure $\hat{P}$ supported on $\{u_{0},u_{1}\}$ with mean $\mu$ is in $\mathcal{H}_{\mu}$.
+We have $\mathbb{E}_{\hat{P}}[E]>1$.
+Indeed, this expected value corresponds to the vertical coordinate of the point $U^{\star}$, the intersection of the line connecting $U_{0}$ and $U_{1}$ with the vertical line at $x=\mu$.
+This is a contradiction if $E$ is an e-variable, in which case it must be that $\beta_{0}\geq\beta_{1}$.
 $F_{\mu}\;:\;x\mapsto\begin{cases}1+\frac{1}{\mu}(x-\mu)&\text{if $x\geq\mu$;}\\ 1+\frac{1}{\mu-1}(x-\mu)&\text{if $x<\mu$.}\end{cases}$
 For any $x\in[0,1]$ there is $P_{x}\in\mathcal{H}_{\mu}$ such that $P_{x}(\{x\})=1/F_{\mu}(x)>0$. Moreover, $F_{\mu}\succeq E$ for every $E\in\mathcal{E}_{\mathcal{H}_{\mu}}$.
-Proof 5
+Proof.
 For $x\geq\mu$, let $P_{x}=\frac{\mu}{x}\delta_{x}+(1-\frac{\mu}{x})\delta_{0}$.
@@
 With a similar reasoning, if $x<\mu$ we can find a measure $P_{x}$ in $\mathcal{H}_{\mu}$, supported on $\{x,1\}$, with mass $1/F_{\mu}(x)$ on $x$.
-Now, to check that $F$ majorises all the e-variables, fix $E\in\mathcal{E}_{\mathcal{H}_{\mu}}$ and $x\in[0,1]$.
+Now, to check that $F_{\mu}$ majorises all the e-variables, fix $E\in\mathcal{E}_{\mathcal{H}_{\mu}}$ and $x\in[0,1]$.
 Consider a $P_{x}\in\mathcal{H}_{\mu}$ with mass $1/F_{\mu}(x)$ on $x$.
-Then, $1\geq\langle P_{x},E\rangle\geq P_{x}(x)E(x)=E(x)/F_{\mu}(x)$, and we conclude.
-$\qed$
-Theorem 1
+Then, $1\geq\mathbb{E}_{P_{x}}[E]\geq P_{x}(x)E(x)=E(x)/F_{\mu}(x)$, and we conclude.
+∎
+Theorem 1.
 For any $\mu\in(0,1)$, the coin-betting e-class $\mathcal{E}^{\mathrm{cb}}_{\mu}$ is the optimal e-class for $\mathcal{H}_{\mu}$.
-Proof 6
-First, let us show that $\mathcal{E}^{\mathrm{cb}}_{\mu}$ is a majorising e-class. Fix an arbitrary $E\in\mathcal{E}_{\mathcal{H}_{\mu}}$. Define the sets
-$B_{0}=\left\{\beta\in I_{\mu}\;:\;\inf_{x\in[0,\mu]}\big{(}E_{\beta}(x)-E(x)\big{)}\geq 0\right\}\qquad\text{and}\qquad B_{1}=\left\{\beta\in I_{\mu}\;:\;\inf_{x\in[\mu,1]}\big{(}E_{\beta}(x)-E(x)\big{)}\geq 0\right\}\,,$
-where we recall that $I_{\mu}=[(1-\mu)^{-1},\mu^{-1}]$ and $E_{\beta}:x\mapsto 1+\beta(x-\mu)$.
+Proof.
+First, let us show that $\mathcal{E}^{\mathrm{cb}}_{\mu}$ is a majorising e-class.
+For a pictorial representation of this part of the proof, the reader is invited to look at Figure 1.
+Fix an arbitrary $E\in\mathcal{E}_{\mathcal{H}_{\mu}}$.
+Define the sets
+$B_{0}=\left\{\beta\in I_{\mu}\;:\;\inf_{x\in[0,\mu)}\big{(}E_{\beta}(x)-E(x)\big{)}\geq 0\right\}\quad\text{and}\quad B_{1}=\left\{\beta\in I_{\mu}\;:\;\inf_{x\in(\mu,1]}\big{(}E_{\beta}(x)-E(x)\big{)}\geq 0\right\}\,,$
+where we recall that $I_{\mu}=[(\mu-1)^{-1},\mu^{-1}]$ and $E_{\beta}:x\mapsto 1+\beta(x-\mu)$.
 Both sets are closed and convex (as they are intersections of closed and convex sets).
-By Lemma 5, $E_{(\mu-1)^{-1}}\in B_{0}$ and $E_{\mu^{-1}}\in B_{1}$, so $B_{0}=[(\mu-1)^{-1},\beta_{0}]$ and $B_{1}=[\beta_{1},\mu]$, for some $\beta_{0}$ and $\beta_{1}$ in $I_{\mu}$.
+By Lemma 4, $(\mu-1)^{-1}\in B_{0}$ and $\mu^{-1}\in B_{1}$, so $B_{0}=[(\mu-1)^{-1},\beta_{0}]$ and $B_{1}=[\beta_{1},\mu^{-1}]$, for some $\beta_{0}$ and $\beta_{1}$ in $I_{\mu}$.
 We will now show that $B_{0}\cap B_{1}\neq\varnothing$, or equivalently that $\beta_{0}\geq\beta_{1}$.
@@
 Then, $\beta^{\star}\notin B_{0}$ and $\beta^{\star}\notin B_{1}$.
-In particular, there are $x_{0}\in[0,\mu]$ and $x_{1}\in[\mu,1]$ such that $E(x_{0})>E_{\beta^{\star}}(x_{0})$ and $E(x_{1})>E_{\beta^{\star}}(x_{1})$.
-Notice that since $\delta_{\mu}\in\mathcal{H}_{\mu}$, it must be that $E(\mu)\leq 1$, and so $x_{0}<\mu$ and $x_{1}>\mu$.
-Hence, $\mu\in(x_{0},x_{1})$ and so there is $\hat{P}\in\mathcal{H}_{\mu}$ with support $\{x_{0},x_{1}\}$.
-Notice that $\langle\hat{P},E_{\beta^{\star}}\rangle=1+\beta^{\star}(\langle\hat{P},X\rangle-\mu)=1$.
-But $E$ is strictly larger than $E_{\beta^{\star}}$ on $\mathrm{Supp}(\hat{P})$, and so $\langle\hat{P},E\rangle>1$, which is a contradiction since $E$ is an e-variable.
+In particular, there are $u_{0}\in[0,\mu)$ and $u_{1}\in(\mu,1]$ such that $E(u_{0})>E_{\beta^{\star}}(u_{0})$ and $E(u_{1})>E_{\beta^{\star}}(u_{1})$.
+As $\mu\in(u_{0},u_{1})$, there is $\hat{P}\in\mathcal{H}_{\mu}$ with support $\{u_{0},u_{1}\}$.
+Note that $\mathbb{E}_{\hat{P}}[E_{\beta^{\star}}]=1+\beta^{\star}(\mathbb{E}_{\hat{P}}[X]-\mu)=1$.
+But $E$ is strictly larger than $E_{\beta^{\star}}$ on $\mathrm{Supp}(\hat{P})$, and so $\mathbb{E}_{\hat{P}}[E]>1$, which is a contradiction since $E$ is an e-variable.
 We conclude that $\beta_{0}\geq\beta_{1}$ and so there exists $\hat{\beta}\in B_{0}\cap B_{1}$.
-By construction, $E_{\hat{\beta}}\in\mathcal{E}^{\mathrm{cb}}_{\mu}$ and $E_{\hat{\beta}}\succeq E$.
-Because the choice of $E$ was arbitrary, $\mathcal{E}^{\mathrm{cb}}_{\mu}$ is a majorising e-class.
+By construction, $E_{\hat{\beta}}\in\mathcal{E}^{\mathrm{cb}}_{\mu}$ and $E_{\hat{\beta}}(x)\geq E(x)$ for all $x\in[0,1]$ different from $\mu$.
+For $x=\mu$, by Lemma 4 we have $E(\mu)\leq F_{\mu}(\mu)=1=E_{\hat{\beta}}(\mu)$, and so $E\preceq E_{\hat{\beta}}$.
+As the choice of $E$ was arbitrary, $\mathcal{E}^{\mathrm{cb}}_{\mu}$ is a majorising e-class.
 Now that we have established that $\mathcal{E}^{\mathrm{cb}}_{\mu}$ is a majorising e-class, by Lemma 3 we only need to show that all its elements are maximal to conclude that it is optimal.
@@
 By Lemma 4, there is $P\in\mathcal{H}_{\mu}$ such that $P(\{x\})>0$.
-Since $E\in\mathcal{E}^{\mathrm{cb}}_{\mu}$, we have $\langle P,E\rangle=1$, and so $0\leq P(\{x\})(\hat{E}(x)-E(x))\leq\langle P,\hat{E}-E\rangle=\langle P,\hat{E}\rangle-1\leq 0$.
+Since $E\in\mathcal{E}^{\mathrm{cb}}_{\mu}$, we have $\langle P,E\rangle=1$, and so $0\leq P(\{x\})(\hat{E}(x)-E(x))\leq\mathbb{E}_{P}[\hat{E}-E]=\mathbb{E}_{P}[\hat{E}]-1\leq 0$.
 Since $P(\{x\})>0$, we get $\hat{E}(x)=E(x)$ and, $x$ being arbitrary, $\hat{E}=E$.
 Hence $E$ is maximal, as it has no strict majoriser.
-$\qed$
+∎
+5 Optimality beyond single-round e-variables
+Up to this point, we have focused on hypotheses defined as subsets of the space $\mathcal{P}_{[0,1]}$, which cannot capture dependencies across multiple rounds.
+In such setting, the dependency structure among observations is fixed and assumed a priori, rather than being subject to testing.
+Yet, e-variable-based testing naturally extends to more general hypotheses that span multiple rounds and can account for sequential or dependent data structures.
+In the second part of the paper, we turn our attention to this richer framework.
+Remarkably, a slight adaptation of the proof strategy used for Theorem 1 and depicted in Figure 1 allows us to show that coin-betting-based testing is also optimal in the setting where the sequence of observations has a fixed conditional mean $\mu$.
+We will first show that this is the case with multi-round e-variables for a fixed time horizon, and then extend the result to testing with e-processes.
+Conversely, we will show that when testing the i.i.d.
+assumption, the coin-betting approach no longer yields the optimal e-class.
+For technical reasons (primarily related to measurability, see discussion in Section 7), for the rest of this work we restrict ourselves to the case of a random variable taking values in a finite (known) set.
+Therefore, henceforth we let $\mathcal{X}\subseteq\mathbb{R}$ denote a finite set.
+For simplicity, we assume that $\min\mathcal{X}=0$ and $\max\mathcal{X}=1$, although all our results extend readily to any case where the minimum and maximum values are finite.
+We note that since $\mathcal{X}$ has finite cardinality, every function $\mathcal{X}^{t}\to\mathbb{R}$ is automatically Borel measurable.
+5.1 Optimal e-class for multi-round e-variables
+For any $\mu\in(0,1)$ and $T\geq 1$, we let
+$\mathcal{H}_{\mu}^{T}=\big{\{}P\in\mathcal{P}_{\mathcal{X}^{T}}\,:\,\mathbb{E}_{P}[X_{t}|X^{t-1}]=\mu\,,\;\forall t\in[1:T]\big{\}}\,,$
+where $\mathbb{E}_{P}[X_{1}|X^{0}]=\mathbb{E}_{P}[X_{1}]$.
+Note that $\mathcal{H}_{\mu}^{T}$ is a hypothesis on $\mathcal{X}^{T}$.
+As such, its set of e-variables will consist of Borel functions from $\mathcal{X}^{T}$ to $\mathbb{R}$.
+We denote the set of all e-values relative to $\mathcal{H}_{\mu}^{T}$ as $\mathcal{E}_{\mu}^{T}$.
+In a coin-betting game, the wealth of the player at round $T$ is a non-negative function of the observations, which takes the form $M_{T}(x^{T})=\prod_{t=1}^{n}E_{t}(x_{t})$, where $E_{t}$ is in the form $E_{\lambda_{t}}$, with $\lambda_{t}$ chosen as a function of the past observations $x^{t-1}$.
+It is straightforward to check that under the hypothesis $\mathcal{H}_{\mu}^{T}$, the expectation of $M_{T}$ is always exactly one ($M_{T}$ is a martingale), which shows that $M_{T}$ is an e-variable.
+Next, we show that the e-class of e-variables in this form coincides is the optimal e-class for $\mathcal{H}_{\mu}^{T}$.
+We let $\Lambda^{T}_{\mu}$ be a set of $T$-tuples of functions defined as
+$\Lambda^{T}_{\mu}=\big{\{}\lambda^{T}=(\lambda_{1},\dots,\lambda_{T})\,:\,\text{$\lambda_{t}$ is a function from $\mathcal{X}^{t-1}$ to $I_{\mu}$}\big{\}}\,.$
+Given $\lambda^{T}\in\Lambda^{T}_{\mu}$, we define the e-variable $E_{\lambda^{T}}\in\mathcal{E}^{T}_{\mu}$ as
+$E_{\lambda^{T}}(x^{T})=\prod_{t=1}^{T}\big{(}1+\lambda_{t}(x^{t-1})(x_{t}-\mu)\big{)}\,.$
+We define the T round coin-betting e-class $\mathcal{E}^{\mathrm{cb},T}_{\mu}=\{E_{\lambda^{T}}\,:\,\lambda^{T}\in\Lambda_{\mu}^{T}\}$.
+Theorem 2.
+Fix any $T\geq 1$ and $\mu\in(0,1)$. $\mathcal{E}^{\mathrm{cb},T}_{\mu}$ is the optimal e-class for $\mathcal{H}_{\mu}^{T}$.
+The proof of Theorem 2 shares many similarity with that of Theorem 1.
+In particular, we use the same technique to establish the maximality of e-variables of the form $E_{\lambda^{T}}$.
+A geometric argument analogous to the one illustrated in Figure 1 is then employed, as part of an induction recursion, to prove that the e-class of interest is majorising.
+However, the detailed proof being somewhat lengthy, we defer it to the appendix.
+5.2 A remark on the i.i.d. case
+From what we have shown so far, in a $T$ round game coin-betting yields all and only the maximal e-variables, when testing the hypothesis that the conditional mean is some fixed $\mu\in(0,1)$.
+Yet, this is not any more the case if we consider a more restrictive hypotheses, stating that the observations are i.i.d.
+and with mean $\mu$.
+More concretely, let
+^$\hat{\mathcal{H}}_{\mu}^{T}=\big{\{}P=Q^{\otimes T}\,:\,Q\in\mathcal{H}_{\mu}\big{\}}\,$
+[…]
+To give a concrete example of the mismatch between the e-variables for the conditional and independent scenarios, we consider the “highly symmetrical” case of $\mathcal{X}=\{0,1/2,1\}$, with $\mu=1/2$ and $T=2$.
+We show at the end appendix that the e-variables in this setting are the functions $E:\{0,1/2,0\}^{2}\to[0,+\infty)$ that satisfy
+$\xi_{0}\leq 1\,;\qquad\xi_{1}\leq 1+\sqrt{(1-\xi_{0})(1-\xi_{2})}\,;\qquad\xi_{2}\leq 1\,,$
+where $\xi_{0}=E(1/2,1/2)$, $\xi_{1}=(E(1,1/2)+E(1/2,1)+E(1/2,0)+E(0,1/2))/4$, and $\xi_{2}=(E(1,1)+E(1,0)+E(0,1)+E(0,0))/4$.
+Clearly, the optimal e-class here consists of those e-variables satisfying the constraint on $\xi_{1}$ with equality.
+We note that all the e-variables in $\mathcal{E}_{1/2}^{T}$ satisfy $\xi_{0}=\xi_{1}=\xi_{2}=1$.
+As expected, these are indeed maximal e-variables for ^$\hat{\mathcal{H}}_{1/2}^{2}$, but there are maximal e-variables for ^$\hat{\mathcal{H}}_{1/2}^{2}$ that are not in $\mathcal{E}_{1/2}^{T}$.
+The function that equals $4$ on $x^{2}=(1,1/2)$, and 00 everywhere else, is an example of a (maximal) e-variable for ^$\hat{\mathcal{H}}_{1/2}^{2}$, which is not an e-variable for $\mathcal{H}_{1/2}^{2}$.
+We leave as a non-trivial open question to characterise the e-variables for ^$\hat{\mathcal{H}}^{T}_{\mu}$ in the general case.
+6 E-processes for testing the fixed conditional mean
+A key advantage of many testing procedures involving e-variables is anytime validity.
+One way to achieve this is through the round-wise testing-by-betting approach, which gives rise to a super-martingale (cf.
+Lemma 1).
+However, as we have already pointed out, this is not the most general approach.
+Rather than only focusing on super-martingales, one can instead use any non-negative process whose expectation is upper bounded by one, regardless the stopping time.
+These are known as e-processes, a generalization of super-martingales, which can enable powerful tests even in scenarios where martingale-based methods lack power (Ramdas et al., 2022b).
+As in the previous section, here $\mathcal{X}$ denotes a finite subset of $[0,1]$, with $\min\mathcal{X}=0$ and $\max\mathcal{X}=1$.
+We denote as $\mathcal{X}^{\infty}$ the space of sequences $(x_{t})_{t\geq 1}\subseteq\mathcal{X}$.
+For convenience, we will sometimes write $x^{\infty}$ to mean $(x_{t})_{t\geq 1}$.
+We endow the set $\mathcal{X}^{\infty}$ with the product sigma-field generated by cylinder sets, and we denote as $\mathcal{P}_{\mathcal{X}^{\infty}}$ the space of probability measures on $\mathcal{X}^{\infty}$.
+For a fixed $\mu\in(0,1)$, in this section we will focus on the hypothesis class
+$\mathcal{H}_{\mu}^{\infty}=\big{\{}P\in\mathcal{P}_{\mathcal{X}^{\infty}}\,:\,\mathbb{E}_{P}[X_{t}|X^{t-1}]=\mu\,,\;\forall t\geq 1\big{\}}\,.$
+This time, we will not aim to study the e-variables for $\mathcal{H}^{\infty}_{\mu}$, as these would be functions that take a whole sequence as argument.
+Conversely, we are interested in tools that allow us to stop the test if we think enough data-points have been observed to decide whether or not the hypothesis has to be rejected.
+To make this rigorous we first need some definitions.
+First, we call a finite stopping time a function $\tau:\mathcal{X}^{\infty}\to\mathbb{N}$ such that, whenever for some $x^{\infty}$ and $y^{\infty}$ in $\mathcal{X}^{\infty}$ and some $t\in\mathbb{N}$ we have $\tau(x^{\infty})=\tau(y^{\infty})=t$, then $x^{t}=y^{t}$.
+[^8] It is straightforward to verify that in our setting this definition is equivalent to the perhaps more classical notion of finite e-process as a $\tau$ such that the sets $\{\tau\leq t\}$ are $\mathcal{F}_{t}$ measurable.
+We say that $\tau$ is bounded by $T\geq 0$ if its image is contained in $[0:T]$.
+We let $\mathcal{T}_{T}$ denote the set of all stopping times bounded by $T$.
+Definition 6.
+An e-process (for $\mathcal{H}_{\mu}^{\infty}$) is a sequence $(E_{t})_{t\geq 0}$, of non-negative functions $E_{t}:\mathcal{X}^{t}\to[0,+\infty)$, such that, for every $P\in\mathcal{H}^{\infty}_{\mu}$, any $T\geq 0$ and any $\tau\in\mathcal{T}_{T}$,
+$\mathbb{E}_{P}[E_{\tau}]\leq 1\,.$
+We denote as $\mathcal{E}_{\mu}^{\infty}$ the set of all the e-processes for $\mathcal{H}_{\mu}^{\infty}$.
+We call e-process class any subset of $\mathcal{E}_{\mu}^{\infty}$.
+Similarly to what we have done for the e-variables, we say that the e-process $E$ majorises the e-process $E^{\prime}$ (we write $E\succeq E^{\prime}$) if for every $t\geq 0$ we have $E_{t}\succeq E^{\prime}_{t}$.
+We say that an e-process class $\mathcal{E}$ is majorising if for every $E\in\mathcal{E}^{\infty}_{\mu}$, there is $E^{\prime}\in\mathcal{E}$ such that $E\preceq E^{\prime}$.
+We say that $\mathcal{E}$ is optimal if it is majorising and is contained in every other majorising e-process class.
+We call an e-process $E$ maximal if it is not strictly majorised by any other e-process.
+With the same reasoning as in Lemma 3, one can show that the optimal e-process class exists if, and only if, the class of all the maximal e-processes is majorising.
+In such case the optimal class is unique and coincides precisely with all and only the majorising e-processes.
+We let $\Lambda_{\mu}^{\infty}$ be the set of all sequences $\lambda^{\infty}=(\lambda_{t})_{t\geq 1}$, such that $\lambda_{t}:\mathcal{X}^{t-1}\to I_{\mu}$.
+For any $\lambda^{\infty}\in\Lambda_{\mu}^{\infty}$, we define $E^{\lambda^{\infty}}=(E^{\lambda^{\infty}}_{T})_{T\geq 0}$ via
+$E^{\lambda^{\infty}}_{T}(x^{T})=E_{\lambda^{T}}(x^{T})=\prod_{t=1}^{T}\big{(}1+\lambda_{t}(x^{t-1})(x_{t}-\mu)\big{)}\,,$
+for $T\geq 1$, and $E_{0}=1$.
+These are the e-processes associated to the wealth of a coin-betting player.
+As $E^{\lambda^{\infty}}$ defines a martingale under $\mathcal{H}_{\mu}^{\infty}$, it is also an e-process for such hypothesis.
+We define the coin-betting e-process class $\mathcal{E}^{\mathrm{cb},\infty}_{\mu}=\{E^{\lambda^{\infty}}\,:\,\lambda^{\infty}\in\Lambda^{\infty}_{\mu}\}$.
+The next result shows that for $\mathcal{H}_{\mu}^{\infty}$ the coin betting approach is optimal also at the level of e-processes.
+The proof (reported in the appendix) always follows the ideas introduced in the proof of Theorem 1.
+Theorem 3.
+For any $\mu\in(0,1)$, $\mathcal{E}^{\mathrm{cb},\infty}_{\mu}$ is the optimal e-process class for $\mathcal{H}_{\mu}^{\infty}$.
 7 Perspectives
 Theorem 1 gives a rigorous sense to the claim that the coin-betting formulation is optimal among the e-variable-based approaches to build confidence sequences for $\mu^{\star}$.
-To formalise this assertion, we introduced the notions of majorising and optimal e-classes, which, in the author’s view, are of independent interest in the context of sequential testing, beyond the scope of this letter.
+Theorem 2 and Theorem 3 extended this optimality result to testing the fixed conditional mean of a sequence of bounded observations.
+To formalise these claims, we introduced the notions of majorising and optimal e-classes, which, in the author’s view, are of independent interest in the context of sequential testing, beyond the scope of this paper.
 The main novelty of these concepts lies in defining “optimality” in terms of e-classes, rather than individual e-variables.
 This perspective contrasts with the notion of log-optimality of a single e-variable with respect to an alternative hypothesis, widely discussed in the literature (e.g., Koolen and Grünwald, 2022; Grünwald et al., 2024; Larsson et al., 2024).
-Rather than log-optimality, what probably best matches our way of “ranking” e-variables is the idea of admissibility introduced by Ramdas et al.
-(2022).
-Roughly speaking, let us say that a wealth-process (namely, a non-negative adapted process starting at one that is a super-martingale under every $P\in\mathcal{H}$) is admissible if it is not dominated by any other wealth-process.
-This concept is tightly related to our notion of maximal e-variables, as it is not hard to see that any wealth-process $(M_{n})_{n\geq 1}$ on $\mathcal{H}_{\mu}$ is admissible if, and only if, it can be associated with a sequence $(E_{t})_{t\geq 1}$ of maximal e-variables such that $M_{n}=\prod_{t=1}^{n}E_{t}(X_{t}^{\prime})$, with $(X_{t}^{\prime})_{t\geq 1}$ a sequence generated by some $P\in\mathcal{H}_{\mu}$.
-Thus, in our setting, Theorem 1 provides a characterisation of the set of all admissible wealth-processes for $\mathcal{H}_{\mu}$.
-Let us emphasise that the tools from Ramdas et al.
-(2022) alone are insufficient to achieve this, as the necessary conditions for admissibility provided by their work rely on the existence of a “reference measure” for the hypothesis, [^3] A hypothesis $\mathcal{H}$ admits a “reference measure” if there exists a Borel measure $Q$ such that $P\ll Q$ for all $P\in\mathcal{H}$.
-a property that $\mathcal{H}_{\mu}$ does not satisfy.
-More precisely, their results are enough to prove that every e-variable in the coin-betting e-class is maximal, but not that $\mathcal{E}^{\mathrm{cb}}_{\mu}$ is a majorising e-class.
+The concept of maximality for e-variables, as introduced in this paper, is essentially equivalent to the classical statistical notion of admissibility.
+An e-variable is considered maximal if no other e-variable strictly dominates it.
+Likewise, the idea of a majorising e-class corresponds to the classical concept of a complete class, with the optimal e-class representing the minimal complete class of e-variables.
+For a detailed discussion of these ideas in traditional statistics we refer to Lehmann and Romano (2022).
+In this work, however, we adopt terminology from partially ordered set (poset) theory to highlight the fact that with e-variables these properties follows from the classical partial order of dominance among functions.
+The concept of admissibility in the context of e-variables and e-processes was previously introduced by Ramdas et al.
+(2022a), whose definition is closely aligned with our notion of maximality.
+The key distinction lies in the type of dominance considered: their framework often relies on almost sure dominance, whereas our definition of maximality requires everywhere dominance.
+[^9] The two definitions becomes of course equivalent when $\mathcal{X}$ is finite, but they are not in the case of $\mathcal{X}=[0,1]$, which is the setting of our analysis on single-round e-variables.
+This stricter requirement is motivated by the will of developing a theory that remains valid even in the absence of a known alternative, where any value in $\mathcal{X}$ might occur at the next observation.
+Although Ramdas et al.
+(2022a) provide some necessary and sufficient conditions for admissibility, we emphasize that their results are insufficient to establish the optimality of the coin-betting e-class.
+Specifically, their approach assumes the existence of a reference measure for the hypothesis, [^10] A hypothesis $\mathcal{H}$ admits a ”reference measure” if there exists a Borel measure $Q$ such that $P\ll Q$ for all $P\in\mathcal{H}$.
+a condition not satisfied by $\mathcal{H}_{\mu}$.
+More precisely, their results can show that each individual e-variable in the coin-betting e-class is maximal, but not that $\mathcal{E}^{\mathrm{cb}}_{\mu}$ forms a majorising e-class.
+This distinction, though subtle, is crucial: proving the optimality of coin-betting requires analyzing the collective properties of a set of e-variables (or equivalently, wealth processes), rather than evaluating them in isolation.
+It is precisely the fact that $\mathcal{E}^{\mathrm{cb}}_{\mu}$ is a majorising e-class that guarantees nothing is lost by relying on the coin-betting approach.
 This point is subtle but essential: establishing the optimality of coin-betting requires studying the properties of a set of e-variables (or equivalently, wealth-processes) rather than focusing on single e-variables or wealth-processes individually.
@@
 Then, ¯$\bar{\mathcal{E}}_{\mathcal{H}^{\prime}_{\mu}}=\{1\}$, while the optimal e-class exists and consists of all functions in the form $E_{\beta}:x\mapsto 1+\beta(x-\mu)$, with $\beta\in[0,\mu^{-1}]$.
-To simplify the exposition, we have focused on random variables valued in $[0,1]$.
-However, all the results presented can be extended to any bounded closed interval.
-More broadly, the author anticipates that similar conclusions hold for more general settings, such as estimating the mean of bounded vector-valued random variables.
-Addressing this case requires more advanced analytic tools and lies beyond the scope of this letter.
-As a final remark, this paper does not aim to delve into the design of effective coin-betting strategies for obtaining tight confidence sequences.
+Although one might argue that for each of the maximal e-variable there is at least one measure that bring the expectation to one (Grünwald, 2024 calls such property sharpness), sharpness alone is not enough to ensure maximality ($\mathbf{1}_{\mu}$ is a sharp e-variable for $\mathcal{H}_{\mu}$, but it is clearly not maximal).
+Two papers (Clerico, 2024; Larsson et al., 2025) have appeared after the first preprint of this work, both extending the results on characterising single-round e-variables for hypotheses defined via linear constraints, a framework that includes the testing for the mean as a special case.
+Their analyses, however, rely on considerably more abstract and technically advanced proof techniques.
+Conversely, a key strength of the present paper is to provide a simple and neat argument (illustrated in Figure 1) that works well in the simple setting considered.
+Interestingly, this same argument is also at the base of the proofs of Theorem 2 and Theorem 3.
+We also remark that a similar strategy was adopted in the proof of Lemma 2 in Wang (2025).
+That said, for the multi-round setting, the approach does give rise to non-trivial measurability issues when $\mathcal{X}$ is uncountable.
+These difficulties already emerge in the two-round case: although one can show that every e-variable must be dominated by a product of the form $(1+\lambda_{1}x_{1})(1+\lambda_{2}(x_{1})x_{2})$, this is constructed in a way that gives no guarantee on the measurability of $\lambda_{2}$.
+Whether this issue can be resolved remains an open question.
+Nonetheless, even in the simplified setting we adopt by requiring $\mathcal{X}$ to have finite cardinality [^11] A countable $\mathcal{X}$ would also suffice to solve the measurability issue, though the final step of the proof of Theorem 3 would require more delicate handling., this work is, to the best of the author’s knowledge, the first to fully characterise all e-processes for the natural problem of testing a fixed conditional mean.
+To conclude, let us remark that this paper does not aim to delve into the design of effective coin-betting strategies for obtaining tight confidence sequences.
 For this, we refer the interested reader to the thorough analysis and discussion by Orabona and Jun (2023) and Waudby-Smith and Ramdas (2023).
+Finally, although we have focused on random variables taking values in $[0,1]$ for clarity of exposition, all the results presented readily extend to any bounded closed interval.
 Acknowledgements
-The author would like to thank Gergely Neu, Peter Daniel Grünwald, Nishant Ajay Mehta, Hamish Edward Flynn, and Claudia Maria Chanu, for the insightful discussions that inspired this work.
-The author acknowledges the use of GPT-4o during the redaction of this letter, in order to polish the presentation.
+The author would like to thank Gergely Neu, Peter D.
+Grünwald, Nishant A.
+Mehta, Hamish E.
+Flynn, and Claudia M.
+Chanu, for the insightful discussions that inspired this work.
+I would also like to thank Nick W.
+Koning and Aditya Ramdas for their constructive feedback, which helped improve this work following its initial appearance, and Wouter M.
+Koolen for helpful discussions on e-processes.
+The author acknowledges the use of GPT-4o during the redaction of this paper, in order to polish the presentation.
 All AI-generated text was subsequently reviewed and edited by the author, who takes full responsibility for the content of this manuscript.
 This project was funded by the European Research Council (ERC), under the European Union’s Horizon 2020 research and innovation programme (grant agreement 950180).
+Appendix
+Proof of Theorem 2
+Proof.
+We start by showing that the class of functions considered is an e-class, and that all its elements are maximal e-variables.
+So, fix any $\lambda^{T}_{\mu}\in\Lambda^{T}$.
+As all the $\lambda_{t}$ are valued in $I_{\mu}$, it is clear that $E_{\lambda^{T}}$ is non-negative.
+Moreover, for any $P\in\mathcal{H}_{\mu}^{T}$ it is easily checked that $\mathbb{E}_{P}[E_{\lambda^{T}}]=1$, which ensures that $E_{\lambda^{T}}$ is an e-variable.
+For the maximality we can use the same argument as in the proof of Theorem 1.
+Let $\hat{E}\in\mathcal{E}^{T}_{\mu}$ be such that $\hat{E}\succeq E_{\lambda^{T}}$.
+Fix any $x^{T}\in\mathcal{X}^{T}$.
+By Lemma 4, for each $t$ we can find a $P_{t}\in\mathcal{H}^{1}_{\mu}$ such that $P_{t}(\{x_{t}\})>0$.
+Let $P=P_{1}\otimes\dots\otimes P_{T}$.
+Then we have that $P(\{x^{T}\})>0$.
+In particular, we have that $0\leq P(\{x^{T}\})(\hat{E}(x^{T})-E_{\lambda^{T}}(x^{T}))\leq\mathbb{E}_{P}[\hat{E}-E_{\lambda^{T}}]=\mathbb{E}_{P}[\hat{E}]-1\leq 0$.
+Since $P(\{x^{T}\})>0$, $\hat{E}(x^{T})=E_{\lambda^{T}}(x^{T})$, and we conclude that $E_{\lambda^{T}}$ is maximal.
+To show that the e-class that we consider is majorising we proceed by induction on $T$.
+From Theorem 1 we know that the statement is true for $T=1$.
+For $T\geq 2$.
+let us assume that the claim holds for $T-1$, and let us show that it is also true for $T$.
+Let $E$ be an e-variable for $\mathcal{H}^{T}_{\mu}$.
+First, we show that, for any $Q\in\mathcal{H}^{T-1}_{\mu}$ and any $x\in\mathcal{X}$, we have
+$\mathbb{E}_{Q}[E(x,X_{2},\dots,X_{T})]\leq F_{\mu}(x)\,,$
+with $F_{\mu}$ as in Lemma 4.
+Fix $Q$ and $x\geq\mu$ in $\mathcal{X}$.
+Define $V_{x}=P_{x}\otimes Q$, where $P_{x}=\frac{\mu}{x}\delta_{x}+(1-\frac{\mu}{x})\delta_{0}$.
+Clearly, $V_{x}\in\mathcal{H}_{\mu}^{T}$.
+We have that $1\geq\mathbb{E}_{V_{x}}[E]=\tfrac{\mu}{x}\mathbb{E}_{Q}[E(x,X_{2},\dots,X_{T})]+\big{(}1-\tfrac{\mu}{x}\big{)}\mathbb{E}_{Q}[E(0,X_{2},\dots,X_{T})]$, and so $\mathbb{E}_{Q}[E(x,X_{2},\dots,X_{T})]\leq x/\mu=F_{\mu}(x)$.
+The case for $x\leq\mu$ is essentially analogous.
+In particular, the following two sets are non-empty:
+$B_{0}=\big{\{}\beta\in I_{\mu}\,:\,\forall x\in\mathcal{X}\cap[0,\mu)\,,\,\forall Q\in\mathcal{H}_{\mu}^{T-1}\,,\,\mathbb{E}_{Q}[E(x,X_{2},\dots,X_{T})]\leq 1+\beta(x-\mu)\big{\}}\,;$
+$B_{1}=\big{\{}\beta\in I_{\mu}\,:\,\forall x\in\mathcal{X}\cap(\mu,1]\,,\,\forall Q\in\mathcal{H}_{\mu}^{T-1}\,,\,\mathbb{E}_{Q}[E(x,X_{2},\dots,X_{T})]\leq 1+\beta(x-\mu)\big{\}}\,.$
+$B_{0}$ and $B_{1}$ are intersections of closed intervals (one per each allowed $x$ and $Q$), and as such they must be closed intervals.
+We can find $\beta_{0}$ and $\beta_{1}$ such that $B_{0}=[(\mu-1)^{-1},\beta_{0}]$ and $B_{1}=[\beta_{1},\mu^{-1}]$.
+We will show that $\beta_{0}\geq\beta_{1}$, and hence that $B_{0}\cap B_{1}\neq\varnothing$.
+The argument is essentially the same used in the proof of Theorem 1 (cf.
+also Figure 1).
+Assume that $\beta_{0}<\beta_{1}$, and let $\beta^{\star}\in(\beta_{0},\beta_{1})$.
+As $\beta^{\star}\notin B_{0}$, there is $Q_{0}\in\mathcal{H}^{T-1}_{\mu}$ and $u_{0}<\mu$ in $\mathcal{X}$ such that $\mathbb{E}_{Q_{0}}[E(u_{0},X_{2},\dots,X_{T})]>1+\beta^{\star}(u_{0}-\mu)$.
+Similarly, there must be $Q_{1}\in\mathcal{H}_{\mu}^{T-1}$ and $u_{1}>\mu$ in $\mathcal{X}$, such that $\mathbb{E}_{Q_{1}}[E(u_{1},X_{2},\dots,X_{T})]>1+\beta^{\star}(u_{1}-\mu)$, since $\beta^{\star}\notin B_{1}$.
+Let $V=\frac{u_{1}-\mu}{u_{1}-u_{0}}\delta_{u_{0}}\otimes Q_{0}+\frac{\mu-u_{0}}{u_{1}-u_{0}}\delta_{u_{1}}\otimes Q_{1}$.
+Then, $V\in\mathcal{H}^{T}_{\mu}$.
+Moreover,
+$\mathbb{E}_{V}[E]>\tfrac{u_{1}-\mu}{u_{1}-u_{0}}\big{(}1+\beta^{\star}(u_{0}-\mu)\big{)}+\tfrac{\mu-u_{0}}{u_{1}-u_{0}}\big{(}1+\beta^{\star}(u_{1}-\mu)\big{)}=1\,,$
+which is a contradiction since $E\in\mathcal{E}^{T}_{\mu}$.
+We have thus established that $B_{0}\cap B_{1}$ is non-empty, and in particular there is $\lambda_{1}\in I_{\mu}$ such that, for every $x\in\mathcal{X}$ and every $Q\in\mathcal{H}_{\mu}^{T-1}$,
+$\mathbb{E}_{Q}[E(x,X_{2},\dots,X_{T})]\leq 1+\lambda_{1}(x-\mu)\,.$
+Fix $y\in\mathcal{X}$ such that $1+\lambda_{1}(y-\mu)\neq 0$.
+Define $E_{y}:\mathcal{X}^{T-1}\to\mathbb{R}$ via $E_{y}(x^{2:T})=E(y,x^{2:T})/(1+\lambda_{1}(y-\mu))$, where we recall that $x^{2:T}$ is a shorthand for $(x_{2},\dots,x_{T})$.
+$E_{y}$ is non-negative, since by definition of $I_{\mu}$ we know that $1+\lambda(y-\mu)$ is non-negative for any $\lambda\in I_{\mu}$.
+From what we have already shown, $E_{y}$ is an e-variable for $\mathcal{H}^{T-1}_{\mu}$.
+Thanks to the inductive hypothesis, for each $t\in[2:T]$ there exist a function $\tilde{\lambda}_{t}^{y}:\mathcal{X}^{t-2}\to I_{\mu}$ such that $E_{y}(x^{2:T})\leq\prod_{t=2}^{T}(1+\tilde{\lambda}_{t}^{y}(x^{2:(t-1)}))(x_{t}-\mu))$, for any $x^{2:T}\in\mathcal{X}^{T-1}$.
+For all $t\in[2:T]$, and every $x^{t-1}\in\mathcal{X}^{t-1}$, define
+$\lambda_{t}(x^{t-1})=\begin{cases}0&\text{if $1+\lambda_{1}(x_{1}-\mu)=0$};\\ \big{(}1+\lambda_{1}(x_{1}-\mu)\big{)}\tilde{\lambda}_{t}^{x_{1}}(x^{2:(t-1)})&\text{otherwise.}\end{cases}$
+So far, we have proved that for any $x^{T}\in\mathcal{X}^{T}$, such that $1+\lambda_{1}(x_{1}-\mu)\neq 0$, we have that $E(x^{T})\leq E_{\lambda^{T}}(x^{T})$.
+We only need to check that this is still true when $01+\lambda_{1}(x_{1}-\mu)=$ 0.
+In such case, $E$ is $Q$-almost surely null for all $Q\in\mathcal{H}^{T-1}_{\mu}$, which implies that $E(x^{T})=0$, and so we conclude.
+∎
+Proof of Theorem 3
+Proof.
+We start by recalling and introducing some definitions.
+For $T\geq 1$, $\Lambda^{T}_{\mu}$ is the set whose elements are $T$-tuple $\lambda^{T}=(\lambda_{1},\dots,\lambda_{T})$, with $\lambda^{t}:\mathcal{X}^{t-1}\to I_{\mu}$.
+We endow $\Lambda^{T}_{\mu}$ with the product topology of point-wise convergence.
+Specifically, a sequence $(\lambda^{t})_{t\geq 0}\to\tilde{\lambda}^{t}$ if, and only if, $\lambda_{t}\to\tilde{\lambda}_{t}$ point-wise for all $t\leq T$.
+For convenience, we will henceforth refer to this topology as the $\Lambda$-topology.
+It is straightforward to check that under the $\Lambda$-topology $\Lambda_{\mu}^{T}$ is compact, since $I_{\mu}$ is bounded and $\mathcal{X}$ has finite cardinality.
+Our proof strategy will rely on an inductive reasoning.
+For this, it is convenient to introduce the notion of a $T$-e-process, for $T\geq 1$.
+This is a sequence $(E_{t})_{t\geq 0}$ of non-negative functions $E_{t}:\mathcal{X}^{t}\to[0,+\infty)$, such that $\mathbb{E}_{P}[E_{\tau}]\leq 1$ for any $\tau\in\mathcal{T}_{T}$, and any $P\in\mathcal{H}_{\mu}^{\infty}$.
+We remark that an e-process is a $T$-e-process for all $T\geq 1$.
+We are now ready to start the proof.
+First, we will show that $\mathcal{E}^{\mathrm{cb},\infty}_{\mu}$ is a majorising e-process class.
+For any $T\geq 1$ and any $T$-e-process $E$, let $\tilde{\Lambda}^{T}_{E,\mu}\subseteq\Lambda^{T}_{\mu}$ denote the set of the $\lambda^{T}$ such that $E_{t}\preceq E_{\lambda^{t}}$, for all $t\in[1:T]$.
+We note that this is a closed and compact set under the $\Lambda$-topology that we introduced earlier.
+We now show that for every $T\geq 1$ and every $T$-e-process $E$, the set $\tilde{\Lambda}^{T}_{E,\mu}$ is non-empty.
+We prove this claim by induction.
+Since $E_{1}$ must be an e-variable for $\mathcal{H}_{\mu}^{1}$, as a consequence of Theorem 1, $\tilde{\Lambda}^{1}_{E,\mu}$ is non-empty.
+Now, fix $T\geq 2$ and a $T$-e-process $E$, and assume that, for any $(T-1)$-e-process $E^{\prime}$, $\tilde{\Lambda}^{T-1}_{E^{\prime},\mu}$ is non-empty.
+We will show that this implies that $\tilde{\Lambda}_{E,\mu}^{T}$ is non-empty.
+Fix an arbitrary $x\in\mathcal{X}$ and define the process $E^{x}=(E^{x}_{t})_{t\geq 1}$ such that $E_{t}^{x}(x^{t})=E_{t+1}(x,x_{1},\dots,x_{t})$, for any $x^{t}\in\mathcal{X}^{t}$.
+Define the sets
+$B_{0}=\big{\{}\beta\in I_{\mu}\,:\,\forall x\in\mathcal{X}\cap[0,\mu),\forall Q\in\mathcal{H}^{\infty}_{\mu},\forall\tau\in\mathcal{T}_{T-1},\,\mathbb{E}_{Q}[E^{x}_{\tau}]\leq 1+\beta(x-\mu)\big{\}}\,;$
+$B_{1}=\big{\{}\beta\in I_{\mu}\,:\,\forall x\in\mathcal{X}\cap(\mu,1],\forall Q\in\mathcal{H}^{\infty}_{\mu},\forall\tau\in\mathcal{T}_{T-1},\,\mathbb{E}_{Q}[E^{x}_{\tau}]\leq 1+\beta(x-\mu)\big{\}}\,.$
+We claim that $B_{0}$ and $B_{1}$ are non-empty.
+Indeed, we now show that $\mu^{-1}\in B_{1}$.
+Analogously one can prove that $(\mu-1)^{-1}\in B_{0}$.
+The argument here follows closely the one depicted in Figure 1.
+Fix $x\in\mathcal{X}\cap[0,\mu)$, $Q\in\mathcal{H}_{\mu}^{\infty}$, and $\tau\in\mathcal{T}_{T-1}$.
+Define $\tau_{x}:\mathcal{X}^{\infty}\to\mathbb{N}$ as $\tau_{x}(x^{\infty})=1+\tau(x^{2:\infty})$, if $x_{1}=x$, otherwise $\tau_{x}(x^{\infty})=1$.
+It is easily checked that $\tau\in\mathcal{T}_{T}$.
+Also, we let $P_{x}=\frac{\mu}{x}\delta_{x}+(1-\frac{\mu}{x}\delta_{0})$, and we define $V_{x}=P_{x}\otimes Q$ [^12] Note that both $V_{x}$ and $Q$ are measures on $\mathcal{P}_{\mathcal{X}^{\infty}}$.
+Here we mean that $V_{x}(\{x^{\infty}\})=P_{x}(\{x_{1}\})Q(\{x^{2:\infty}\})$ for all $x^{\infty}\in\mathcal{X}^{\infty}$.
+Clearly, $V_{x}\in\mathcal{H}_{\mu}^{\infty}$.
+Since $E$ is a $T$-e-process, we have
+$1\geq\mathbb{E}_{V_{x}}[E_{\tau_{x}}]=(1-\tfrac{\mu}{x})E_{1}(0)+\tfrac{\mu}{x}\mathbb{E}_{Q}[E^{x}_{\tau}]\geq\tfrac{\mu}{x}\mathbb{E}_{Q}[E^{x}_{\tau}]\,.$
+So, $E_{Q}[E^{x}_{\tau}]\leq\frac{x}{\mu}=1+\mu^{-1}(x-\mu)$, and so $\mu^{-1}\in B_{1}$.
+Since both $B_{0}$ and $B_{1}$ can be written as intersections of non-empty closed intervals, we can find $\beta_{0}$ and $\beta_{1}$ such that $B_{0}=[(\mu-1)^{-1},\beta_{0}]$ and $B_{1}=[\beta_{1},\mu^{-1}]$.
+As usual, we want to show that $B_{0}\cap B_{1}$ is non-empty, or equivalently that $\beta_{1}\leq\beta_{0}$.
+Assume that this was not the case and there is $\beta^{\star}\in(\beta_{0},\beta_{1})$.
+Since $\beta^{\star}\notin B_{0}$, there must be $u_{0}<\mu$ in $\mathcal{X}$, $Q_{0}\in\mathcal{H}_{\mu}^{\infty}$, and $\tau_{0}\in\mathcal{T}_{T-1}$ such that $\mathbb{E}_{Q_{0}}[E^{u_{0}}_{\tau_{0}}]>1+\beta^{\star}(u_{0}-\mu)$.
+Similarly, as $\beta^{\star}\notin B_{1}$, one can find $u_{1}>\mu$ in $\mathcal{X}$, $Q_{1}\in\mathcal{H}_{\mu}^{\infty}$, and $\tau_{1}\in\mathcal{T}_{T-1}$, such that $\mathbb{E}_{Q_{1}}[E^{u_{1}}_{\tau_{1}}]>1+\beta^{\star}(u_{1}-\mu)$.
+Define $\tau:\mathcal{X}^{\infty}\to\mathcal{N}$ via $\tau(x^{\infty})=1+\tau_{0}(x^{2:\infty})$, if $x_{1}<\mu$, $\tau(x^{\infty})=1+\tau_{1}(x^{2:\infty})$ if $x_{1}>\mu$, and $\tau(x^{\infty})=1$ if $x_{1}=\mu$.
+Also, let $V=\frac{u_{1}-\mu}{u_{1}-u_{0}}\delta_{u_{0}}\otimes Q_{0}+\frac{\mu-u_{0}}{u_{1}-u_{0}}\delta_{u_{1}}\otimes Q_{1}$.
+Then, we have that
+$\mathbb{E}_{V}[E_{\tau}]=\frac{u_{1}-\mu}{u_{1}-u_{0}}\mathbb{E}_{Q_{0}}[E^{u_{0}}_{\tau_{1}}(X_{2},\dots,X_{T})]+\frac{\mu-u_{0}}{u_{1}-u_{0}}\mathbb{E}_{Q_{1}}[E^{u_{1}}_{\tau_{1}}(X_{2},\dots,X_{T})]>1\,.$
+However, this is a contradiction since $\tau\in\mathcal{T}_{T}$, $V\in\mathcal{H}_{\mu}^{\infty}$, and $E$ is a $T$-e-process.
+We can thus conclude that $B_{0}\cap B_{1}\neq\varnothing$ and, in particular, there must be $\lambda_{1}\in I_{\mu}$ such that, for every $x\in\mathcal{X}$, every $Q\in\mathcal{H}_{\mu}^{\infty}$, and every $\tau\in\mathcal{T}_{T-1}$, we have
+$\mathbb{E}_{Q}[E^{x}_{\tau}]\leq 1+\lambda_{1}(x-\mu)\,.$
+Now, for any arbitrary $x\in\mathcal{X}$ such that $1+\lambda_{1}(x-\mu)\neq 0$. Define the process $\tilde{E}^{(x)}=(\tilde{E}^{(x)}_{t})_{t\geq 0}$ via
+$\tilde{E}^{(x)}_{t}(x^{t})=\frac{E_{t+1}(x,x_{1},\dots,x_{t})}{1+\lambda_{1}(x-\mu)}\,.$
+Let us show that $\tilde{E}^{(x)}$ is a $(T-1)$-e-process.
+It is clearly non-negative, as $\lambda_{1}\in I_{\mu}$ implies $1+\lambda_{1}(x-\mu)\geq 0$.
+Fix $Q\in\mathcal{H}_{\mu}^{\infty}$ and $\tau\in\mathcal{T}_{T-1}$.
+From what we have shown, we have
+$\mathbb{E}_{Q}[\tilde{E}^{(x)}_{\tau}]=\frac{\mathbb{E}_{Q}[E^{x}_{\tau}]}{1+\lambda_{1}(x-\mu)}\leq 1\,.$
+In particular, the inductive argument yields the existence of $\tilde{\lambda}_{(x)}^{T-1}\in\Lambda^{T-1}_{\mu}$, such that for all $t\in[1:T-1]$ we have $\tilde{E}^{(x)}\preceq E_{\tilde{\lambda}^{t}_{(x)}}$.
+Let $\lambda^{T}\in\Lambda^{T}_{\mu}$ be defined via $\lambda_{t}(x^{T})=\big{(}1+\lambda_{1}(x_{1}-\mu)\big{)}\tilde{\lambda}_{(x_{1}),t}(x_{2},\dots,x_{T})$, where we set $\tilde{\lambda}_{(x),t}$ identically equal to zero whenever $01+\lambda_{1}(x-\mu)=$ 0.
+We will now show that $\lambda^{T}\in\tilde{\Lambda}_{E,\mu}^{T}$.
+Fix any $x^{T}\in\mathcal{X}^{T}$.
+If $1+\lambda_{1}(x_{1}-\mu)\neq 0$, then for any $t\in[1:T]$ we have that
+$E_{t}(x^{t})=\big{(}1+\lambda_{1}(x_{t}-\mu)\big{)}E^{(x_{1})}_{t-1}(x_{2},\dots,x_{t})\leq\prod_{s=1}^{t}\big{(}1+\lambda_{s}(x^{s-1})(x_{s}-\mu)\big{)}=E_{\lambda^{t}}(x^{t})\,.$
+If $01+\lambda_{1}(x_{1}-\mu)=$ 0, then, for every $t\in[1:T-1]$, $\mathbb{E}_{Q}[E_{t}^{x_{1}}]=0$ for every $Q\in\mathcal{H}^{\infty}$, which implies that $E_{t}^{x_{1}}$ is identically null.
+In particular, $E_{t}(x^{t})=0=E_{\lambda^{t}}(x^{t})$.
+This shows that $\lambda^{T}\in\tilde{\Lambda}_{E,\mu}^{T}$, and so $\tilde{\Lambda}_{E,\mu}^{T}$ is non-empty.
+What we have shown so far allows to say that given an e-process $E$, for every $T\geq 1$ we can find a $\lambda^{T}\in\Lambda^{T}$ such that $E_{t}\leq E_{\lambda^{t}}$ for every $t\in[1:T]$.
+However, this does not tell us yet that there is a $\lambda^{\infty}\in\Lambda^{\infty}_{\mu}$ such that the same property holds.
+To show this, we leverage the compactness of $\tilde{\Lambda}^{T}_{E,\mu}$ under the $\Lambda$-topology.
+For $T^{\prime}\geq T$, let $\tilde{\Lambda}^{T|T^{\prime}}_{E,\mu}$ denote the set of $\lambda^{T}\in\Lambda^{T}$ such that there is $\tilde{\lambda}^{T^{\prime}}\in\tilde{\Lambda}_{E,\mu}^{T^{\prime}}$ satisfying, for all $t\in[1:T]$, $\lambda_{t}=\tilde{\lambda}_{t}$.
+It is easily checked that $\tilde{\Lambda}^{T|T^{\prime}}_{E,\mu}$ is compact in the $\Lambda$-topology.
+Moreover, if $T^{\prime\prime}\geq T^{\prime}$, then $\tilde{\Lambda}^{T|T^{\prime\prime}}_{E,\mu}\subseteq\tilde{\Lambda}^{T|T^{\prime}}_{E,\mu}$.
+Since the intersection of countably many non-empty compact sets is non-empty, we can define $\tilde{\Lambda}^{1|\infty}_{E,\mu}=\bigcap_{T\geq 1}\tilde{\Lambda}^{1|T}_{E,\mu}$, which is a non-empty set.
+Now, fix $\lambda_{1}\in\tilde{\Lambda}^{1|\infty}_{E,\mu}$.
+We have that $E_{1}\preceq E_{\lambda_{1}}$.
+By construction $\lambda_{1}\in\tilde{\Lambda}^{1|2}_{E,\mu}$, and so we can find a $\lambda_{2}$ such that $(\lambda_{1},\lambda_{2})\in\tilde{\Lambda}^{2}_{E,\mu}$.
+We can go on like this indefinitely.
+In particular, we can construct a sequence $\lambda^{\infty}$ such that $\lambda^{T}\in\tilde{\Lambda}^{T}_{E,\mu}$ for all $T\geq 1$.
+This concludes the proof that $\mathcal{E}^{\mathrm{cb},\infty}_{\mu}$ is a majorising e-process class.
+We are left with checking that every e-process in $\mathcal{E}^{\mathrm{cb},\infty}_{\mu}$ is maximal.
+Fix $\lambda^{\infty}\in\Lambda^{\infty}_{\mu}$, and let $E$ be an e-process such that $E\succeq E^{\lambda^{\infty}}$.
+Clearly, we must have $E_{0}=1=E_{0}^{\lambda^{\infty}}$.
+Now, fix $T\geq 1$.
+We have that $E_{T}\succeq E_{\lambda^{T}}$.
+Since $E$ and $E^{\lambda^{\infty}}$ are e-processes, choosing the stopping time that is identically equal to $T$ shows that $E_{T}$ and $E_{\lambda^{T}}$ are $T$-round e-variables.
+In particular, by Theorem 2, $E_{\lambda^{T}}$ is maximal, and so $E_{T}\succeq E_{\lambda^{T}}$ implies that $E_{T}$ and $E_{\lambda^{T}}$ coincide.
+Since this is true for all $T\geq 1$, we conclude.
+∎
+Characterising the e-variables for ^$\hat{\mathcal{H}}_{1/2}^{2}$ with $\mathcal{X}=\{0,1/2,1\}$
+Let $\mathcal{X}=\{0,1/2,1\}$, and $\mu=1/2$.
+Let $Q\in\mathcal{H}_{1/2}^{1}$ and let $q=Q(\{1/2\})$.
+The mean constraints reads $Q(\{1\})+q/2=1/2$, which implies that $Q(\{1\})=(1-q)/2$.
+The normalisation yields $Q(\{0\})=1-q-(1-q)/2=(1-q)/2$.
+This shows that there is a one-to-one correspondence between $[0,1]$ and $\mathcal{H}_{1/2}^{1}$.
+Now, since by definition there is a one-to-one correspondence between $\mathcal{H}_{1/2}^{1}$ and ^$\hat{\mathcal{H}}_{1/2}^{2}$, we obtain that we can parametrise ^$\hat{\mathcal{H}}_{1/2}^{2}$ by $[0,1]$.
+For $q\in[0,1]$, we hence let $P_{q}$ be the (unique) element of $\mathcal{H}_{1/2}^{2}$ that gives mass $q^{2}$ to $(1/2,1/2)$.
+Let $E:\mathcal{X}^{2}\to[0,+\infty)$ be a non-negative function. We have that
+$\mathbb{E}_{P_{q}}[E]=q^{2}\xi_{0}+2q(1-q)\xi_{1}+(1-q)^{2}\xi_{2}=(\xi_{0}+\xi_{2}-2\xi_{1})q^{2}-2(\xi_{0}-\xi_{1})q+\xi_{2}\,,$
+where $\xi_{0}=E(1/2,1/2)$, $\xi_{1}=(E(1,1/2)+E(1/2,1)+E(1/2,0)+E(0,1/2))/4$, and $\xi_{2}=(E(1,1)+E(1,0)+E(0,1)+E(0,0))/4$. Now, $E$ is in ^$\hat{\mathcal{E}}_{1/2}^{2}$ if, and only if,
+$\max_{q\in[0,1]}\big{(}(\xi_{0}+\xi_{2}-2\xi_{1})q^{2}-2(\xi_{0}-\xi_{1})q+\xi_{2}\big{)}\leq 1\,.$
+In particular, checking for $q=0$ and $q=1$ implies that $\xi_{0}\leq 1$ and $\xi_{2}\leq 1$.
+If both $\xi_{0}$ and $\xi_{2}$ are equal to $1$, then the constraint on $\xi_{1}$ becomes $\max_{q\in[0,1]}(1-\xi_{1})(q^{2}-q)\leq 0$, which implies $\xi_{1}\leq 1$.
+We are left to check if when one among $\xi_{0}$ and $\xi_{2}$ is strictly less then $1$ a $\xi_{1}$ can be larger than one.
+In such case, we note that we have $\xi_{2}-\xi_{1}<0$ and $\xi_{0}-\xi_{1}<0$.
+This implies that the parabola is concave and achieve its maximum at $q^{\star}=(\xi_{0}-\xi_{1})/(\xi_{0}-\xi_{1}+\xi_{2}-\xi_{1})\in(0,1)$.
+The maximum is equal to $\xi_{2}(\xi_{1}-\xi_{0})^{2}/(2\xi_{1}-\xi_{0}-\xi_{2})$.
+Asking that this quantity is less than one reduces to the constraint $\xi_{1}\leq 1+\sqrt{(1-\xi_{0})(1-\xi_{2})}$.

### v4

@@
 Abstract
-Confidence sequences are sequences of confidence sets that adapt to incoming data while maintaining validity.
-Recent advances have introduced an algorithmic formulation for constructing some of the tightest confidence sequences for the mean of bounded real random variables.
-These approaches use a coin-betting framework, where a player sequentially bets on differences between potential mean values and observed data.
-This work discusses the optimality of such coin-betting formulation among algorithmic frameworks building on e-variables methods to test and estimate the mean of bounded random variables.
+We consider the problem of testing the mean of a bounded real random variable.
+We introduce a notion of optimal classes for e-variables and e-processes, and establish the optimality of the coin-betting formulation among e-variable-based algorithmic frameworks for testing and estimating the (conditional) mean.
+As a consequence, we provide a direct and explicit characterisation of all valid e-variables and e-processes for this testing problem.
+In the language of classical statistical decision theory, we fully describe the set of all admissible e-variables and e-processes, and identify the corresponding minimal complete class.
+keywords:
+sequential hypothesis testing , e-variables , e-processes , mean estimation , admissibility.
+journal: International Journal of Approximate Reasoning
 1 Introduction
@@
 When the data are observed sequentially, one might want to update this set as new data-points become available.
-However, such procedure can compromise the validity of the statistical guarantee, if this was designed for a fixed sample size.
+However, such procedure may compromise the validity of the statistical guarantee, if this was designed for a fixed sample size.
 To address this issue, Darling and Robbins (1967) introduced the concept of confidence sequence, a data-adaptive sequence of confidence sets whose intersection contains the desired mean with high probability.
-Orabona and Jun (2023) and Waudby-Smith and Ramdas (2023) have recently suggested algorithmic approaches that yield some of the tightest confidence sequences for the mean of a bounded real random variable.
+Orabona and Jun (2023) and Waudby-Smith and Ramdas (2023) have recently explored algorithmic approaches that yield some of the tightest confidence sequences for the mean of a bounded real random variable.
 Both papers propose setting up a series of sequential coin-betting games, one per each mean candidate value $\mu$, where a player sequentially bets on the difference between $\mu$ and the upcoming observation.
-If $\mu$ matches the true mean, the game is fair, making substantial gains unlikely.
-A confidence sequence is then obtained by excluding those values $\mu$ that allowed the player to accumulate significant wealth.
+If $\mu$ matches the true mean, the game is fair, and substantial gains unlikely.
+A confidence sequence is obtained by excluding those values $\mu$ that allowed the player to accumulate significant wealth.
 This coin-betting approach to mean estimation is a particular instance of a broader algorithmic framework for constructing confidence sequences through sequential hypothesis testing, which can be framed in terms of betting games where at each round the player has to select an e-variable (Shafer, 2021; Ramdas et al., 2022a, 2023).
-E-variables, non-negative random variables whose expectation is bounded by one under the tested hypothesis (Grünwald et al., 2024), have recently emerged as a powerful and increasingly popular tool for anytime-valid hypothesis testing.
+E-variables, non-negative random variables whose expectation is bounded by 111 under the tested hypothesis (Grünwald et al., 2024), have recently emerged as a powerful and increasingly popular tool for anytime-valid hypothesis testing.
 By serving as building blocks for constructing non-negative super-martingales, which can be seen as representing the wealth of a player in a betting game, e-variables naturally lend themselves to game-theoretic interpretations (Shafer and Vovk, 2019; Ramdas and Wang, 2024).
-The first part of this work focuses on round-wise testing-by-betting, a scenario where the player iteratively picks a single-round e-variable, whose choice can depend on the past observations.
-Our main goal is to illustrate that, when estimating the mean of a bounded real random variable, no instantiation of such round-wise testing-by-betting game can yield strictly better results than the coin-betting approach, in terms of tightness of the resulting confidence sequence.
-In a sense to be clarified later, coin-betting is optimal, as it represents the “simplest” formulation among those that cannot be strictly outperformed by any other such betting approach.
+The main goal of this work is to illustrate and formalise that, when sequentially testing and estimating the (conditional) mean of a bounded real random variable, no e-variable procedure yields strictly better guarantees than the coin-betting approach.
+In a sense to be clarified later, coin-betting is optimal, as it represents the “simplest” formulation among those that cannot be strictly performed by any other such testing-by-betting approach.
 One main novelty of this work is the introduction of a notion of optimality at the level of sets of e-variables.
-This perspective differs from much of the existing literature, which is primarily focused on the optimality of an individual e-variable or wealth process (e.g., log-optimality in Koolen and Grünwald 2022; Grünwald et al.
+This perspective differs from much of the existing literature, primarily focused on the optimality of an individual e-variable, or wealth process, in the betting game (e.g., log-optimality in Koolen and Grünwald 2022; Grünwald et al.
 2024; Larsson et al.
@@
 2022a).
-An alternative way to frame the first part of this work is as a characterisation of the family of single-round e-variables for testing the mean of a bounded random variable.
-In a similar spirit, since the appearance of the first version of this manuscript, two works have derived more general results for testing hypotheses defined by linear constraints (Clerico, 2024; Larsson et al., 2025), of which the present setting is a particular case.
-However, the approach we present here is direct and tailored to the simple setting considered, making it valuable for building intuition, while the more general results are significantly more abstract.
-We remark that testing via sequential betting with single-round e-variables is not the most general form of sequential testing with e-variables, which typically relies on multi-round e-variables and e-processes (Shafer, 2021; Koolen and Grünwald, 2022; Ramdas et al., 2022b; Ramdas and Wang, 2024).
-These allow for testing hypotheses over the entire data sequence, including assumptions about the dependence structure (e.g., i.i.d.
+We remark that the perspective adopted in this work can be seen as an adaptation, to the setting of e-variables, of the classical statistical problem of identifying a minimal complete class of tests (see Lehmann and Romano, 2022).
+Further discussion of this connection is deferred to Section 7.
+The first part of this work focuses on round-wise testing-by-betting, a scenario where the player iteratively picks a single-round e-variable, whose choice may depend on the past observations.
+This procedure naturally applies to the setting where the observations are known to be independently drawn from a fixed probability distribution, whose mean has to be estimated.
+However, we remark that testing via sequential betting with single-round e-variables is not the most general form of sequential testing with e-variables, which typically relies on multi-round e-variables and e-processes (Shafer, 2021; Koolen and Grünwald, 2022; Ramdas et al., 2022b; Ramdas and Wang, 2024).
+These tools allow for testing hypotheses over the entire data sequence, including assumptions about the dependence structure (e.g., i.i.d.
 or fixed conditional mean).
-In such cases, restricting the player to select a single-round e-variable at each step can be limiting (Koolen and Grünwald, 2022; Ramdas et al., 2022b).
+In such cases, restricting the player to select a single-round e-variable at each step may be highly limiting (Koolen and Grünwald, 2022; Ramdas et al., 2022b).
 The second part of this work considers this broader setting.
-For technical convenience, we only focus on variables with finite discrete support, though we conjecture our conclusions hold more generally.
-We establish that when the sequence has fixed conditional mean, coin betting remains optimal even among testing methods based on multi-round e-variables and e-processes.
+We establish that when the sequence has fixed conditional mean, coin-betting remains optimal even among testing methods based on multi-round e-variables and e-processes.
 However, we also show that this optimality result no longer holds under the more restrictive hypothesis of i.i.d.
 observations.
+It is worth noting that an alternative way to frame the main contribution of this work is as a concrete and direct characterisation of the family of e-variables and e-processes for testing the (conditional) mean of a real bounded random variable.
+More precisely, the e-variables and e-processes for these tests are exactly the non-negative measurable functions or processes that are majorised by a coin-betting e-variable or e-process.
+Since the coin-betting formulation provides a very explicit expression for these objects, our results directly yield a fully explicit description of the full set of e-variables and e-processes for the problem at hand.
+Following a first pre-print of this manuscript, general characterisations for the e-variables when testing hypotheses defined by linear constraints were established by Clerico (2024) and Larsson et al.
+(2025).
+These results directly imply ours for single-round e-variables (Theorem 1), which is also explicitly discussed by these works as an application.
+However, the approach we present here is direct and tailored to the simple setting considered, making it valuable for building intuition, while the more general results are significantly more abstract.
+Moreover, the two aforementioned works focus exclusively on single-round e-variables, whereas our contribution provides a complete and explicit characterisation of both e-variables and e-processes in the sequential conditional mean testing problem.
+For further discussion on these works, see Section 7.
 Notation
 We endow any Borel set $\mathcal{Z}\subseteq\mathbb{R}^{d}$ with the standard topology, and we denote as $\mathcal{P}_{\mathcal{Z}}$ the set of Borel probability measures on $\mathcal{Z}$.
-For $z\in\mathcal{Z}$, $\delta_{z}$ denotes the Dirac unit mass on $z$.
+For $z\in\mathcal{Z}$, $\delta_{z}$ is the Dirac unit mass on $z$.
 For $P\in\mathcal{P}_{\mathcal{Z}}$ and a Borel measurable function $f$ on $\mathcal{Z}$, $\mathbb{E}_{P}[f(Z)]$ (or more compactly $\mathbb{E}_{P}[f]$) denotes the expectation of $f$ under $Z\sim P$.
-Given a sub-sigma-field $\mathcal{G}$, $\mathbb{E}_{P}[Z|\mathcal{G}]$ denotes the conditional expectation.
-Typically, $\mathcal{G}$ will be the sigma-field generated by some random variable $X$.
-In such case, we will write $\mathbb{E}_{P}[Z|X]$.
-We will also consider measures on product spaces.
-For $T\geq 1$, we write $\mathcal{P}_{\mathcal{Z}^{T}}$ the set of Borel probability measure on $\mathcal{Z}^{T}$ (endowed with the product topology).
-We will used $\otimes$ to denote the direct product of measures.
-For instance, given $P$ and $Q$ in $\mathcal{P}_{\mathcal{Z}}$, $P\otimes Q$ will be the element of $\mathcal{P}_{Z^{2}}$ that encode the law of $(Z_{1},Z_{2})$, where $Z_{1}\sim P$ and $Z_{2}\sim Q$ are independent.
-For any two given integers $s$ and $t$, with $s\leq t$, $[s:t]$ denotes the set of integers between $s$ and $t$ (included).
-For an integer $T$, given a vector $(x_{1},\dots,x_{T})$, we often represent it compactly as $x^{T}$ (upper indices).
-At times, we will also use the notation $x^{t:T}$ (with $t\in[1:T]$), to denote the vector $(x_{t},x_{t+1},\dots,x_{T})$.
-Sequences are denoted as $(s_{t})_{t\geq T_{0}}$, with $t$ an integer index and $T_{0}$ its smallest value (typically 00 or $1$).
-Sometimes we will also use the notation $s^{T_{0}:\infty}$ to denote $(s_{t})_{t\geq T_{0}}$, or simply $s^{\infty}$ if $T_{0}$ is clear from the context.
-For high probability statements, $\mathbb{P}$ expresses probability with respect to all the randomness involved.
-For instance, if $(X_{t})_{t\geq 1}$ is a sequence of i.i.d.
-draws from $P\in\mathcal{P}_{\mathcal{X}}$, we write $\mathbb{P}(X_{t}\geq 1/2\,,\,\forall t\geq 1)$, with obvious meaning.
-2 Algorithmic mean estimation
-Fix a Borel set $\mathcal{X}\subseteq\mathbb{R}$.
-Let $(X_{t})_{t\geq 1}$ be a sequence of independent random variables drawn from an unknown fixed distribution $P^{\star}\in\mathcal{P}_{\mathcal{X}}$, whose mean $\mu^{\star}\in(0,1)$ has to be estimated.
-Let $\mathcal{F}=(\mathcal{F}_{t})_{t\geq 0}$ represent the natural filtration generated by this sequence, where $\mathcal{F}_{t}=\sigma(X_{1},\dots,X_{t})$ captures all information available up to time $t$.
-Fix a confidence level parameter $\delta\in(0,1)$.
-A confidence sequence $(S_{t})_{t\geq 1}$ is a sequence of random sets such that the sequence of events ($(\{\mu^{\star}\in S_{t}\})_{t\geq 1}$ is adapted with respect to the filtration $\mathcal{F}$ (i.e., for all $t\geq 1$, $\{\mu^{\star}\in S_{t}\}$ is $\mathcal{F}_{t}$-measurable) and satisfies the property
-$\mathbb{P}\big{(}\mu^{\star}\in S_{t}\,,\;\forall t\geq 1\big{)}\geq 1-\delta\,.$
-Intuitively, this means that $(S_{t})_{t\geq 1}$ provides a set of plausible values for $\mu^{\star}$ at each time step $t$, while ensuring that the true mean remains in these sets indefinitely with high probability.
+Given a sub-sigma-field $\mathcal{G}$, $\mathbb{E}_{P}[Z|\mathcal{G}]$ is the conditional expectation.
+We will be interested in the case of $\mathcal{G}$ being the sigma-field generated by some random variable $X$.
+In such case, we write $\mathbb{E}_{P}[Z|X]$.
+We also consider measures on product spaces.
+For $T\geq 1$, we write $\mathcal{P}_{\mathcal{Z}^{T}}$ for the set of Borel probability measures on $\mathcal{Z}^{T}$ (endowed with the product topology).
+We will use $\otimes$ to denote the direct product of measures.
+For instance, given $P$ and $Q$ in $\mathcal{P}_{\mathcal{Z}}$, $P\otimes Q$ will be the element of $\mathcal{P}_{\mathcal{Z}^{2}}$ that encodes the law of $(Z_{1},Z_{2})$, where $Z_{1}\sim P$ and $Z_{2}\sim Q$ are independent.
+[…]
+2 Algorithmic mean testing via single-round e-variables
+We start by presenting a framework for sequential hypothesis testing, formalised as a betting game.
+We then specialise to testing the mean of a bounded distribution, introducing the coin-betting approach.
+2.1 Sequential testing game
+Let $\mathcal{Z}$ be a non-empty Borel set in $\mathbb{R}^{d}$.
+A hypothesis on $\mathcal{Z}$ is a non-empty subset $\mathcal{H}$ of $\mathcal{P}_{\mathcal{Z}}$, and an e-variable (for $\mathcal{H}$) is a non-negative Borel measurable function $E:\mathcal{Z}\to[0,+\infty)$, such that
+$\mathbb{E}_{P}[E]\leq 1\,,\qquad\forall P\in\mathcal{H}\,.$
+We denote as $\mathcal{E}_{\mathcal{H}}$ the set of all the e-variables with respect to $\mathcal{H}$, and we call e-class any subset of $\mathcal{E}_{\mathcal{H}}$.
+We remark that $\mathcal{E}_{\mathcal{H}}$ is never empty, since the constant function 111 is always an e-variable, for any $\mathcal{H}$.
+Definition 1 (Testing-by-betting game)
+Fix a hypothesis $\mathcal{H}\subseteq\mathcal{P}_{\mathcal{Z}}$ and a non-empty e-class $\mathcal{E}\subseteq\mathcal{E}_{\mathcal{H}}$.
+An $\mathcal{E}$-restricted testing-by-betting game (on $\mathcal{H}$) is the following sequential procedure.
+Each round $t\geq 1$,
+  • 1.
+    the player picks [^1] We assume that $E_{t}$ is picked in a measurable fashion, namely the mapping $(z_{1},\dots z_{t})\mapsto E_{t}(z_{t})$ is Borel measurable.
+an e-variable $E_{t}\in\mathcal{E}$ based solely on the past observations $z_{1},\dots,z_{t-1}$;
+  • 2.
+    the player observes a new data-point $z_{t}\in\mathcal{Z}$;
+  • 3.
+    the player earns a reward $\log E_{t}(z_{t})$.
+If $\mathcal{E}=\mathcal{E}_{\mathcal{H}}$, we speak of unrestricted testing-by-betting game.
+The above game is an instance of e-variable testing, where one designs a test that rejects the hypothesis $\mathcal{H}$ whenever the total reward earned by the player gets excessively high.
+This procedure is justified by the fact that, if the data-points observed during the game were independently drawn from $P\in\mathcal{H}$, then the cumulative reward would be unlikely to grow very large.
+This is formalised by the following proposition.
+Proposition 1
+Let $\mathcal{H}\subseteq\mathcal{P}_{\mathcal{Z}}$ and consider a sequence $(Z_{t})_{t\geq 1}$ of independent draws from $P\in\mathcal{H}$.
+Fix $\delta\in(0,1)$ and $\mathcal{E}\subseteq\mathcal{E}_{\mathcal{H}}$.
+Consider an $\mathcal{E}$-restricted testing-by-betting game, where the observations are the sequence $(Z_{t})_{t\geq 1}$.
+Let $R_{n}=\sum_{t=1}^{n}\log E_{t}(Z_{t})$ represent the player’s cumulative reward at round $n$.
+Then,
+$\mathbb{P}\big{(}R_{n}\leq\log\tfrac{1}{\delta}\,,\;\forall n\geq 1\big{)}\geq 1-\delta\,.$
+Proof 1
+The result follows directly from Ville’s inequality, since $M_{n}=\prod_{t=1}^{n}E_{t}(Z_{t})$ defines a non-negative super-martingale with respect to the natural filtration of the process $(Z_{t})_{t\geq 1}$, with $M_{0}\equiv 1$$\qed$
+The cumulative reward earned by the player serves as a quantitative measure of evidence against the hypothesis $\mathcal{H}$.
+Given a sequence of independent observations known to be drawn from some $P\in\mathcal{P}(\mathcal{Z})$, Proposition 1 justifies the following sequential testing procedure: the null hypothesis “the data generating distribution $P$ is in $\mathcal{H}$” is rejected as soon as the player’s total reward exceeds the threshold $\log(1/\delta)$, for a chosen confidence level $\delta\in(0,1)$.
+In this setup, $\delta$ controls the Type I error rate, ensuring that the probability of wrongly rejecting a true null is at most $\delta$.
+Remarkably, Proposition 1 guarantees this control uniformly over time, allowing the statistician to freely decide when to stop the test.
+For more details on sequential testing by betting, we refer to Ramdas et al.
+(2023), or Chapter 7 of Ramdas and Wang (2024).
+To design a powerful test, we want the rewards to accumulate rapidly whenever the data provide evidence against the null.
+To this regard, the pool $\mathcal{E}$, from which the player can pick the e-variables, plays an important role: excluding useful functions may weaken the test, while including unnecessary ones (e.g., the constant 1/21/21 / 2) adds no value.
+A carefully tailored class can simplify strategy design while preserving statistical power.
+The goal of this paper is to identify the “best” e-class to use when testing for the mean of a bounded real random variable.
+2.2 Testing for the mean and coin-betting
+Fix a Borel set $\mathcal{X}\subseteq[0,1]$, containing 0 and 111.222The main results of this work are actually valid for any Borel set whereof [0,1][0,1][ 0 , 1 ] is the convex closure.
+The requirement that 0 and 111 belong to $\mathcal{X}$ slightly simplifies some proofs (e.g., Lemma 3), which use the fact that 0 and 111 are in $\mathcal{X}$.
+Let $(X_{t})_{t\geq 1}$ be a sequence of independent random variables drawn from an unknown fixed distribution $P^{\star}\in\mathcal{P}_{\mathcal{X}}$, with mean $\mu^{\star}\in(0,1)$.
+To test whether $\mu^{\star}$ equals a given value $\mu\in(0,1)$, we can define the corresponding null hypothesis:
+$\mathcal{H}_{\mu}=\{P\in\mathcal{P}_{\mathcal{X}}:\mathbb{E}_{P}[X]=\mu\}\,,$ (1)
+which is the set of all distributions on $\mathcal{X}$ having mean $\mu$.
+Rejecting $\mathcal{H}_{\mu}$ thus amounts to rejecting the claim that $\mu^{\star}=\mu$.
+Such null hypotheses are closely related to the problem of mean estimation, specifically, to constructing a sequence of intervals that, with high probability, contain the true mean (i.e., a confidence sequence).
+We will make this connection more precise in Section 5.
 We stress here that, for the time being, we assume that the sequence $(X_{t})_{t\geq 1}$ is i.i.d., and we shall not challenge this assumption, regardless of whether or not the observed data appear to exhibit such behaviour.
-In a way, this perspective aligns naturally with the goal of constructing confidence sequences, as simultaneously testing for the mean and the i.i.d.
-assumption might lead to rejecting everything and returning empty confidence sets because “the data do not look i.i.d.”.
-Under the i.i.d.
+In a way, this perspective aligns naturally with the goal of constructing confidence sequences that we will discuss in Section 5, as simultaneously testing for the mean and the i.i.d.
+assumption might lead to rejecting every point in $\mathcal{X}$ and returning empty confidence sets because “the data do not look i.i.d.
+enough”.
+As a matter of facts, under the i.i.d.
 model, mean estimation is well posed: we assume the existence of a fixed distribution $P^{\star}$ independently generating each observation, and our task is to estimate its mean.
-We remark that the independence assumption could be relaxed by only asking that each $X_{t}$ has fixed conditional mean $\mu^{\star}$ (to be estimated) given the past $\mathcal{F}_{t-1}$.
+We remark that the independence assumption could be relaxed by just asking that each $X_{t}$ has fixed conditional mean $\mu^{\star}$ (to be estimated) given the past $\mathcal{F}_{t-1}$.
 The arguments we present next for the i.i.d.
@@
 However, to keep the exposition clearer, we focus here solely on the independent case.
-The conditional setting will be addressed in the second part of this work, where we consider the more general scenario in which dependencies between observations are not assumed but are instead challenged by the statistical test.
-2.1 Coin-betting confidence sequence
-Recently, Orabona and Jun (2023) and Waudby-Smith and Ramdas (2023) obtained some of the tightest known confidence sequences for bounded random variables via an algorithmic approach.
-The idea behind both papers is to set up a sequential coin-betting game for each candidate value $\mu\in(0,1)$ for the mean, and to exclude from the confidence sets those values $\mu$ that allow the player to earn unlikely high rewards.
-Definition 1 (Coin-betting game).
-Fix $\mu\in(0,1)$ and let $I_{\mu}=[(\mu-1)^{-1},\mu^{-1}]$. Consider the following sequential procedure. At each round $t\geq 1$, a player
-  • picks [^2] We implicitly assume a measurable selection of $\beta_{t}$, namely the mapping $(x_{1},\dots x_{t-1})\mapsto\beta_{t}$ is Borel measurable.
+The conditional setting will be addressed in the second part of this work (Section 6), where we consider the more complex scenario in which the nature of the depedences between observations is not given or assumed, but is instead challenged via statistical testing.
+2.2.1 The coin-betting e-class
+Recently, Orabona and Jun (2023) and Waudby-Smith and Ramdas (2023) obtained some of the tightest known confidence sequences for the mean of bounded real random variables via an algorithmic approach based on sequential testing.
+The main idea behind both papers involves the following sequential testing game for $\mathcal{H}_{\mu}$, which can be thought as betting on the outcome of a “continuous” coin (see Orabona and Jun 2023 for a thorough discussion on the “coin-betting” interpretation).
+Definition 2 (Coin-betting game)
+Fix $\mu\in(0,1)$ and let $I_{\mu}=[(\mu-1)^{-1},\mu^{-1}]$.
+[^3] The definition of $I_{\mu}$ ensures that the game’s rewards are well defined, as logarithms of non-negative quantities.
+Consider the following sequential procedure.
+At each round $t\geq 1$, a player
+  • 1.
+    picks [^4] Again, we implicitly assume a measurable selection of $\beta_{t}$, namely the mapping $(x_{1},\dots x_{t-1})\mapsto\beta_{t}$ is Borel measurable.
 $\beta_{t}\in I_{\mu}$, based solely on the past observations $x_{1},\dots,x_{t-1}$;
+  • 2.
     observes a new data-point $x_{t}\in[0,1]$;
-  • gets a reward $\log\big{(}1+\beta_{t}(x_{t}-\mu)\big{)}$.
-Now, for each $\mu\in(0,1)$, consider a coin-betting game where at each round the player observes a data-point from the sequence $(X_{t})_{t\geq 1}$ drawn from $P^{\star}$.
-Let $R_{n}(\mu)=\sum_{t=1}^{n}\log(1+\beta_{t}(X_{t}-\mu))$ denote the cumulative reward at round $n$.
-Then, via standard martingale concentration arguments (which will become clear later) one can show that the sets $S_{n}=\big{\{}\mu\in(0,1)\,:\,R_{n}(\mu)\leq\log\tfrac{1}{\delta}\big{\}}$ define a valid confidence sequence for $\mu^{\star}$.
-2.2 Sequential testing game
-The coin-betting approach is a particular instance of constructing confidence sequences via a sequential hypothesis testing procedure based on e-variables.
-Here, we call hypothesis an arbitrary subset $\mathcal{H}$ of $\mathcal{P}_{[0,1]}$.
-An e-variable (for $\mathcal{H}$) is any non-negative Borel measurable function $E:[0,1]\to[0,+\infty)$, such that
-$\mathbb{E}_{P}[E]\leq 1\,,\qquad\forall P\in\mathcal{H}\,.$
-We denote as $\mathcal{E}_{\mathcal{H}}$ the set of all the e-variables with respect to $\mathcal{H}$, and we call e-class any subset of $\mathcal{E}_{\mathcal{H}}$.
-We remark that $\mathcal{E}_{\mathcal{H}}$ is never empty, since the constant function one is always an e-variable, independently of $\mathcal{H}$.
-Definition 2 (Testing-by-betting game).
-Fix a hypothesis $\mathcal{H}\subseteq\mathcal{P}_{[0,1]}$ and a non-empty e-class $\mathcal{E}\subseteq\mathcal{E}_{\mathcal{H}}$.
-We call $\mathcal{E}$-restricted testing-by-betting game (on $\mathcal{H}$) the following sequential procedure.
-At each round $t\geq 1$,
-  • the player picks [^3] Again, we assume that $E_{t}$ is picked in a measurable fashion, namely the mapping $(x_{1},\dots x_{t})\mapsto E_{t}(x_{t})$ is Borel measurable.
-an e-variable $E_{t}\in\mathcal{E}$ based solely on the past observations $x_{1},\dots,x_{t-1}$;
-  • the player observes a new data-point $x_{t}\in[0,1]$;
-  • the player earns a reward $\log E_{t}(x_{t})$.
-If $\mathcal{E}=\mathcal{E}_{\mathcal{H}}$, we speak of unrestricted testing-by-betting game.
-The above game is an instance of e-variable testing (see Ramdas and Wang, 2024, for a recent overview), where one designs a test that rejects the hypothesis $\mathcal{H}$ whenever the total reward earned by the player gets excessively high.
-This procedure is justified by the fact that, if the data-points observed during the game were independently drawn from $P\in\mathcal{H}$, then the cumulative reward would be unlikely to grow very large.
-Lemma 1.
-Let $\mathcal{H}\subseteq\mathcal{P}_{[0,1]}$ and consider a sequence $(X_{t}^{\prime})_{t\geq 1}$ of independent draws from some $P\in\mathcal{H}$.
-Fix $\delta\in(0,1)$ and $\mathcal{E}\subseteq\mathcal{E}_{\mathcal{H}}$.
-Consider an $\mathcal{E}$-restricted testing-by-betting game, where the observations are the sequence $(X_{t}^{\prime})_{t\geq 1}$.
-Let $R_{n}=\sum_{t=1}^{n}\log E_{t}(X_{t}^{\prime})$ represent the player’s cumulative reward at round $n$.
-We have
-$\mathbb{P}\big{(}R_{n}\leq\log\tfrac{1}{\delta}\,,\;\forall n\geq 1\big{)}\geq 1-\delta\,.$
-Proof.
-The result follows directly from Ville’s inequality, since $M_{n}=\prod_{t=1}^{n}E_{t}(X_{t}^{\prime})$ defines a non-negative super-martingale with respect to the natural filtration of the process $(X_{t}^{\prime})_{t\geq 1}$, with $M_{0}\equiv 1$.
-∎
-2.3 Mean estimation via sequential betting
-We can leverage the testing-by-betting game of Definition 2 to obtain a confidence sequence for the mean $\mu^{\star}\in(0,1)$ of $P^{\star}$. For $\mu\in(0,1)$, define the hypothesis
-$\mathcal{H}_{\mu}=\{P\in\mathcal{P}_{[0,1]}\,:\,\mathbb{E}_{P}[X]=\mu\}\,,$ (1)
-which contains all probability measures on $[0,1]$ with mean $\mu$.
-For each $\mu$, we fix an e-class $\mathcal{E}_{\mu}\subseteq\mathcal{E}_{\mathcal{H}_{\mu}}$ and we consider an $\mathcal{E}_{\mu}$-restricted testing-by-betting game on $\mathcal{H}_{\mu}$, where the player observes $(X_{t})_{t\geq 1}$, the sequence of draws from $P^{\star}$.
-For each one of these games, denote as $R_{n}(\mu)$ the player’s cumulative reward at round $n$.
-Proposition 1.
-The sequence $(S_{n})_{n\geq 1}$, defined as
-$S_{n}=\big{\{}\mu\in(0,1)\;:\;R_{n}(\mu)\leq\log\tfrac{1}{\delta}\big{\}}\,,$
-is a confidence sequence for the mean $\mu^{\star}$ of the data-generating probability measure $P^{\star}$.
-Proof.
-For each $n\geq 1$, we have that $\{\mu^{\star}\in S_{n}\}=\{R_{n}(\mu^{\star})\leq\log\frac{1}{\delta}\}$, which is a $\mathcal{F}_{n}$-measurable event. Moreover,
-$\mathbb{P}\big{(}\mu^{\star}\in S_{n}\,,\;\forall n\geq 1\big{)}=\mathbb{P}\big{(}R_{n}(\mu^{\star})\leq\log\tfrac{1}{\delta}\,,\;\forall n\geq 1\big{)}\geq 1-\delta$
-by Lemma 1, as the observations are drawn from $P^{\star}\in\mathcal{H}_{\mu^{\star}}$. ∎
-We remark that the approach to derive confidence sequences via the coin-betting game, proposed by Orabona and Jun (2023) and Waudby-Smith and Ramdas (2023), is a specific instance of the framework that we have just described.
-Indeed, letting $E_{t}:x\mapsto 1+\beta_{t}(x-\mu)$, it is straightforward to verify that $E_{t}$ is non-negative on $[0,1]$, due to the restriction $\beta_{t}\in I_{\mu}=[(\mu-1)^{-1},\mu^{-1}]$ in Definition 1.
+  • 3.
+    receives the reward $\log\big{(}1+\beta_{t}(x_{t}-\mu)\big{)}$.
+We remark that this coin-betting game is a specific instance of the testing-by-betting game that we have described earlier.
+Indeed, letting $E_{t}:x\mapsto 1+\beta_{t}(x-\mu)$, it is straightforward to verify that $E_{t}$ is non-negative on $\mathcal{X}$, due to the restriction $\beta_{t}\in I_{\mu}=[(\mu-1)^{-1},\mu^{-1}]$ in Definition 2.
 Moreover, for any $P\in\mathcal{H}_{\mu}$, we have $\mathbb{E}_{P}[E_{t}]=1$, which implies that $E_{t}$ is an e-variable for $\mathcal{H}_{\mu}$.
 Finally, the reward in the coin-betting game is precisely equal to $\log E_{t}(x_{t})$.
-Hence, for the hypothesis $\mathcal{H}_{\mu}$, the coin-betting game matches exactly the testing game of Definition 2, restricted to the coin-betting e-class
+Hence, for the hypothesis $\mathcal{H}_{\mu}$, the coin-betting game above matches exactly the testing game of Definition 1, restricted to the coin-betting e-class
 $\mathcal{E}^{\mathrm{cb}}_{\mu}=\big{\{}E_{\beta}:x\mapsto 1+\beta(x-\mu)\,,\;\;\beta\in I_{\mu}\big{\}}\,.$ (2)
-Ideally, one aims to construct tight confidence sets.
+2.2.2 A suboptimal choice: the Hoeffding e-class
+Another perhaps natural e-class to test $\mathcal{H}_{\mu}$ is
+$\mathcal{E}^{\mathrm{Hoeff}}_{\mu}=\{E^{\mathrm{H}}_{\alpha}:x\mapsto e^{\alpha(x-\mu)-\alpha^{2}/8}\,,\;\;\alpha\in\mathbb{R}\}\,,$
+which we will refer to as the Hoeffding e-class. The fact that this is an e-class follows immediately from the well known sub-Gaussian nature of bounded random variables (Hoeffding’s lemma).
+Let us now consider two players, Alice and Bob, playing two different testing-by-betting games for $\mathcal{H}_{\mu}$.
+Alice plays a $\mathcal{E}^{\mathrm{cb}}_{\mu}$-restricted game, while Bob a $\mathcal{E}^{\mathrm{Hoeff}}_{\mu}$-restricted game.
+We will now show that Alice’s game is strictly stronger, from a statistical perspective, than Bob’s one.
+More precisely, if Bob plays first, and picks $E^{\mathrm{H}}_{\alpha}$, there is always a $\lambda_{\alpha}\in I_{\mu}$ such that, picking $E_{\lambda_{\alpha}}$, Alice can be sure of getting a reward that is at least as high as Bob’s one, no matter what the observation that round will be.
+Notably, the converse is not true: there are $\lambda\in I_{\mu}$ such that no e-variable $E^{\mathrm{H}}_{\alpha}\in\mathcal{E}^{\mathrm{Hoeff}}_{\mu}$ dominates $E_{\lambda}$ for every possible value of $x$.
+This is formalised in the next statement, whose simple proof relies on elementary calculus (see A.1).
+Proposition 2
+Fix $\mu\in(0,1)$.
+For each $\alpha\in\mathbb{R}$ there is $\lambda_{\alpha}\in I_{\mu}$ such that $E_{\lambda_{\alpha}}(x)\geq E^{\mathrm{H}}_{\alpha}(x)$ for all $x\in\mathcal{X}$.
+On the other hand, fix any non-zero $\lambda\in I_{\mu}$.
+For every $\alpha\in\mathbb{R}$, there is at least a point $x_{\alpha}\in\mathcal{X}$ such that $E^{\mathrm{H}}_{\alpha}(x_{\alpha})<E_{\lambda}(x_{\alpha})$.
+As a consequence of the above discussion, when testing $\mathcal{H}_{\mu}$ it is always “better” to restrict the player to $\mathcal{E}^{\mathrm{cb}}_{\mu}$ rather than to $\mathcal{E}^{\mathrm{Hoeff}}_{\mu}$.
+We will make this intuition more formal in the next section.
+3 Majorising e-classes and optimal e-class
+Ideally, one aims to set up a powerful testing procedure, capable of rejecting the null as soon as there is enough evidence against it.
 However, achieving this depends on the strategy employed in the testing games.
-For example, stubbornly playing $E_{t}\equiv 1$ at all rounds in every game would produce valid confidence sets ($S_{n}=(0,1)$, for every $n$) of no practical interest.
-Indeed, a player’s strategy is most effective when it can rapidly increase the cumulative reward, whenever possible, as this allows to fast exclude redundant points from the confidence sets.
-In short, the highest the rewards, the tightest the confidence sequence.
-With this in mind, it is clear that carelessly restricting the player’s choice to a subset $\mathcal{E}_{\mu}$ of $\mathcal{E}_{\mathcal{H}_{\mu}}$ could be highly detrimental, as it might force the player to adopt poor strategies (e.g., consider the trivial case $\mathcal{E}_{\mu}=\{1\}$).
+For example, stubbornly playing $E_{t}\equiv 1$ at all rounds in every game would produce powerless tests of no practical interest.
+Indeed, a player’s strategy is most effective when it can rapidly increase the cumulative reward, whenever possible.
+In short, the highest the rewards, the more powerful the statistical test.
+We have already seen at the end of the previous section that there is no point in considering a $\mathcal{E}^{\mathrm{Hoeff}}_{\mu}$-restricted game, as this is always outperformed by the $\mathcal{E}^{\mathrm{cb}}_{\mu}$-restricted one.
+An even worse option would be the trivial restriction to =$\mathcal{E}=\{1\}$, which can never lead to rejection.
+With this in mind, it is clear that carelessly restricting the player’s choice to a subset of $\mathcal{E}_{\mathcal{H}_{\mu}}$ could be highly detrimental, as it might force the player to adopt poor strategies.
 This point naturally raises the question: “Does restricting the player to the coin-betting e-class (2) loosen the confidence sequence?”.
 Interestingly, for the round-wise testing-by-betting framework that we are considering the answer turns out to be negative.
-3 Majorising e-classes and optimal e-class
-We introduce a partial ordering on the set of real functions on a set $\mathcal{X}$.
-Given two functions $f,f^{\prime}:\mathcal{X}\to\mathbb{R}$, we say that $f$ majorises $f^{\prime}$, and write $f\succeq f^{\prime}$, if $f(x)\geq f^{\prime}(x)$ for all $x\in\mathcal{X}$.
-If $f\succeq f^{\prime}$ and there exists a point $x\in\mathcal{X}$ such that $f(x)>f^{\prime}(x)$, we say that $f$ is a strict majoriser of $f^{\prime}$, and we write $f\succ f^{\prime}$.
-Definition 3.
-$E\in\mathcal{E}_{\mathcal{H}}$ is called a maximal e-variable when there is no $E^{\prime}\in\mathcal{E}_{\mathcal{H}}$ such that $E^{\prime}\succ E$.
+In order to make this statement rigorous we now introduce the concepts of majorising and optimal e-class.
+First, we endow the set of real functions on a set $\mathcal{Z}$ with a partial ordering.
+Given two functions $f,f^{\prime}:\mathcal{Z}\to\mathbb{R}$, we say that $f$ majorises $f^{\prime}$, and write $f\succeq f^{\prime}$, if $f(z)\geq f^{\prime}(z)$ for all $z\in\mathcal{Z}$.
+If $f\succeq f^{\prime}$ and there is a $z\in\mathcal{Z}$ such that $f(z)>f^{\prime}(z)$, we say that $f$ is a strict majoriser of $f^{\prime}$, and we write $f\succ f^{\prime}$.
+Definition 3
+An e-variable $E\in\mathcal{E}_{\mathcal{H}}$ is called maximal if there is no $E^{\prime}\in\mathcal{E}_{\mathcal{H}}$ such that $E^{\prime}\succ E$.
 Next, we introduce a way to compare different e-classes.
-Definition 4.
+Definition 4
 Given two e-classes $\mathcal{E}$ and $\mathcal{E}^{\prime}$, we say that $\mathcal{E}$ majorises $\mathcal{E}^{\prime}$ if, for any $E^{\prime}\in\mathcal{E}^{\prime}$, there is an e-variable $E\in\mathcal{E}$ such that $E\succeq E^{\prime}$.
-We call majorising e-class any e-class that majorises $\mathcal{E}_{\mathcal{H}}$.
-Lemma 2.
+An e-class is said to be a majorising e-class if it majorises $\mathcal{E}_{\mathcal{H}}$.
+Lemma 1
 Every majorising e-class contains all the maximal e-variables.
-Proof.
+Proof 2
 Let $E\in\mathcal{E}_{\mathcal{H}}$ be maximal and $\mathcal{E}$ a majorising e-class.
 There must be $E^{\prime}\in\mathcal{E}$ such that $E^{\prime}\succeq E$, but since $E$ is maximal it has to be that $E=E^{\prime}$.
-Hence, $E\in\mathcal{E}$.
-∎
+Hence, $E\in\mathcal{E}$$\qed$
 The significance of the notion of majorising e-class for our problem is straightforward: if $\mathcal{E}$ majorises $\mathcal{E}^{\prime}$, then any strategy in an $\mathcal{E}^{\prime}$-restricted game can be matched or outperformed (in terms of rewards) by a corresponding strategy in the $\mathcal{E}$-restricted game, regardless of the sequence of observations.
 This allows us to compare how the restriction to different e-classes affects the testing-by-betting game of Definition 1.
-Definition 5.
-If a majorising e-class is contained in every other majorising e-class, we call it optimal.
+Notably, Proposition 2 implies that the coin-betting e-class always majorises the Hoeffding e-class.
+Definition 5
+If a majorising e-class is contained in every other majorising e-class, it is called optimal.
 For any $\mathcal{H}$, a majorising e-class always exists, as $\mathcal{E}_{\mathcal{H}}$ itself is a majorising e-class.
 However, an optimal e-class may not exist.
-[^4] We refer the reader to the follow-up work Clerico (2024) for an example of non-existence of the optimal e-class.
-Proposition 2 therein shows that =$\mathcal{H}=\{P\in\mathcal{P}_{[0,1]}\,:\,P(\{0\})\geq 1/2\}\cup\{U_{[0,1]}\}$ (with $U_{[0,1]}$ the uniform distribution on $[0,1]$) is a hypothesis for which the set of maximal e-variables is not a majorising e-class.
+[^5] We refer to the follow-up work Clerico (2024) for an example of non-existence of the optimal e-class.
+Proposition 2 therein shows that =$\mathcal{H}=\{P\in\mathcal{P}_{\mathcal{X}}\,:\,P(\{0\})\geq 1/2\}\cup\{U_{[0,1]}\}$ (with $U_{[0,1]}$ the uniform distribution on [0,1][0,1][ 0 , 1 ]) is a hypothesis for which the set of maximal e-variables is not a majorising e-class.
 Hence, $\mathcal{H}$ does not admit an optimal e-class.
-Next, we state a sufficient and necessary condition for the existence of optimal e-classes.
-Lemma 3.
+Next, we state a sufficient and necessary condition for its existence.
+Lemma 2
 An optimal e-class exists if, and only if, the set of all maximal e-variables is a majorising e-class.
 If an optimal e-class exists, it is unique, it corresponds to the set of all maximal e-variables, and it is the only majorising e-class whose elements are all maximal.
-Proof.
+Proof 3
 Denote as $\hat{\mathcal{E}}$ the set of all the maximal e-variables.
-Now, assume that there exists an optimal e-class $\mathcal{E}$.
-First, we show that all its elements must be maximal.
+Assume that there exists an optimal e-class $\mathcal{E}$.
+Let us show that all its elements are maximal.
 For $E\in\mathcal{E}$, consider any element $E^{\prime}\in\mathcal{E}_{\mathcal{H}}$ such that $E^{\prime}\succeq E$.
@@
 Since $E^{\prime}\succeq E$, it is clear that $\mathcal{E}^{\prime}$ majorises $\mathcal{E}$.
-It follows that $\mathcal{E}^{\prime}$ is a majorising e-class, as $\mathcal{E}$ is.
+So, $\mathcal{E}^{\prime}$ is a majorising e-class, as $\mathcal{E}$ is.
 Since $\mathcal{E}$ is optimal, $\mathcal{E}\subseteq\mathcal{E}^{\prime}$, which implies $E=E^{\prime}$.
 In particular, $E$ does not have any strict majoriser, and so it is maximal.
-Hence, $\mathcal{E}\subseteq\hat{\mathcal{E}}$.
+In particular, $\mathcal{E}\subseteq\hat{\mathcal{E}}$.
 As $\mathcal{E}\supseteq\hat{\mathcal{E}}$ by Lemma 1, we conclude that $\mathcal{E}=\hat{\mathcal{E}}$, and so $\hat{\mathcal{E}}$ is a majorising e-class.
@@
 So, $\hat{\mathcal{E}}$ is contained in all the majorising e-classes, and hence it is optimal.
-Now, the final statements are a trivial consequence of what was shown above and Lemma 2. ∎
+Now, the remaining statements are a trivial consequence of what was shown above and Lemma 1$\qed$
 Let us emphasise once more that, from our discussion thus far, it is clear that restricting the testing-by-betting game of Definition 1 to a majorising e-class does not hinder the performance of the player, as for any unrestricted strategy $(E_{t})_{t\geq 1}$ they can always pick a restricted strategy $(E_{t}^{\prime})_{t\geq 1}$, whose cumulative rewards inevitably match or outperform those of $(E_{t})_{t\geq 1}$, regardless of the sequence of observations.
@@
 Conversely, when a player selects an e-variable from the optimal e-class, no other choice can be guaranteed to be better before observing $x_{t}$, since the player’s pick is a maximal e-variable.
-As a straightforward consequence, within this round-wise testing-by-betting approach, the optimal approach to obtain confidence sequences consists in restricting game to the optimal e-class, for each mean candidate value $\mu$.
+As a straightforward consequence, within this round-wise testing-by-betting approach, the optimal approach to test $\mathcal{H}_{\mu}$ consists in restricting game to the optimal e-class.
 We show next that this coincides precisely with the coin-betting formulation.
@@
 We now show that the optimal e-class for $\mathcal{H}_{\mu}$ exists and coincides with the coin-betting e-class $\mathcal{E}^{\mathrm{cb}}_{\mu}$, defined in (2).
-Lemma 4.
-Fix $\mu\in(0,1)$ and consider the function $F_{\mu}:[0,1]\to[1,+\infty)$ defined as
+First, let us show that each e-variable is majorised by the function $F_{\mu}=\max(E_{\mu^{-1}},E_{(\mu-1)^{-1}})$.
+Lemma 3
+Fix $\mu\in(0,1)$ and consider the function $F_{\mu}:\mathcal{X}\to[1,+\infty)$ defined as
 […]
@@
 $F_{\mu}\;:\;x\mapsto\begin{cases}1+\frac{1}{\mu}(x-\mu)&\text{if $x\geq\mu$;}\\ 1+\frac{1}{\mu-1}(x-\mu)&\text{if $x<\mu$.}\end{cases}$
-For any $x\in[0,1]$ there is $P_{x}\in\mathcal{H}_{\mu}$ such that $P_{x}(\{x\})=1/F_{\mu}(x)>0$. Moreover, $F_{\mu}\succeq E$ for every $E\in\mathcal{E}_{\mathcal{H}_{\mu}}$.
-Proof.
-For $x\geq\mu$, let $P_{x}=\frac{\mu}{x}\delta_{x}+(1-\frac{\mu}{x})\delta_{0}$.
+For any $x\in\mathcal{X}$ there is $P_{x}\in\mathcal{H}_{\mu}$ such that $P_{x}(\{x\})=1/F_{\mu}(x)>0$. Moreover, $F_{\mu}\succeq E$ for all $E\in\mathcal{E}_{\mathcal{H}_{\mu}}$.
+Proof 4
+For $x\in\mathcal{X}\cap[\mu,1]$, let $P_{x}=\frac{\mu}{x}\delta_{x}+(1-\frac{\mu}{x})\delta_{0}$.
 Then, $P_{x}\in\mathcal{H}_{\mu}$ and $P_{x}(\{x\})=\mu/x=1/F_{\mu}(x)$.
-With a similar reasoning, if $x<\mu$ we can find a measure $P_{x}$ in $\mathcal{H}_{\mu}$, supported on $\{x,1\}$, with mass $1/F_{\mu}(x)$ on $x$.
-Now, to check that $F_{\mu}$ majorises all the e-variables, fix $E\in\mathcal{E}_{\mathcal{H}_{\mu}}$ and $x\in[0,1]$.
-Consider a $P_{x}\in\mathcal{H}_{\mu}$ with mass $1/F_{\mu}(x)$ on $x$.
-Then, $1\geq\mathbb{E}_{P_{x}}[E]\geq P_{x}(x)E(x)=E(x)/F_{\mu}(x)$, and we conclude.
-∎
-Theorem 1.
+Similarly, if $x<\mu$ we can find a measure $P_{x}$ in $\mathcal{H}_{\mu}$, supported on $\{x,1\}$, with mass $1/F_{\mu}(x)$ on $x$.
+To check that $F_{\mu}$ majorises all the e-variables, fix $E\in\mathcal{E}_{\mathcal{H}_{\mu}}$ and $x\in\mathcal{X}$.
+Let $P_{x}\in\mathcal{H}_{\mu}$ have mass $1/F_{\mu}(x)$ on $x$.
+Then, $1\geq\mathbb{E}_{P_{x}}[E]\geq P_{x}(\{x\})E(x)=E(x)/F_{\mu}(x)$, and we conclude.
+$\qed$
+Theorem 1
 For any $\mu\in(0,1)$, the coin-betting e-class $\mathcal{E}^{\mathrm{cb}}_{\mu}$ is the optimal e-class for $\mathcal{H}_{\mu}$.
-Proof.
+Proof 5
 First, let us show that $\mathcal{E}^{\mathrm{cb}}_{\mu}$ is a majorising e-class.
@@
 Define the sets
-$B_{0}=\left\{\beta\in I_{\mu}\;:\;\inf_{x\in[0,\mu)}\big{(}E_{\beta}(x)-E(x)\big{)}\geq 0\right\}\quad\text{and}\quad B_{1}=\left\{\beta\in I_{\mu}\;:\;\inf_{x\in(\mu,1]}\big{(}E_{\beta}(x)-E(x)\big{)}\geq 0\right\}\,,$
+$B_{0}=\left\{\beta\in I_{\mu}\;:\;\inf_{x\in\mathcal{X}\cap[0,\mu)}\big{(}E_{\beta}(x)-E(x)\big{)}\geq 0\right\}\;\text{and}\,\;B_{1}=\left\{\beta\in I_{\mu}\;:\;\inf_{x\in\mathcal{X}\cap(\mu,1]}\big{(}E_{\beta}(x)-E(x)\big{)}\geq 0\right\}\,,$
 where we recall that $I_{\mu}=[(\mu-1)^{-1},\mu^{-1}]$ and $E_{\beta}:x\mapsto 1+\beta(x-\mu)$.
@@
 Then, $\beta^{\star}\notin B_{0}$ and $\beta^{\star}\notin B_{1}$.
-In particular, there are $u_{0}\in[0,\mu)$ and $u_{1}\in(\mu,1]$ such that $E(u_{0})>E_{\beta^{\star}}(u_{0})$ and $E(u_{1})>E_{\beta^{\star}}(u_{1})$.
+In particular, there are $u_{0}<\mu$ and $u_{1}<\mu$, in $\mathcal{X}$, such that $E(u_{0})>E_{\beta^{\star}}(u_{0})$ and $E(u_{1})>E_{\beta^{\star}}(u_{1})$.
 As $\mu\in(u_{0},u_{1})$, there is $\hat{P}\in\mathcal{H}_{\mu}$ with support $\{u_{0},u_{1}\}$.
@@
 But $E$ is strictly larger than $E_{\beta^{\star}}$ on $\mathrm{Supp}(\hat{P})$, and so $\mathbb{E}_{\hat{P}}[E]>1$, which is a contradiction since $E$ is an e-variable.
-We conclude that $\beta_{0}\geq\beta_{1}$ and so there exists $\hat{\beta}\in B_{0}\cap B_{1}$.
-By construction, $E_{\hat{\beta}}\in\mathcal{E}^{\mathrm{cb}}_{\mu}$ and $E_{\hat{\beta}}(x)\geq E(x)$ for all $x\in[0,1]$ different from $\mu$.
-For $x=\mu$, by Lemma 4 we have $E(\mu)\leq F_{\mu}(\mu)=1=E_{\hat{\beta}}(\mu)$, and so $E\preceq E_{\hat{\beta}}$.
+So, $\beta_{0}\geq\beta_{1}$, and there exists $\hat{\beta}\in B_{0}\cap B_{1}$.
+By construction, $E_{\hat{\beta}}\in\mathcal{E}^{\mathrm{cb}}_{\mu}$ and $E_{\hat{\beta}}(x)\geq E(x)$ for all $x\in\mathcal{X}$ different from $\mu$.
+If $x=\mu$ by Lemma 3 we have $E(\mu)\leq F_{\mu}(\mu)=1=E_{\hat{\beta}}(\mu)$, so $E\preceq E_{\hat{\beta}}$.
 As the choice of $E$ was arbitrary, $\mathcal{E}^{\mathrm{cb}}_{\mu}$ is a majorising e-class.
-Now that we have established that $\mathcal{E}^{\mathrm{cb}}_{\mu}$ is a majorising e-class, by Lemma 3 we only need to show that all its elements are maximal to conclude that it is optimal.
+Once established that $\mathcal{E}^{\mathrm{cb}}_{\mu}$ is a majorising e-class, by Lemma 2 we only need to show that all its elements are maximal.
 Fix $E\in\mathcal{E}^{\mathrm{cb}}_{\mu}$, and consider an e-variable $\hat{E}\in\mathcal{E}_{\mathcal{H}_{\mu}}$ such that $\hat{E}\succeq E$.
-Fix any $x\in[0,1]$.
+Fix any $x\in\mathcal{X}$.
 By Lemma 3, there is $P\in\mathcal{H}_{\mu}$ such that $P(\{x\})>0$.
@@
 Since $P(\{x\})>0$, we get $\hat{E}(x)=E(x)$ and, $x$ being arbitrary, $\hat{E}=E$.
-Hence $E$ is maximal, as it has no strict majoriser.
-∎
+Hence, $E$ is maximal, as it has no strict majoriser.
+$\qed$
+5 From mean testing to confidence sequences
+Before moving to the more complex setting of multi-round e-variables and e-processes, which allow for tests that challenge the dependence structure of the observations, we first illustrate how sequential testing can be directly applied to the problem of mean estimation.
+In particular, we show how this framework naturally gives rise to confidence sequences, a sequential counterpart to classical confidence intervals that dates at least back to Darling and Robbins (1967).
+We consider the following approach to constructing confidence sequences via hypothesis testing: at each time step, we test each candidate value $\mu$ for the mean, and include in the confidence set those values that are not rejected.
+For further discussion on the connection between sequential testing and confidence sequences, we refer to Ramdas et al.
+(2022a).
+As usual, $(X_{t})_{t\geq 1}$ is a sequence of independent draws from $P^{\star}\in\mathcal{P}(\mathcal{X})$, whose mean $\mu^{\star}\in(0,1)$ has to be estimated.
+Let $\mathcal{F}=(\mathcal{F}_{t})_{t\geq 0}$ represent the natural filtration generated by $(X_{t})_{t\geq 1}$, where $\mathcal{F}_{t}=\sigma(X_{1},\dots,X_{t})$ captures all information available up to time $t$.
+Fix a confidence level parameter $\delta\in(0,1)$.
+A confidence sequence $(S_{t})_{t\geq 1}$ is a sequence of random sets [^6] Here, by random sets we simply mean sets that depend on the sequence of random observations $(X_{t})_{t\geq 1}$.
+such that the sequence of events ($(\{\mu^{\star}\in S_{t}\})_{t\geq 1}$ is adapted to the filtration $\mathcal{F}$ (i.e., for all $t\geq 1$, $\{\mu^{\star}\in S_{t}\}$ is $\mathcal{F}_{t}$-measurable) and satisfies
+$\mathbb{P}\big{(}\mu^{\star}\in S_{t}\,,\;\forall t\geq 1\big{)}\geq 1-\delta\,.$
+Intuitively, this means that $(S_{t})_{t\geq 1}$ provides a set of plausible values for $\mu^{\star}$ at each time step $t$, while ensuring that the true mean remains in these sets indefinitely with high probability.
+We can leverage the testing-by-betting game of Definition 1 to obtain a confidence sequence for the mean $\mu^{\star}$ of $P^{\star}$.
+For $\mu\in(0,1)$, we define the hypothesis $\mathcal{H}_{\mu}$ as in (1), namely $\mathcal{H}_{\mu}$ contains all probability measures on $\mathcal{X}$ with mean $\mu$.
+For each $\mu$, we fix an e-class $\mathcal{E}_{\mu}\subseteq\mathcal{E}_{\mathcal{H}_{\mu}}$ and we consider an $\mathcal{E}_{\mu}$-restricted testing-by-betting game on $\mathcal{H}_{\mu}$, where the player observes $(X_{t})_{t\geq 1}$, the sequence of draws from $P^{\star}$.
+For each one of these games, denote as $R_{n}(\mu)$ the player’s cumulative reward at round $n$.
+We can then construct a confidence sequence as follows.
+Proposition 3
+The sequence $(S_{n})_{n\geq 1}$, defined as
+$S_{n}=\big{\{}\mu\in(0,1)\;:\;R_{n}(\mu)\leq\log\tfrac{1}{\delta}\big{\}}\,,$
+is a confidence sequence for the mean $\mu^{\star}$ of the data-generating probability measure $P^{\star}$.
+Proof 6
+For each $n\geq 1$, we have that $\{\mu^{\star}\in S_{n}\}=\{R_{n}(\mu^{\star})\leq\log\frac{1}{\delta}\}$, which is a $\mathcal{F}_{n}$-measurable event. Moreover,
+$\mathbb{P}\big{(}\mu^{\star}\in S_{n}\,,\;\forall n\geq 1\big{)}=\mathbb{P}\big{(}R_{n}(\mu^{\star})\leq\log\tfrac{1}{\delta}\,,\;\forall n\geq 1\big{)}\geq 1-\delta$
+by Proposition 1, as the observations are drawn from $P^{\star}\in\mathcal{H}_{\mu^{\star}}$. $\qed$
+It is worth stressing that the strength of the resulting confidence sequence depends directly on the power of the underlying sequential tests for each hypothesis $\mathcal{H}_{\mu}$: more powerful tests yield tighter confidence sets.
+In particular, the optimality result for the coin-betting e-class established in Section 4 implies that, for each $\mu$, the testing game should be restricted to the class $\mathcal{E}^{\mathrm{cb}}_{\mu}$.
+With this choice, we recover the coin-betting framework to derive confidence sequences leveraged by Orabona and Jun (2023) and Waudby-Smith and Ramdas (2023).
+Both Orabona and Jun (2023) and Waudby-Smith and Ramdas (2023) propose explicit strategies for placing bets in the coin-betting game, leading to concrete confidence sequences.
+While the present work’s focus is on defining the optimal betting game rather than designing specific strategies, it is still useful to look more closely at the approach of Orabona and Jun (2023) to illustrate how the optimality of the coin-betting e-class simplifies the construction of a strategy for the game of Definition 1.
+They employ a coin-betting strategy based on the universal portfolio algorithm, a special instance of online learning Bayesian aggregation techniques (see, e.g., Chapters 9 and 10 of Cesa-Bianchi and Lugosi, 2006).
+In this approach, a prior distribution $\rho_{1}$ is fixed over the decision space $I_{\mu}$, and then updated sequentially using observed data.
+Specifically, at round $t$, one defines
+$\mathrm{d}\rho_{t}(\lambda)=\zeta_{t}^{-1}\prod_{i=1}^{t-1}(1+\lambda x_{i})\mathrm{d}\rho_{1}(\lambda)\,,$
+where $\zeta_{t}$ is the normalising constant ensuring that $\rho_{t}$ is a probability measure on $I_{\mu}$.
+The bet $\lambda_{t}$ is then chosen as the posterior mean: $\lambda_{t}=\int_{I_{\mu}}\lambda\mathrm{d}\rho_{t}(\lambda)$.
+This strategy can be interpreted as performing Bayesian averaging over the coin-betting e-class $\mathcal{E}^{\mathrm{cb}}_{\mu}$, leveraging its simple and low-dimensional parametric structure.
+This is a concrete example of how knowing explicitly the optimal e-class can help designing powerful testing-by-betting strategies.
+Indeed, such approach would not be feasible on the full class of all e-variables, an infinite dimensional functional space where even the definition of a prior can become problematic, if we did not know that the prior should be supported on the coin-betting e-class.
+Importantly, the optimality of $\mathcal{E}^{\mathrm{cb}}_{\mu}$ ensures that no statistical power is sacrificed by restricting to this class.
+Notably, a similarly structured averaging strategy over a suboptimal e-class, such as the Hoeffding e-class or its convex hull, would remain well defined and computationally tractable, but lead to strictly worse performance, as shown by Proposition 2.
 6 Optimality beyond single-round e-variables
-Up to this point, we have focused on hypotheses defined as subsets of the space $\mathcal{P}_{[0,1]}$, which cannot capture dependencies across multiple rounds.
-In such setting, the dependency structure among observations is fixed and assumed a priori, rather than being subject to testing.
+Up to this point, we have focused on hypotheses defined as subsets of the space $\mathcal{P}_{\mathcal{X}}$, which cannot capture depedences across multiple rounds.
+In such setting, the dependence structure among observations was fixed and assumed a priori, rather than being subject to testing.
 Yet, e-variable-based testing naturally extends to more general hypotheses that span multiple rounds and can account for sequential or dependent data structures.
 In the second part of the paper, we turn our attention to this richer framework.
-Remarkably, a slight adaptation of the proof strategy used for Theorem 1 and depicted in Figure 1 allows us to show that coin-betting-based testing is also optimal in the setting where the sequence of observations has a fixed conditional mean $\mu$.
-We will first show that this is the case with multi-round e-variables for a fixed time horizon, and then extend the result to testing with e-processes.
+Remarkably, an adaptation of the proof strategy used for Theorem 1 and depicted in Figure 1 allows us to show that coin-betting-based testing is also optimal in the setting where the sequence of observations has a fixed conditional mean $\mu$.
+We show that this is the case with multi-round e-variables for a fixed time horizon, and then extend the result to testing with e-processes.
 Conversely, we will show that when testing the i.i.d.
 assumption, the coin-betting approach no longer yields the optimal e-class.
-For technical reasons (primarily related to measurability, see discussion in Section 7), for the rest of this work we restrict ourselves to the case of a random variable taking values in a finite (known) set.
-Therefore, henceforth we let $\mathcal{X}\subseteq\mathbb{R}$ denote a finite set.
-For simplicity, we assume that $\min\mathcal{X}=0$ and $\max\mathcal{X}=1$, although all our results extend readily to any case where the minimum and maximum values are finite.
-We note that since $\mathcal{X}$ has finite cardinality, every function $\mathcal{X}^{t}\to\mathbb{R}$ is automatically Borel measurable.
-5.1 Optimal e-class for multi-round e-variables
+6.1 Optimality with multi-round e-variables
 For any $\mu\in(0,1)$ and $T\geq 1$, we let
@@
 As such, its set of e-variables will consist of Borel functions from $\mathcal{X}^{T}$ to $\mathbb{R}$.
-We denote the set of all e-values relative to $\mathcal{H}_{\mu}^{T}$ as $\mathcal{E}_{\mu}^{T}$.
-In a coin-betting game, the wealth of the player at round $T$ is a non-negative function of the observations, which takes the form $M_{T}(x^{T})=\prod_{t=1}^{n}E_{t}(x_{t})$, where $E_{t}$ is in the form $E_{\lambda_{t}}$, with $\lambda_{t}$ chosen as a function of the past observations $x^{t-1}$.
+We denote the set of all e-variablues relative to $\mathcal{H}_{\mu}^{T}$ as $\mathcal{E}_{\mu}^{T}$.
+Let us consider a test where all the $T$ observations $x^{T}$ are seen at once, where a single e-variable $E\in\mathcal{E}_{\mu}^{T}$ needs to be selected.
+If $E(x^{T})\geq 1/\delta$, the null $\mathcal{H}^{T}_{\mu}$ is rejected.
+We remark that this setting departs from the sequential testing-by-betting games discussed in the previous sections, as now the data set size $T$ is fixed in advance and the player makes a single decision before seeing the entire data set.
+In a way, this is as if we where playing a single round in the game of Definition 1, with $\mathcal{Z}=\mathcal{X}^{T}$.
+[^8] Of course, this can also be extended in a sequential game, where each round a new block of $T$ data points is observed.
+Despite this difference, the connection to the $T$-round coin-betting game of Definition 2 remains strong.
+Indeed, in the coin-betting setup that we have considered earlier, the wealth of the player at round $T$ is a non-negative function of the observations, which takes the form $M_{T}(x^{T})=\prod_{t=1}^{n}E_{t}(x_{t})$, where $E_{t}$ is in the form $E_{\lambda_{t}}$, with $\lambda_{t}$ chosen as a function of the past observations $x^{t-1}$.
 It is straightforward to check that under the hypothesis $\mathcal{H}_{\mu}^{T}$, the expectation of $M_{T}$ is always exactly one ($M_{T}$ is a martingale), which shows that $M_{T}$ is an e-variable.
-Next, we show that the e-class of e-variables in this form coincides is the optimal e-class for $\mathcal{H}_{\mu}^{T}$.
+Next, we show that the e-class of e-variables in this form coincides with the optimal e-class for $\mathcal{H}_{\mu}^{T}$.
 We let $\Lambda^{T}_{\mu}$ be a set of $T$-tuples of functions defined as
-$\Lambda^{T}_{\mu}=\big{\{}\lambda^{T}=(\lambda_{1},\dots,\lambda_{T})\,:\,\text{$\lambda_{t}$ is a function from $\mathcal{X}^{t-1}$ to $I_{\mu}$}\big{\}}\,.$
+$\Lambda^{T}_{\mu}=\big{\{}\lambda^{T}=(\lambda_{1},\dots,\lambda_{T})\,:\,\text{$\lambda_{t}$ is a Borel function from $\mathcal{X}^{t-1}$ to $I_{\mu}$}\big{\}}\,.$
 Given $\lambda^{T}\in\Lambda^{T}_{\mu}$, we define the e-variable $E_{\lambda^{T}}\in\mathcal{E}^{T}_{\mu}$ as
@@
 We define the T round coin-betting e-class $\mathcal{E}^{\mathrm{cb},T}_{\mu}=\{E_{\lambda^{T}}\,:\,\lambda^{T}\in\Lambda_{\mu}^{T}\}$.
-Theorem 2.
+Theorem 2
 Fix any $T\geq 1$ and $\mu\in(0,1)$. $\mathcal{E}^{\mathrm{cb},T}_{\mu}$ is the optimal e-class for $\mathcal{H}_{\mu}^{T}$.
-The proof of Theorem 2 shares many similarity with that of Theorem 1.
+The proof of Theorem 2 shares many similarities with that of Theorem 1.
 In particular, we use the same technique to establish the maximality of e-variables of the form $E_{\lambda^{T}}$.
-A geometric argument analogous to the one illustrated in Figure 1 is then employed, as part of an induction recursion, to prove that the e-class of interest is majorising.
-However, the detailed proof being somewhat lengthy, we defer it to the appendix.
+A geometric argument analogous to the one illustrated in Figure 1 can then be employed, as part of an induction recursion, to prove that the e-class of interest is majorising.
+However, the detailed proof is somewhat lengthy and involves some technical subtleties when dealing with measurability.
+We hence defer it to A.3.
 6.2 A remark on the i.i.d. case
-From what we have shown so far, in a $T$ round game coin-betting yields all and only the maximal e-variables, when testing the hypothesis that the conditional mean is some fixed $\mu\in(0,1)$.
-Yet, this is not any more the case if we consider a more restrictive hypotheses, stating that the observations are i.i.d.
+From what we have established so far, coin-betting yields all and only the maximal e-variables when testing the $T$-round hypothesis $\mathcal{H}_{\mu}^{T}$ that the conditional mean is some fixed $\mu\in(0,1)$.
+Yet, this is not any more the case if we consider a more restrictive hypothesis, stating that the observations are i.i.d.
 and with mean $\mu$.
-More concretely, let
+More concretely, let us define
 ^$\hat{\mathcal{H}}_{\mu}^{T}=\big{\{}P=Q^{\otimes T}\,:\,Q\in\mathcal{H}_{\mu}\big{\}}\,$
 […]
-To give a concrete example of the mismatch between the e-variables for the conditional and independent scenarios, we consider the “highly symmetrical” case of $\mathcal{X}=\{0,1/2,1\}$, with $\mu=1/2$ and $T=2$.
-We show at the end appendix that the e-variables in this setting are the functions $E:\{0,1/2,0\}^{2}\to[0,+\infty)$ that satisfy
+To give a concrete example of the mismatch between the e-variables for the conditional and independent scenarios, let us consider the “highly symmetrical” case $\mathcal{X}=\{0,1/2,1\}$, with $\mu=1/2$ and $T=2$.
+We show in A.2 that the e-variables in this setting are the $E:\{0,1/2,0\}^{2}\to[0,+\infty)$ satisfying
 $\xi_{0}\leq 1\,;\qquad\xi_{1}\leq 1+\sqrt{(1-\xi_{0})(1-\xi_{2})}\,;\qquad\xi_{2}\leq 1\,,$
 where $\xi_{0}=E(1/2,1/2)$, $\xi_{1}=(E(1,1/2)+E(1/2,1)+E(1/2,0)+E(0,1/2))/4$, and $\xi_{2}=(E(1,1)+E(1,0)+E(0,1)+E(0,0))/4$.
-Clearly, the optimal e-class here consists of those e-variables satisfying the constraint on $\xi_{1}$ with equality.
+Clearly, the optimal e-class here consists of those e-variables that satisfy $\xi_{0}=\xi_{1}=\xi_{2}=1$, or $\xi_{1}=1+\sqrt{(1-\xi_{0})(1-\xi_{2})}$ with $\xi_{0}<1$ and $\xi_{2}<1$.
 We note that all the e-variables in $\mathcal{E}_{1/2}^{T}$ satisfy $\xi_{0}=\xi_{1}=\xi_{2}=1$.
-As expected, these are indeed maximal e-variables for ^$\hat{\mathcal{H}}_{1/2}^{2}$, but there are maximal e-variables for ^$\hat{\mathcal{H}}_{1/2}^{2}$ that are not in $\mathcal{E}_{1/2}^{T}$.
-The function that equals $4$ on $x^{2}=(1,1/2)$, and 00 everywhere else, is an example of a (maximal) e-variable for ^$\hat{\mathcal{H}}_{1/2}^{2}$, which is not an e-variable for $\mathcal{H}_{1/2}^{2}$.
-We leave as a non-trivial open question to characterise the e-variables for ^$\hat{\mathcal{H}}^{T}_{\mu}$ in the general case.
-6 E-processes for testing the fixed conditional mean
-A key advantage of many testing procedures involving e-variables is anytime validity.
-One way to achieve this is through the round-wise testing-by-betting approach, which gives rise to a super-martingale (cf.
-Lemma 1).
+As expected, these are indeed maximal e-variables for ^$\hat{\mathcal{H}}_{1/2}^{2}$.
+Yet, there are maximal e-variables for ^$\hat{\mathcal{H}}_{1/2}^{2}$ that are not in $\mathcal{E}_{1/2}^{T}$.
+The function that equals 444 on $x^{2}=(1,1/2)$, and 0 everywhere else, is an example of a (maximal) e-variable for ^$\hat{\mathcal{H}}_{1/2}^{2}$, which is not an e-variable for $\mathcal{H}_{1/2}^{2}$.
+We leave as a (non-trivial) open question to characterise the e-variables for ^$\hat{\mathcal{H}}^{T}_{\mu}$ in the general case.
+6.3 Optimal classes of e-processes
+A key advantage of many testing procedures involving e-variables is allowing early stopping.
+One way to achieve this is through the round-wise testing-by-betting approach that we presented in Section 2, which gives rise to a super-martingale (cf.
+Proposition 1).
 However, as we have already pointed out, this is not the most general approach.
-Rather than only focusing on super-martingales, one can instead use any non-negative process whose expectation is upper bounded by one, regardless the stopping time.
+Rather than only focusing on super-martingales, one can in principle use any non-negative process whose expectation is upper bounded by one, regardless the stopping time.
 These are known as e-processes, a generalization of super-martingales, which can enable powerful tests even in scenarios where martingale-based methods lack power (Ramdas et al., 2022b).
-As in the previous section, here $\mathcal{X}$ denotes a finite subset of $[0,1]$, with $\min\mathcal{X}=0$ and $\max\mathcal{X}=1$.
 We denote as $\mathcal{X}^{\infty}$ the space of sequences $(x_{t})_{t\geq 1}\subseteq\mathcal{X}$.
-For convenience, we will sometimes write $x^{\infty}$ to mean $(x_{t})_{t\geq 1}$.
-We endow the set $\mathcal{X}^{\infty}$ with the product sigma-field generated by cylinder sets, and we denote as $\mathcal{P}_{\mathcal{X}^{\infty}}$ the space of probability measures on $\mathcal{X}^{\infty}$.
-For a fixed $\mu\in(0,1)$, in this section we will focus on the hypothesis class
-$\mathcal{H}_{\mu}^{\infty}=\big{\{}P\in\mathcal{P}_{\mathcal{X}^{\infty}}\,:\,\mathbb{E}_{P}[X_{t}|X^{t-1}]=\mu\,,\;\forall t\geq 1\big{\}}\,.$
+For convenience, we will often write $x^{\infty}$ for $(x_{t})_{t\geq 1}$.
+We endow $\mathcal{X}^{\infty}$ with the product sigma-field generated by cylinder sets, and we denote as $\mathcal{P}_{\mathcal{X}^{\infty}}$ the space of probability measures on $\mathcal{X}^{\infty}$.
+For a fixed $\mu\in(0,1)$, in this section we focus on the hypothesis class
+$\mathcal{H}_{\mu}^{\infty}=\big{\{}P\in\mathcal{P}_{\mathcal{X}^{\infty}}\,:\,\mathbb{E}_{P}[X_{t}|X^{t-1}]=\mu\,,\;\forall t\geq 1\big{\}}\,,$
+where again we use the convention $\mathbb{E}_{P}[X_{1}|X^{0}]=\mathbb{E}_{P}[X_{1}]$.
 This time, we will not aim to study the e-variables for $\mathcal{H}^{\infty}_{\mu}$, as these would be functions that take a whole sequence as argument.
@@
 To make this rigorous we first need some definitions.
-First, we call a finite stopping time a function $\tau:\mathcal{X}^{\infty}\to\mathbb{N}$ such that, whenever for some $x^{\infty}$ and $y^{\infty}$ in $\mathcal{X}^{\infty}$ and some $t\in\mathbb{N}$ we have $\tau(x^{\infty})=\tau(y^{\infty})=t$, then $x^{t}=y^{t}$.
-[^8] It is straightforward to verify that in our setting this definition is equivalent to the perhaps more classical notion of finite e-process as a $\tau$ such that the sets $\{\tau\leq t\}$ are $\mathcal{F}_{t}$ measurable.
-We say that $\tau$ is bounded by $T\geq 0$ if its image is contained in $[0:T]$.
-We let $\mathcal{T}_{T}$ denote the set of all stopping times bounded by $T$.
-Definition 6.
-An e-process (for $\mathcal{H}_{\mu}^{\infty}$) is a sequence $(E_{t})_{t\geq 0}$, of non-negative functions $E_{t}:\mathcal{X}^{t}\to[0,+\infty)$, such that, for every $P\in\mathcal{H}^{\infty}_{\mu}$, any $T\geq 0$ and any $\tau\in\mathcal{T}_{T}$,
-$\mathbb{E}_{P}[E_{\tau}]\leq 1\,.$
-We denote as $\mathcal{E}_{\mu}^{\infty}$ the set of all the e-processes for $\mathcal{H}_{\mu}^{\infty}$.
-We call e-process class any subset of $\mathcal{E}_{\mu}^{\infty}$.
+First, we define a finite stopping time as a measurable [^13] Here we endow $\mathbb{N}$ with the discrete sigma-field.
+function $\tau:\mathcal{X}^{\infty}\to\mathbb{N}$ such that, for any $t\geq 0$ the set $\{\tau=t\}$ is measurable with respect to the sigma-field $\mathcal{F}_{t}$, generated by the projection on the first $t$ components of the sequences in $\mathcal{X}^{\infty}$.
+We let $\mathcal{T}$ denote the set of all finite stopping times.
+Definition 6
+Let $E=(E_{t})_{t\geq 0}$ be a sequence of non-negative Borel functions $E_{t}:\mathcal{X}^{t}\to[0,+\infty)$.
+We say that $E$ is an e-process (for $\mathcal{H}_{\mu}^{\infty}$) when, for any $P\in\mathcal{H}^{\infty}_{\mu}$ and any $\tau\in\mathcal{T}$, $\mathbb{E}_{P}[E_{\tau}]\leq 1$.
+Sequential testing using e-processes proceeds as follows (see Ramdas and Wang 2024 for more details).
+Before observing any data, an e-process $E=(E_{t})_{t\geq 0}$ for $\mathcal{H}_{\mu}^{\infty}$ is fixed.
+[^14] Note that although the e-process is fixed before observing any data, each $E_{t}$ is a function of the past observations $x^{t-1}$, allowing the test to adapt to the data as they become available.
+Then, at each round $t$, the player’s wealth is defined as $\log E_{t}(x^{t})$.
+$\mathcal{H}_{\mu}^{\infty}$ is rejected if this wealth ever exceeds $\log 1/\delta$.
+The definition of an e-process ensures that this yields a sequential test with Type I error controlled at level $\delta$.
+We remark that this testing procedure extends the testing-by-betting framework of Definition 1 to the more expressive setting where the null hypothesis takes into account the entire dependence structure of the data sequence.
+We denote as $\mathcal{E}_{\mu}^{\infty}$ the set of all e-processes for $\mathcal{H}_{\mu}^{\infty}$.
+An e-process class is any subset of $\mathcal{E}_{\mu}^{\infty}$.
 Similarly to what we have done for the e-variables, we say that the e-process $E$ majorises the e-process $E^{\prime}$ (we write $E\succeq E^{\prime}$) if for every $t\geq 0$ we have $E_{t}\succeq E^{\prime}_{t}$.
-We say that an e-process class $\mathcal{E}$ is majorising if for every $E\in\mathcal{E}^{\infty}_{\mu}$, there is $E^{\prime}\in\mathcal{E}$ such that $E\preceq E^{\prime}$.
-We say that $\mathcal{E}$ is optimal if it is majorising and is contained in every other majorising e-process class.
-We call an e-process $E$ maximal if it is not strictly majorised by any other e-process.
-With the same reasoning as in Lemma 3, one can show that the optimal e-process class exists if, and only if, the class of all the maximal e-processes is majorising.
-In such case the optimal class is unique and coincides precisely with all and only the majorising e-processes.
-We let $\Lambda_{\mu}^{\infty}$ be the set of all sequences $\lambda^{\infty}=(\lambda_{t})_{t\geq 1}$, such that $\lambda_{t}:\mathcal{X}^{t-1}\to I_{\mu}$.
+If $E\succeq E^{\prime}$ and there is a $t\geq 0$ such that $E_{t}\succ E^{\prime}_{t}$, then $E$ strictly majorises $E^{\prime}$, and we write $E\succ E^{\prime}$.
+We call an e-process maximal if it is not strictly majorised by any other e-process.
+We say that an e-process class $\mathcal{E}$ is majorising if, for every $E\in\mathcal{E}^{\infty}_{\mu}$, there is $E^{\prime}\in\mathcal{E}$ such that $E\preceq E^{\prime}$.
+We say that $\mathcal{E}$ is optimal if it is majorising and all its elements are maximal.
+We remark that if an optimal majorising e-process class exists, then it is unique, it consists with the set of all the maximal e-processes, and it is included in every other majorising e-process class.
+We let $\Lambda_{\mu}^{\infty}$ be the set of all sequences $\lambda^{\infty}=(\lambda_{t})_{t\geq 1}$, with $\lambda_{t}:\mathcal{X}^{t-1}\to I_{\mu}$ Borel for all $t\geq 1$.
 For any $\lambda^{\infty}\in\Lambda_{\mu}^{\infty}$, we define $E^{\lambda^{\infty}}=(E^{\lambda^{\infty}}_{T})_{T\geq 0}$ via
@@
 for $T\geq 1$, and $E_{0}=1$.
-These are the e-processes associated to the wealth of a coin-betting player.
+These are the sequences associated with the wealth of a coin-betting player.
 As $E^{\lambda^{\infty}}$ defines a martingale under $\mathcal{H}_{\mu}^{\infty}$, it is also an e-process for such hypothesis.
-We define the coin-betting e-process class $\mathcal{E}^{\mathrm{cb},\infty}_{\mu}=\{E^{\lambda^{\infty}}\,:\,\lambda^{\infty}\in\Lambda^{\infty}_{\mu}\}$.
-The next result shows that for $\mathcal{H}_{\mu}^{\infty}$ the coin betting approach is optimal also at the level of e-processes.
-The proof (reported in the appendix) always follows the ideas introduced in the proof of Theorem 1.
-Theorem 3.
-For any $\mu\in(0,1)$, $\mathcal{E}^{\mathrm{cb},\infty}_{\mu}$ is the optimal e-process class for $\mathcal{H}_{\mu}^{\infty}$.
+We hence define the coin-betting e-process class
+$\mathcal{E}^{\mathrm{cb},\infty}_{\mu}=\{E^{\lambda^{\infty}}\,:\,\lambda^{\infty}\in\Lambda^{\infty}_{\mu}\}\,.$
+The next result shows that the coin-betting approach is optimal at the level of e-processes for $\mathcal{H}_{\mu}^{\infty}$.
+The proof (see A.4) follows the ideas introduced in the proofs of Theorems 1 and 2.
+Theorem 3
+For any $\mu\in(0,1)$ the coin-betting e-process class is the optimal e-process class for $\mathcal{H}_{\mu}^{\infty}$.
 7 Perspectives
-Theorem 1 gives a rigorous sense to the claim that the coin-betting formulation is optimal among the e-variable-based approaches to build confidence sequences for $\mu^{\star}$.
-Theorem 2 and Theorem 3 extended this optimality result to testing the fixed conditional mean of a sequence of bounded observations.
-To formalise these claims, we introduced the notions of majorising and optimal e-classes, which, in the author’s view, are of independent interest in the context of sequential testing, beyond the scope of this paper.
+Theorem 1 gives a rigorous sense to the claim that the coin-betting formulation is optimal among the e-variable-based approaches to test the mean and build confidence sequences given a sequence of independent draws from an unknown supported on [0,1][0,1][ 0 , 1 ].
+Theorems 2 and 3 extend this optimality result when testing the hypothesis of a fixed conditional mean.
+To formalise these claims, we introduced the notions of majorising and optimal e-classes, which may be of independent interest in the context of sequential testing, beyond the scope of this paper.
 The main novelty of these concepts lies in defining “optimality” in terms of e-classes, rather than individual e-variables.
 This perspective contrasts with the notion of log-optimality of a single e-variable with respect to an alternative hypothesis, widely discussed in the literature (e.g., Koolen and Grünwald, 2022; Grünwald et al., 2024; Larsson et al., 2024).
+Notably, log-optimality is defined in terms of an alternative hypothesis, against which the null is tested.
+In such setting there can be indeed a single (up to null sets under the alternative hypothesis) best e-variable.
+However, here we adopt a different perspective, where no alternative is defined.
+It is not hard to see that if the optimal e-class exists, whenever one considers an alternative that allows to define a log-optimal e-variable, a “version” [^15] As the log-optimal e-variable is defined up to null sets under the alternative, there might be elements of this family of e-variables that are not maximal.
+However, there is always a maximal e-variable among them if the optimal e-class exists.
+of this e-variable must lie in the optimal e-class.
+Interestingly, in the case of the coin-betting e-class, for each $E\in\mathcal{E}^{\mathrm{cb}}_{\mu}$ one can find an alternative hypothesis such that $E$ is log-optimal.
+[^16] Specifically, Theorem 1 in Grünwald et al.
+(2024) implies that, for $\lambda\in I_{\mu}$, $E_{\lambda}$ is log-optimal for the point alternative $\{Q\}$, with $Q=(1-\mu)E(0)\delta_{0}+\mu E(1)\delta_{1}$, since $E_{\lambda}=\mathrm{d}Q/\mathrm{d}P$ on the support of $Q$, where $P=(1-\mu)\delta_{0}+\mu\delta_{1}$ is in $\mathcal{H}_{\mu}$.
 The concept of maximality for e-variables, as introduced in this paper, is essentially equivalent to the classical statistical notion of admissibility.
 An e-variable is considered maximal if no other e-variable strictly dominates it.
-Likewise, the idea of a majorising e-class corresponds to the classical concept of a complete class, with the optimal e-class representing the minimal complete class of e-variables.
+Likewise, the idea of a majorising e-class corresponds to the notion of a complete class, with the optimal e-class representing the minimal complete class of e-variables.
+In this sense, the problem of identifying the optimal e-class can be seen as an instance of characterising the minimal complete class of tests, a question largely studied in classical statistics.
 For a detailed discussion of these ideas in traditional statistics we refer to Lehmann and Romano (2022).
-In this work, however, we adopt terminology from partially ordered set (poset) theory to highlight the fact that with e-variables these properties follows from the classical partial order of dominance among functions.
+In this work, however, we adopt terminology from partially ordered set (poset) theory to highlight the fact that with e-variables these properties follows from the standard dominance partial ordering among functions.
 The concept of admissibility in the context of e-variables and e-processes was previously introduced by Ramdas et al.
 (2022a), whose definition is closely aligned with our notion of maximality.
-The key distinction lies in the type of dominance considered: their framework often relies on almost sure dominance, whereas our definition of maximality requires everywhere dominance.
-[^9] The two definitions becomes of course equivalent when $\mathcal{X}$ is finite, but they are not in the case of $\mathcal{X}=[0,1]$, which is the setting of our analysis on single-round e-variables.
+However, a key distinction lies in the type of dominance considered: their framework often relies on almost sure dominance, whereas our definition of maximality requires everywhere dominance.
 This stricter requirement is motivated by the will of developing a theory that remains valid even in the absence of a known alternative, where any value in $\mathcal{X}$ might occur at the next observation.
 Although Ramdas et al.
-(2022a) provide some necessary and sufficient conditions for admissibility, we emphasize that their results are insufficient to establish the optimality of the coin-betting e-class.
-Specifically, their approach assumes the existence of a reference measure for the hypothesis, [^10] A hypothesis $\mathcal{H}$ admits a ”reference measure” if there exists a Borel measure $Q$ such that $P\ll Q$ for all $P\in\mathcal{H}$.
-a condition not satisfied by $\mathcal{H}_{\mu}$.
-More precisely, their results can show that each individual e-variable in the coin-betting e-class is maximal, but not that $\mathcal{E}^{\mathrm{cb}}_{\mu}$ forms a majorising e-class.
-This distinction, though subtle, is crucial: proving the optimality of coin-betting requires analyzing the collective properties of a set of e-variables (or equivalently, wealth processes), rather than evaluating them in isolation.
-It is precisely the fact that $\mathcal{E}^{\mathrm{cb}}_{\mu}$ is a majorising e-class that guarantees nothing is lost by relying on the coin-betting approach.
-This point is subtle but essential: establishing the optimality of coin-betting requires studying the properties of a set of e-variables (or equivalently, wealth-processes) rather than focusing on single e-variables or wealth-processes individually.
-Crucially, it is precisely the fact that $\mathcal{E}^{\mathrm{cb}}_{\mu}$ is a majorising e-class that ensures that nothing is lost by relying on the coin-betting approach.
+(2022a) provide some necessary and sufficient conditions for admissibility, we emphasise that their results are insufficient to establish the optimality of the coin-betting e-class.
+Specifically, their necessary conditions assume the existence of a reference measure for the considered hypothesis, [^17] A hypothesis $\mathcal{H}$ admits a ”reference measure” if there exists a Borel measure $Q$ such that $P\ll Q$ for all $P\in\mathcal{H}$.
+a requirement not satisfied by $\mathcal{H}_{\mu}$.
+More precisely, their findings imply that each individual e-variable in the coin-betting e-class is maximal, but not that $\mathcal{E}^{\mathrm{cb}}_{\mu}$ forms a majorising e-class.
+This distinction, though subtle, is crucial: proving the optimality of coin-betting requires analyzing the collective properties of a set of e-variables, rather than evaluating them in isolation.
+It is precisely the fact that $\mathcal{E}^{\mathrm{cb}}_{\mu}$ is a majorising e-class that guarantees that nothing is lost by relying on the coin-betting approach.
 To the author’s knowledge, this is the first work to rigorously examine the optimality of the coin-betting formulation, as previous discussion on the topic has relied solely on heuristic arguments.
-For instance, Waudby-Smith and Ramdas (2023) justify restricting to the coin-betting e-class by noticing that $\mathcal{E}^{\mathrm{cb}}_{\mu}$ is precisely the set ¯$\bar{\mathcal{E}}_{\mathcal{H}_{\mu}}$ of e-variables whose expectation equals one under every $P\in\mathcal{H}_{\mu}$.
-However, this property does not generally imply the optimality of the e-class, as it merely indicates that all its elements are maximal.
+For instance, Waudby-Smith and Ramdas (2023) justify restricting to the coin-betting e-class by noting that $\mathcal{E}^{\mathrm{cb}}_{\mu}$ is precisely the set ¯$\bar{\mathcal{E}}_{\mathcal{H}_{\mu}}$ of e-variables whose expectation equals 111 under every $P\in\mathcal{H}_{\mu}$.
+However, in general this property (often called exactness) merely implies that all the elements in the e-class are maximal, and not its optimality.
 As a simple counterexample, consider $\mathcal{H}^{\prime}_{\mu}=\{P\in\mathcal{P}\,:\,\langle P,X\rangle\leq\mu\}$.
-Then, ¯$\bar{\mathcal{E}}_{\mathcal{H}^{\prime}_{\mu}}=\{1\}$, while the optimal e-class exists and consists of all functions in the form $E_{\beta}:x\mapsto 1+\beta(x-\mu)$, with $\beta\in[0,\mu^{-1}]$.
-Although one might argue that for each of the maximal e-variable there is at least one measure that bring the expectation to one (Grünwald, 2024 calls such property sharpness), sharpness alone is not enough to ensure maximality ($\mathbf{1}_{\mu}$ is a sharp e-variable for $\mathcal{H}_{\mu}$, but it is clearly not maximal).
-Two papers (Clerico, 2024; Larsson et al., 2025) have appeared after the first preprint of this work, both extending the results on characterising single-round e-variables for hypotheses defined via linear constraints, a framework that includes the testing for the mean as a special case.
-Their analyses, however, rely on considerably more abstract and technically advanced proof techniques.
+Then, ¯$\bar{\mathcal{E}}_{\mathcal{H}^{\prime}_{\mu}}=\{1\}$, while the optimal e-class exists and consists of all functions in the form $E_{\beta}:x\mapsto 1+\beta(x-\mu)$, with $\beta\in[0,\mu^{-1}]$ (see Clerico, 2024).
+Although one might argue that for each maximal e-variable there is at least one measure that brings the expectation to 111 (Grünwald 2024 calls such property sharpness), this feature alone is not enough to ensure maximality ($\mathbf{1}_{\mu}$, the function equal to 111 on $\mu$ and 0 elsewhere, is a sharp e-variable for $\mathcal{H}_{\mu}$, but it is clearly not maximal).
+Two papers (Clerico, 2024; Larsson et al., 2025) have appeared after the first preprint of this work, both characterising single-round e-variables for hypotheses defined via linear constraints, a framework that includes the testing for the mean via single-round e-variables as a special case.
+Their results directly imply our Theorem 1 (but not Theorems 2 and 3) [^18] Interestingly, one could see the multi-round hypothesis $\mathcal{H}_{\mu}^{T}$ as a linearly constrained hypothesis, in the sense of Larsson et al.
+(2025).
+However, it does not seem to be trivial to directly derive Theorem 2 from their results..
+However, their analyses rely on considerably more abstract and technically advanced proof techniques.
+Clerico (2024) first proves results for finite domains and then extends them to the uncountable case by compactness and density arguments, while Larsson et al.
+(2024) leverages and extends powerful duality tools from the theory of functional lattices.
 Conversely, a key strength of the present paper is to provide a simple and neat argument (illustrated in Figure 1) that works well in the simple setting considered.
 Interestingly, this same argument is also at the base of the proofs of Theorem 2 and Theorem 3.
-We also remark that a similar strategy was adopted in the proof of Lemma 2 in Wang (2025).
-That said, for the multi-round setting, the approach does give rise to non-trivial measurability issues when $\mathcal{X}$ is uncountable.
-These difficulties already emerge in the two-round case: although one can show that every e-variable must be dominated by a product of the form $(1+\lambda_{1}x_{1})(1+\lambda_{2}(x_{1})x_{2})$, this is constructed in a way that gives no guarantee on the measurability of $\lambda_{2}$.
-Whether this issue can be resolved remains an open question.
-Nonetheless, even in the simplified setting we adopt by requiring $\mathcal{X}$ to have finite cardinality [^11] A countable $\mathcal{X}$ would also suffice to solve the measurability issue, though the final step of the proof of Theorem 3 would require more delicate handling., this work is, to the best of the author’s knowledge, the first to fully characterise all e-processes for the natural problem of testing a fixed conditional mean.
-To conclude, let us remark that this paper does not aim to delve into the design of effective coin-betting strategies for obtaining tight confidence sequences.
+We remark that a similar strategy was adopted in the proof of Lemma 2 in Wang (2025).
+Another interesting connection with this paper’s approach and the literature is recent work on admissible merging for e-variables (Vovk and Wang, 2024; Wang, 2025).
+A thorough exploration of these connections is an interesting avenue for future research.
+In particular, Wang (2025) implies that the coin-betting single-round e-variables cannot be majorised by any e-variable that is a monotonic function.
+However, the current paper approach relies on this monotonicity assumption, and cannot hence imply directly Theorem 1.
+On the other hand, Vovk and Wang (2024) considered conditional sequential hypotheses, of which conditional mean testing is a special case.
+Although their framework adopts a slightly different perspective, some of their results may imply, or be equivalent to, a weaker version of Theorem 3, where a finite time horizon $T$ is fixed and the e-processes are defined as vectors $(E_{t})_{t\in[0:T]}$, rather than sequences.
+As our proof techniques follow a different route compared to these two works, we believe our findings offer an alternative perspective that contribute to a broader understanding of the problem and could potentially help generalise or strengthen their results.
+To conclude, we remark that the present paper does not aim to delve into the design of effective coin-betting strategies.
 For this, we refer the interested reader to the thorough analysis and discussion by Orabona and Jun (2023) and Waudby-Smith and Ramdas (2023).
-Finally, although we have focused on random variables taking values in $[0,1]$ for clarity of exposition, all the results presented readily extend to any bounded closed interval.
+However, let us stress once more that characterising the optimal e-class simplifies the taks of designing such strageties, by provides the minimal set where e-values shall be picked.
+We discussed at the end of Section 5 an explicit example where this can be turned into a practical advantage, namely when using a Bayesian aggregation strategy that places a prior over the e-class considered.
+The tractable, parametric form of the coin-betting e-class makes such approach feasible, while its optimality ensures that no statistical power is lost.
+As a final comment, although we have focused on random variables taking values in $\mathcal{X}\subseteq[0,1]$ for clarity of exposition, most results presented can be easily extended to any bounded closed real set.
 Acknowledgements
@@
 Koolen for helpful discussions on e-processes.
-The author acknowledges the use of GPT-4o during the redaction of this paper, in order to polish the presentation.
-All AI-generated text was subsequently reviewed and edited by the author, who takes full responsibility for the content of this manuscript.
+Finally, I thank the anonymous reviewers for the insightful comments and feedback that helped improving this work.
+GPT-4o was used during the redaction of this paper to polish the presentation.
+All AI-generated text was reviewed and edited by the author, who takes full responsibility for the content of this manuscript.
 This project was funded by the European Research Council (ERC), under the European Union’s Horizon 2020 research and innovation programme (grant agreement 950180).
-Appendix
-Proof of Theorem 2
-Proof.
-We start by showing that the class of functions considered is an e-class, and that all its elements are maximal e-variables.
-So, fix any $\lambda^{T}_{\mu}\in\Lambda^{T}$.
-As all the $\lambda_{t}$ are valued in $I_{\mu}$, it is clear that $E_{\lambda^{T}}$ is non-negative.
-Moreover, for any $P\in\mathcal{H}_{\mu}^{T}$ it is easily checked that $\mathbb{E}_{P}[E_{\lambda^{T}}]=1$, which ensures that $E_{\lambda^{T}}$ is an e-variable.
-For the maximality we can use the same argument as in the proof of Theorem 1.
-Let $\hat{E}\in\mathcal{E}^{T}_{\mu}$ be such that $\hat{E}\succeq E_{\lambda^{T}}$.
-Fix any $x^{T}\in\mathcal{X}^{T}$.
-By Lemma 4, for each $t$ we can find a $P_{t}\in\mathcal{H}^{1}_{\mu}$ such that $P_{t}(\{x_{t}\})>0$.
-Let $P=P_{1}\otimes\dots\otimes P_{T}$.
-Then we have that $P(\{x^{T}\})>0$.
-In particular, we have that $0\leq P(\{x^{T}\})(\hat{E}(x^{T})-E_{\lambda^{T}}(x^{T}))\leq\mathbb{E}_{P}[\hat{E}-E_{\lambda^{T}}]=\mathbb{E}_{P}[\hat{E}]-1\leq 0$.
-Since $P(\{x^{T}\})>0$, $\hat{E}(x^{T})=E_{\lambda^{T}}(x^{T})$, and we conclude that $E_{\lambda^{T}}$ is maximal.
-To show that the e-class that we consider is majorising we proceed by induction on $T$.
-From Theorem 1 we know that the statement is true for $T=1$.
-For $T\geq 2$.
-let us assume that the claim holds for $T-1$, and let us show that it is also true for $T$.
-Let $E$ be an e-variable for $\mathcal{H}^{T}_{\mu}$.
-First, we show that, for any $Q\in\mathcal{H}^{T-1}_{\mu}$ and any $x\in\mathcal{X}$, we have
-$\mathbb{E}_{Q}[E(x,X_{2},\dots,X_{T})]\leq F_{\mu}(x)\,,$
-with $F_{\mu}$ as in Lemma 4.
-Fix $Q$ and $x\geq\mu$ in $\mathcal{X}$.
-Define $V_{x}=P_{x}\otimes Q$, where $P_{x}=\frac{\mu}{x}\delta_{x}+(1-\frac{\mu}{x})\delta_{0}$.
-Clearly, $V_{x}\in\mathcal{H}_{\mu}^{T}$.
-We have that $1\geq\mathbb{E}_{V_{x}}[E]=\tfrac{\mu}{x}\mathbb{E}_{Q}[E(x,X_{2},\dots,X_{T})]+\big{(}1-\tfrac{\mu}{x}\big{)}\mathbb{E}_{Q}[E(0,X_{2},\dots,X_{T})]$, and so $\mathbb{E}_{Q}[E(x,X_{2},\dots,X_{T})]\leq x/\mu=F_{\mu}(x)$.
-The case for $x\leq\mu$ is essentially analogous.
-In particular, the following two sets are non-empty:
-$B_{0}=\big{\{}\beta\in I_{\mu}\,:\,\forall x\in\mathcal{X}\cap[0,\mu)\,,\,\forall Q\in\mathcal{H}_{\mu}^{T-1}\,,\,\mathbb{E}_{Q}[E(x,X_{2},\dots,X_{T})]\leq 1+\beta(x-\mu)\big{\}}\,;$
-$B_{1}=\big{\{}\beta\in I_{\mu}\,:\,\forall x\in\mathcal{X}\cap(\mu,1]\,,\,\forall Q\in\mathcal{H}_{\mu}^{T-1}\,,\,\mathbb{E}_{Q}[E(x,X_{2},\dots,X_{T})]\leq 1+\beta(x-\mu)\big{\}}\,.$
-$B_{0}$ and $B_{1}$ are intersections of closed intervals (one per each allowed $x$ and $Q$), and as such they must be closed intervals.
-We can find $\beta_{0}$ and $\beta_{1}$ such that $B_{0}=[(\mu-1)^{-1},\beta_{0}]$ and $B_{1}=[\beta_{1},\mu^{-1}]$.
-We will show that $\beta_{0}\geq\beta_{1}$, and hence that $B_{0}\cap B_{1}\neq\varnothing$.
-The argument is essentially the same used in the proof of Theorem 1 (cf.
-also Figure 1).
-Assume that $\beta_{0}<\beta_{1}$, and let $\beta^{\star}\in(\beta_{0},\beta_{1})$.
-As $\beta^{\star}\notin B_{0}$, there is $Q_{0}\in\mathcal{H}^{T-1}_{\mu}$ and $u_{0}<\mu$ in $\mathcal{X}$ such that $\mathbb{E}_{Q_{0}}[E(u_{0},X_{2},\dots,X_{T})]>1+\beta^{\star}(u_{0}-\mu)$.
-Similarly, there must be $Q_{1}\in\mathcal{H}_{\mu}^{T-1}$ and $u_{1}>\mu$ in $\mathcal{X}$, such that $\mathbb{E}_{Q_{1}}[E(u_{1},X_{2},\dots,X_{T})]>1+\beta^{\star}(u_{1}-\mu)$, since $\beta^{\star}\notin B_{1}$.
-Let $V=\frac{u_{1}-\mu}{u_{1}-u_{0}}\delta_{u_{0}}\otimes Q_{0}+\frac{\mu-u_{0}}{u_{1}-u_{0}}\delta_{u_{1}}\otimes Q_{1}$.
-Then, $V\in\mathcal{H}^{T}_{\mu}$.
-Moreover,
-$\mathbb{E}_{V}[E]>\tfrac{u_{1}-\mu}{u_{1}-u_{0}}\big{(}1+\beta^{\star}(u_{0}-\mu)\big{)}+\tfrac{\mu-u_{0}}{u_{1}-u_{0}}\big{(}1+\beta^{\star}(u_{1}-\mu)\big{)}=1\,,$
-which is a contradiction since $E\in\mathcal{E}^{T}_{\mu}$.
-We have thus established that $B_{0}\cap B_{1}$ is non-empty, and in particular there is $\lambda_{1}\in I_{\mu}$ such that, for every $x\in\mathcal{X}$ and every $Q\in\mathcal{H}_{\mu}^{T-1}$,
-$\mathbb{E}_{Q}[E(x,X_{2},\dots,X_{T})]\leq 1+\lambda_{1}(x-\mu)\,.$
-Fix $y\in\mathcal{X}$ such that $1+\lambda_{1}(y-\mu)\neq 0$.
-Define $E_{y}:\mathcal{X}^{T-1}\to\mathbb{R}$ via $E_{y}(x^{2:T})=E(y,x^{2:T})/(1+\lambda_{1}(y-\mu))$, where we recall that $x^{2:T}$ is a shorthand for $(x_{2},\dots,x_{T})$.
-$E_{y}$ is non-negative, since by definition of $I_{\mu}$ we know that $1+\lambda(y-\mu)$ is non-negative for any $\lambda\in I_{\mu}$.
-From what we have already shown, $E_{y}$ is an e-variable for $\mathcal{H}^{T-1}_{\mu}$.
-Thanks to the inductive hypothesis, for each $t\in[2:T]$ there exist a function $\tilde{\lambda}_{t}^{y}:\mathcal{X}^{t-2}\to I_{\mu}$ such that $E_{y}(x^{2:T})\leq\prod_{t=2}^{T}(1+\tilde{\lambda}_{t}^{y}(x^{2:(t-1)}))(x_{t}-\mu))$, for any $x^{2:T}\in\mathcal{X}^{T-1}$.
-For all $t\in[2:T]$, and every $x^{t-1}\in\mathcal{X}^{t-1}$, define
-$\lambda_{t}(x^{t-1})=\begin{cases}0&\text{if $1+\lambda_{1}(x_{1}-\mu)=0$};\\ \big{(}1+\lambda_{1}(x_{1}-\mu)\big{)}\tilde{\lambda}_{t}^{x_{1}}(x^{2:(t-1)})&\text{otherwise.}\end{cases}$
-So far, we have proved that for any $x^{T}\in\mathcal{X}^{T}$, such that $1+\lambda_{1}(x_{1}-\mu)\neq 0$, we have that $E(x^{T})\leq E_{\lambda^{T}}(x^{T})$.
-We only need to check that this is still true when $01+\lambda_{1}(x_{1}-\mu)=$ 0.
-In such case, $E$ is $Q$-almost surely null for all $Q\in\mathcal{H}^{T-1}_{\mu}$, which implies that $E(x^{T})=0$, and so we conclude.
-∎
-Proof of Theorem 3
-Proof.
-We start by recalling and introducing some definitions.
-For $T\geq 1$, $\Lambda^{T}_{\mu}$ is the set whose elements are $T$-tuple $\lambda^{T}=(\lambda_{1},\dots,\lambda_{T})$, with $\lambda^{t}:\mathcal{X}^{t-1}\to I_{\mu}$.
-We endow $\Lambda^{T}_{\mu}$ with the product topology of point-wise convergence.
-Specifically, a sequence $(\lambda^{t})_{t\geq 0}\to\tilde{\lambda}^{t}$ if, and only if, $\lambda_{t}\to\tilde{\lambda}_{t}$ point-wise for all $t\leq T$.
-For convenience, we will henceforth refer to this topology as the $\Lambda$-topology.
-It is straightforward to check that under the $\Lambda$-topology $\Lambda_{\mu}^{T}$ is compact, since $I_{\mu}$ is bounded and $\mathcal{X}$ has finite cardinality.
-Our proof strategy will rely on an inductive reasoning.
-For this, it is convenient to introduce the notion of a $T$-e-process, for $T\geq 1$.
-This is a sequence $(E_{t})_{t\geq 0}$ of non-negative functions $E_{t}:\mathcal{X}^{t}\to[0,+\infty)$, such that $\mathbb{E}_{P}[E_{\tau}]\leq 1$ for any $\tau\in\mathcal{T}_{T}$, and any $P\in\mathcal{H}_{\mu}^{\infty}$.
-We remark that an e-process is a $T$-e-process for all $T\geq 1$.
-We are now ready to start the proof.
-First, we will show that $\mathcal{E}^{\mathrm{cb},\infty}_{\mu}$ is a majorising e-process class.
-For any $T\geq 1$ and any $T$-e-process $E$, let $\tilde{\Lambda}^{T}_{E,\mu}\subseteq\Lambda^{T}_{\mu}$ denote the set of the $\lambda^{T}$ such that $E_{t}\preceq E_{\lambda^{t}}$, for all $t\in[1:T]$.
-We note that this is a closed and compact set under the $\Lambda$-topology that we introduced earlier.
-We now show that for every $T\geq 1$ and every $T$-e-process $E$, the set $\tilde{\Lambda}^{T}_{E,\mu}$ is non-empty.
-We prove this claim by induction.
-Since $E_{1}$ must be an e-variable for $\mathcal{H}_{\mu}^{1}$, as a consequence of Theorem 1, $\tilde{\Lambda}^{1}_{E,\mu}$ is non-empty.
-Now, fix $T\geq 2$ and a $T$-e-process $E$, and assume that, for any $(T-1)$-e-process $E^{\prime}$, $\tilde{\Lambda}^{T-1}_{E^{\prime},\mu}$ is non-empty.
-We will show that this implies that $\tilde{\Lambda}_{E,\mu}^{T}$ is non-empty.
-Fix an arbitrary $x\in\mathcal{X}$ and define the process $E^{x}=(E^{x}_{t})_{t\geq 1}$ such that $E_{t}^{x}(x^{t})=E_{t+1}(x,x_{1},\dots,x_{t})$, for any $x^{t}\in\mathcal{X}^{t}$.
-Define the sets
-$B_{0}=\big{\{}\beta\in I_{\mu}\,:\,\forall x\in\mathcal{X}\cap[0,\mu),\forall Q\in\mathcal{H}^{\infty}_{\mu},\forall\tau\in\mathcal{T}_{T-1},\,\mathbb{E}_{Q}[E^{x}_{\tau}]\leq 1+\beta(x-\mu)\big{\}}\,;$
-$B_{1}=\big{\{}\beta\in I_{\mu}\,:\,\forall x\in\mathcal{X}\cap(\mu,1],\forall Q\in\mathcal{H}^{\infty}_{\mu},\forall\tau\in\mathcal{T}_{T-1},\,\mathbb{E}_{Q}[E^{x}_{\tau}]\leq 1+\beta(x-\mu)\big{\}}\,.$
-We claim that $B_{0}$ and $B_{1}$ are non-empty.
-Indeed, we now show that $\mu^{-1}\in B_{1}$.
-Analogously one can prove that $(\mu-1)^{-1}\in B_{0}$.
-The argument here follows closely the one depicted in Figure 1.
-Fix $x\in\mathcal{X}\cap[0,\mu)$, $Q\in\mathcal{H}_{\mu}^{\infty}$, and $\tau\in\mathcal{T}_{T-1}$.
-Define $\tau_{x}:\mathcal{X}^{\infty}\to\mathbb{N}$ as $\tau_{x}(x^{\infty})=1+\tau(x^{2:\infty})$, if $x_{1}=x$, otherwise $\tau_{x}(x^{\infty})=1$.
-It is easily checked that $\tau\in\mathcal{T}_{T}$.
-Also, we let $P_{x}=\frac{\mu}{x}\delta_{x}+(1-\frac{\mu}{x}\delta_{0})$, and we define $V_{x}=P_{x}\otimes Q$ [^12] Note that both $V_{x}$ and $Q$ are measures on $\mathcal{P}_{\mathcal{X}^{\infty}}$.
-Here we mean that $V_{x}(\{x^{\infty}\})=P_{x}(\{x_{1}\})Q(\{x^{2:\infty}\})$ for all $x^{\infty}\in\mathcal{X}^{\infty}$.
-Clearly, $V_{x}\in\mathcal{H}_{\mu}^{\infty}$.
-Since $E$ is a $T$-e-process, we have
-$1\geq\mathbb{E}_{V_{x}}[E_{\tau_{x}}]=(1-\tfrac{\mu}{x})E_{1}(0)+\tfrac{\mu}{x}\mathbb{E}_{Q}[E^{x}_{\tau}]\geq\tfrac{\mu}{x}\mathbb{E}_{Q}[E^{x}_{\tau}]\,.$
-So, $E_{Q}[E^{x}_{\tau}]\leq\frac{x}{\mu}=1+\mu^{-1}(x-\mu)$, and so $\mu^{-1}\in B_{1}$.
-Since both $B_{0}$ and $B_{1}$ can be written as intersections of non-empty closed intervals, we can find $\beta_{0}$ and $\beta_{1}$ such that $B_{0}=[(\mu-1)^{-1},\beta_{0}]$ and $B_{1}=[\beta_{1},\mu^{-1}]$.
-As usual, we want to show that $B_{0}\cap B_{1}$ is non-empty, or equivalently that $\beta_{1}\leq\beta_{0}$.
-Assume that this was not the case and there is $\beta^{\star}\in(\beta_{0},\beta_{1})$.
-Since $\beta^{\star}\notin B_{0}$, there must be $u_{0}<\mu$ in $\mathcal{X}$, $Q_{0}\in\mathcal{H}_{\mu}^{\infty}$, and $\tau_{0}\in\mathcal{T}_{T-1}$ such that $\mathbb{E}_{Q_{0}}[E^{u_{0}}_{\tau_{0}}]>1+\beta^{\star}(u_{0}-\mu)$.
-Similarly, as $\beta^{\star}\notin B_{1}$, one can find $u_{1}>\mu$ in $\mathcal{X}$, $Q_{1}\in\mathcal{H}_{\mu}^{\infty}$, and $\tau_{1}\in\mathcal{T}_{T-1}$, such that $\mathbb{E}_{Q_{1}}[E^{u_{1}}_{\tau_{1}}]>1+\beta^{\star}(u_{1}-\mu)$.
-Define $\tau:\mathcal{X}^{\infty}\to\mathcal{N}$ via $\tau(x^{\infty})=1+\tau_{0}(x^{2:\infty})$, if $x_{1}<\mu$, $\tau(x^{\infty})=1+\tau_{1}(x^{2:\infty})$ if $x_{1}>\mu$, and $\tau(x^{\infty})=1$ if $x_{1}=\mu$.
-Also, let $V=\frac{u_{1}-\mu}{u_{1}-u_{0}}\delta_{u_{0}}\otimes Q_{0}+\frac{\mu-u_{0}}{u_{1}-u_{0}}\delta_{u_{1}}\otimes Q_{1}$.
-Then, we have that
-$\mathbb{E}_{V}[E_{\tau}]=\frac{u_{1}-\mu}{u_{1}-u_{0}}\mathbb{E}_{Q_{0}}[E^{u_{0}}_{\tau_{1}}(X_{2},\dots,X_{T})]+\frac{\mu-u_{0}}{u_{1}-u_{0}}\mathbb{E}_{Q_{1}}[E^{u_{1}}_{\tau_{1}}(X_{2},\dots,X_{T})]>1\,.$
-However, this is a contradiction since $\tau\in\mathcal{T}_{T}$, $V\in\mathcal{H}_{\mu}^{\infty}$, and $E$ is a $T$-e-process.
-We can thus conclude that $B_{0}\cap B_{1}\neq\varnothing$ and, in particular, there must be $\lambda_{1}\in I_{\mu}$ such that, for every $x\in\mathcal{X}$, every $Q\in\mathcal{H}_{\mu}^{\infty}$, and every $\tau\in\mathcal{T}_{T-1}$, we have
-$\mathbb{E}_{Q}[E^{x}_{\tau}]\leq 1+\lambda_{1}(x-\mu)\,.$
-Now, for any arbitrary $x\in\mathcal{X}$ such that $1+\lambda_{1}(x-\mu)\neq 0$. Define the process $\tilde{E}^{(x)}=(\tilde{E}^{(x)}_{t})_{t\geq 0}$ via
-$\tilde{E}^{(x)}_{t}(x^{t})=\frac{E_{t+1}(x,x_{1},\dots,x_{t})}{1+\lambda_{1}(x-\mu)}\,.$
-Let us show that $\tilde{E}^{(x)}$ is a $(T-1)$-e-process.
-It is clearly non-negative, as $\lambda_{1}\in I_{\mu}$ implies $1+\lambda_{1}(x-\mu)\geq 0$.
-Fix $Q\in\mathcal{H}_{\mu}^{\infty}$ and $\tau\in\mathcal{T}_{T-1}$.
-From what we have shown, we have
-$\mathbb{E}_{Q}[\tilde{E}^{(x)}_{\tau}]=\frac{\mathbb{E}_{Q}[E^{x}_{\tau}]}{1+\lambda_{1}(x-\mu)}\leq 1\,.$
-In particular, the inductive argument yields the existence of $\tilde{\lambda}_{(x)}^{T-1}\in\Lambda^{T-1}_{\mu}$, such that for all $t\in[1:T-1]$ we have $\tilde{E}^{(x)}\preceq E_{\tilde{\lambda}^{t}_{(x)}}$.
-Let $\lambda^{T}\in\Lambda^{T}_{\mu}$ be defined via $\lambda_{t}(x^{T})=\big{(}1+\lambda_{1}(x_{1}-\mu)\big{)}\tilde{\lambda}_{(x_{1}),t}(x_{2},\dots,x_{T})$, where we set $\tilde{\lambda}_{(x),t}$ identically equal to zero whenever $01+\lambda_{1}(x-\mu)=$ 0.
-We will now show that $\lambda^{T}\in\tilde{\Lambda}_{E,\mu}^{T}$.
-Fix any $x^{T}\in\mathcal{X}^{T}$.
-If $1+\lambda_{1}(x_{1}-\mu)\neq 0$, then for any $t\in[1:T]$ we have that
-$E_{t}(x^{t})=\big{(}1+\lambda_{1}(x_{t}-\mu)\big{)}E^{(x_{1})}_{t-1}(x_{2},\dots,x_{t})\leq\prod_{s=1}^{t}\big{(}1+\lambda_{s}(x^{s-1})(x_{s}-\mu)\big{)}=E_{\lambda^{t}}(x^{t})\,.$
-If $01+\lambda_{1}(x_{1}-\mu)=$ 0, then, for every $t\in[1:T-1]$, $\mathbb{E}_{Q}[E_{t}^{x_{1}}]=0$ for every $Q\in\mathcal{H}^{\infty}$, which implies that $E_{t}^{x_{1}}$ is identically null.
-In particular, $E_{t}(x^{t})=0=E_{\lambda^{t}}(x^{t})$.
-This shows that $\lambda^{T}\in\tilde{\Lambda}_{E,\mu}^{T}$, and so $\tilde{\Lambda}_{E,\mu}^{T}$ is non-empty.
-What we have shown so far allows to say that given an e-process $E$, for every $T\geq 1$ we can find a $\lambda^{T}\in\Lambda^{T}$ such that $E_{t}\leq E_{\lambda^{t}}$ for every $t\in[1:T]$.
-However, this does not tell us yet that there is a $\lambda^{\infty}\in\Lambda^{\infty}_{\mu}$ such that the same property holds.
-To show this, we leverage the compactness of $\tilde{\Lambda}^{T}_{E,\mu}$ under the $\Lambda$-topology.
-For $T^{\prime}\geq T$, let $\tilde{\Lambda}^{T|T^{\prime}}_{E,\mu}$ denote the set of $\lambda^{T}\in\Lambda^{T}$ such that there is $\tilde{\lambda}^{T^{\prime}}\in\tilde{\Lambda}_{E,\mu}^{T^{\prime}}$ satisfying, for all $t\in[1:T]$, $\lambda_{t}=\tilde{\lambda}_{t}$.
-It is easily checked that $\tilde{\Lambda}^{T|T^{\prime}}_{E,\mu}$ is compact in the $\Lambda$-topology.
-Moreover, if $T^{\prime\prime}\geq T^{\prime}$, then $\tilde{\Lambda}^{T|T^{\prime\prime}}_{E,\mu}\subseteq\tilde{\Lambda}^{T|T^{\prime}}_{E,\mu}$.
-Since the intersection of countably many non-empty compact sets is non-empty, we can define $\tilde{\Lambda}^{1|\infty}_{E,\mu}=\bigcap_{T\geq 1}\tilde{\Lambda}^{1|T}_{E,\mu}$, which is a non-empty set.
-Now, fix $\lambda_{1}\in\tilde{\Lambda}^{1|\infty}_{E,\mu}$.
-We have that $E_{1}\preceq E_{\lambda_{1}}$.
-By construction $\lambda_{1}\in\tilde{\Lambda}^{1|2}_{E,\mu}$, and so we can find a $\lambda_{2}$ such that $(\lambda_{1},\lambda_{2})\in\tilde{\Lambda}^{2}_{E,\mu}$.
-We can go on like this indefinitely.
-In particular, we can construct a sequence $\lambda^{\infty}$ such that $\lambda^{T}\in\tilde{\Lambda}^{T}_{E,\mu}$ for all $T\geq 1$.
-This concludes the proof that $\mathcal{E}^{\mathrm{cb},\infty}_{\mu}$ is a majorising e-process class.
-We are left with checking that every e-process in $\mathcal{E}^{\mathrm{cb},\infty}_{\mu}$ is maximal.
-Fix $\lambda^{\infty}\in\Lambda^{\infty}_{\mu}$, and let $E$ be an e-process such that $E\succeq E^{\lambda^{\infty}}$.
-Clearly, we must have $E_{0}=1=E_{0}^{\lambda^{\infty}}$.
-Now, fix $T\geq 1$.
-We have that $E_{T}\succeq E_{\lambda^{T}}$.
-Since $E$ and $E^{\lambda^{\infty}}$ are e-processes, choosing the stopping time that is identically equal to $T$ shows that $E_{T}$ and $E_{\lambda^{T}}$ are $T$-round e-variables.
-In particular, by Theorem 2, $E_{\lambda^{T}}$ is maximal, and so $E_{T}\succeq E_{\lambda^{T}}$ implies that $E_{T}$ and $E_{\lambda^{T}}$ coincide.
-Since this is true for all $T\geq 1$, we conclude.
-∎
-Characterising the e-variables for ^$\hat{\mathcal{H}}_{1/2}^{2}$ with $\mathcal{X}=\{0,1/2,1\}$
+Appendix A
+A.1 Proof of Proposition 2
+Proof 7
+For the first claim, without loss of generality we let $\mathcal{X}=[0,1]$, which implis the desired result for any $\mathcal{X}$.
+For any $\alpha\in\mathbb{R}$, let $S_{\alpha}:x\mapsto E^{\mathrm{H}}_{\alpha}(0)+(E^{\mathrm{H}}_{\alpha}(1)-E^{\mathrm{H}}_{\alpha}(0))x$ be the straight line that intersects $E^{\mathrm{H}}_{\alpha}$ at $x=0$ and $x=1$.
+Since $E^{\mathrm{H}}_{\alpha}$ is a convex function on [0,1][0,1][ 0 , 1 ], we have $E^{\mathrm{H}}_{\alpha}(x)\leq S_{\alpha}(x)$ for all $x\in[0,1]$.
+Let us show that $E_{\lambda_{\alpha}}(x)\geq S_{\alpha}(x)$ for all $x$.
+As $E_{\lambda_{\alpha}}$ and $S_{\alpha}$ represents parallel straight lines, it suffices to show inequality for a single point, say $x=\mu$ where $E_{\lambda_{\alpha}}$ equals 111.
+Hence, it is enough to prove that $1+(e^{\alpha}-1)\mu\leq e^{\alpha\mu+\alpha^{2}/8}$.
+As $\mu\in(0,1)$, we always have $1+(e^{\alpha}-1)\mu>0$, and we can define $u:\mathbb{R}\to\mathbb{R}$ as
+$u(\alpha)=\frac{\alpha^{2}}{8}+\mu\alpha-\log\big{(}1+\mu(e^{\alpha}-1)\big{)}\,.$
+Clearly, if $u$ is non-negative, then $1+(e^{\alpha}-1)\mu\leq e^{\alpha\mu+\alpha^{2}/8}$ for all $\alpha\in\mathbb{R}$, and so we obtain the desired claim.
+First, note that $u$ is twice differentiable, and we can explicitly compute its first and second derivatives:
+$u^{\prime}(\alpha)=\frac{\alpha}{4}+\mu-\frac{\mu e^{\alpha}}{1+\mu(e^{\alpha}-1)}\quad\text{ and }\quad u^{\prime\prime}(\alpha)=\frac{1}{4}-\frac{\mu e^{\alpha}}{1+\mu(e^{\alpha}-1)}\left(1-\frac{\mu e^{\alpha}}{1+\mu(e^{\alpha}-1)}\right)\,.$
+For any $\xi\in\mathbb{R}$ we have $\xi(1-\xi)\leq 1/4$, so $u^{\prime\prime}(\alpha)\geq 0$ for all $\alpha$, and $u$ is convex.
+Moreover $u^{\prime}(0)=0$ and $u(0)=0$, so that 0 is the minimum of $u$, which must then be non-negative.
+For the second claim, fix any $\lambda\neq 0$.
+If $\alpha\neq 0$, then $E^{\mathrm{H}}_{\alpha}(\mu)=e^{-\alpha^{2}/8}<1=E_{\lambda}(\mu)$.
+If $\alpha=0$, then $E^{\mathrm{H}}_{\alpha}$ is identically equal to 111, and $\max(E_{\lambda}(0),E_{\lambda}(1))>1$, so we conclude.$\qed$
+A.2 Characterising the e-variables for ^$\hat{\mathcal{H}}_{1/2}^{2}$ with $\mathcal{X}=\{0,1/2,1\}$
 Let $\mathcal{X}=\{0,1/2,1\}$, and $\mu=1/2$.
 Let $Q\in\mathcal{H}_{1/2}^{1}$ and let $q=Q(\{1/2\})$.
-The mean constraints reads $Q(\{1\})+q/2=1/2$, which implies that $Q(\{1\})=(1-q)/2$.
+The mean constraint reads $Q(\{1\})+q/2=1/2$, yielding $Q(\{1\})=(1-q)/2$.
 The normalisation yields $Q(\{0\})=1-q-(1-q)/2=(1-q)/2$.
-This shows that there is a one-to-one correspondence between $[0,1]$ and $\mathcal{H}_{1/2}^{1}$.
-Now, since by definition there is a one-to-one correspondence between $\mathcal{H}_{1/2}^{1}$ and ^$\hat{\mathcal{H}}_{1/2}^{2}$, we obtain that we can parametrise ^$\hat{\mathcal{H}}_{1/2}^{2}$ by $[0,1]$.
-For $q\in[0,1]$, we hence let $P_{q}$ be the (unique) element of $\mathcal{H}_{1/2}^{2}$ that gives mass $q^{2}$ to $(1/2,1/2)$.
+This shows that there is a one-to-one correspondence between [0,1][0,1][ 0 , 1 ] and $\mathcal{H}_{1/2}^{1}$.
+Now, since by definition there is a one-to-one correspondence between $\mathcal{H}_{1/2}^{1}$ and ^$\hat{\mathcal{H}}_{1/2}^{2}$, we obtain that we can parametrise ^$\hat{\mathcal{H}}_{1/2}^{2}$ by [0,1][0,1][ 0 , 1 ].
+For $q\in[0,1]$, we hence let $P_{q}$ be the (unique) element of $\mathcal{H}_{1/2}^{2}$ that gives mass $q^{2}$ to (1/2,1/2)(1/2,1/2)( 1 / 2 , 1 / 2 ).
 Let $E:\mathcal{X}^{2}\to[0,+\infty)$ be a non-negative function. We have that
@@
 In particular, checking for $q=0$ and $q=1$ implies that $\xi_{0}\leq 1$ and $\xi_{2}\leq 1$.
-If both $\xi_{0}$ and $\xi_{2}$ are equal to $1$, then the constraint on $\xi_{1}$ becomes $\max_{q\in[0,1]}(1-\xi_{1})(q^{2}-q)\leq 0$, which implies $\xi_{1}\leq 1$.
-We are left to check if when one among $\xi_{0}$ and $\xi_{2}$ is strictly less then $1$ a $\xi_{1}$ can be larger than one.
+If both $\xi_{0}$ and $\xi_{2}$ are equal to 111, then the constraint on $\xi_{1}$ becomes $\max_{q\in[0,1]}(1-\xi_{1})(q^{2}-q)\leq 0$, which implies $\xi_{1}\leq 1$.
+We are left to check whether $\xi_{1}$ can be larger than 111 when at least one among $\xi_{0}$ and $\xi_{2}$ is strictly smaller than 111.
 In such case, we note that we have $\xi_{2}-\xi_{1}<0$ and $\xi_{0}-\xi_{1}<0$.
@@
 Asking that this quantity is less than one reduces to the constraint $\xi_{1}\leq 1+\sqrt{(1-\xi_{0})(1-\xi_{2})}$.
+A.3 Proof of Theorem 2
+Proof 8
+First let us show that all the e-variables in $\mathcal{E}^{\mathrm{cb},T}_{\mu}$ are maximal.
+This can be proved essentially with the same argument we used for Theorem 1.
+Specifically, fix $E\in\mathcal{E}^{\mathrm{cb},T}_{\mu}$ and let $E^{\prime}\succeq E$ be an e-variable for $\mathcal{H}_{\mu}^{T}$.
+Fix any $x^{T}\in\mathcal{X}^{T}$.
+It is easy to see that there is $P_{x^{T}}\in\mathcal{H}_{\mu}^{T}$ such that $P_{x^{T}}(\{x^{T}\})>0$.
+Then, $0\leq P_{x^{T}}(\{x^{T}\})(E^{\prime}(x^{T})-E(x^{T}))\leq\mathbb{E}_{P_{x^{T}}}[E^{\prime}-E]=\mathbb{E}_{P_{x^{T}}}[E^{\prime}]-1\leq 0$.
+Hence, $E^{\prime}(x^{T})=E(x^{T})$, so $E=E^{\prime}$, since $x^{T}$ was arbitrary, and $E$ is maximal.
+The fact that $\mathcal{E}^{\mathrm{cb},T}_{\mu}$ is a majorising e-class follows directly from the more general Theorem 3.
+Indeed, let $E\in\mathcal{E}_{\mu}^{T}$.
+Then we can consider an e-process $\tilde{E}=(\tilde{E}_{t})_{t\geq 0}\in\mathcal{E}_{\mu}^{\infty}$, with $\tilde{E}_{T}=E$, and $\tilde{E}_{t}=0$ for $t\neq 0$.
+By Theorem 3, there is $\lambda^{\infty}$ such that $E^{\lambda^{\infty}}\succeq\tilde{E}$.
+In particular $E=\tilde{E}_{T}\preceq E_{\lambda^{T}}\in\mathcal{E}^{\mathrm{cb},T}_{\mu}$, which show that $\mathcal{E}^{\mathrm{cb},T}_{\mu}$ is a majorising e-class.
+However, since the general proof of Theorem 3 is rather technical, we provide a detailed argument for Theorem 2 in the case $T=2$ as a simplified example, which might help build intuition for the proof of Theorem 3.
+As we have already shown that all the e-variables in $\mathcal{E}^{\mathrm{cb},2}$ are maximal, we are left with checking that $\mathcal{E}^{\mathrm{cb},2}$ is a majorising e-class.
+So, fix an e-variable $E$ for $\mathcal{H}_{\mu}^{2}$.
+We need to show that there is a $\lambda^{2}\in\Lambda^{2}_{\mu}$ such that $E_{\lambda^{2}}\succeq E$.
+We start by showing that there is a $\lambda_{1}\in I_{\mu}$ such that, for any $Q\in\mathcal{H}^{1}_{\mu}$ and $x\in\mathcal{X}$,
+$\mathbb{E}_{Q}[E(x,X)]\leq 1+\lambda_{1}(x-\mu)\,.$
+For this we will essentially use the idea that was at the core of the proof of Theorem 1.
+For any $Q$ and $x\geq\mu$, we can define $V_{x,Q}=P_{x}\otimes Q$, where $P_{x}=\frac{\mu}{x}\delta_{x}+(1-\frac{\mu}{x})\delta_{0}$.
+Then, $V_{x,Q}\in\mathcal{H}_{\mu}^{2}$.
+We have that $1\geq\mathbb{E}_{V_{x,Q}}[E]=\tfrac{\mu}{x}\mathbb{E}_{Q}[E(x,X)]+\big{(}1-\tfrac{\mu}{x}\big{)}\mathbb{E}_{Q}[E(0,X)]$, and so $\mathbb{E}_{Q}[E(x,X)]\leq x/\mu=F_{\mu}(x)$, with $F_{\mu}$ as in Lemma 3.
+A similar argument can be used to show that $\mathbb{E}_{Q}[E(x,X)]\leq F_{\mu}(x)$ is true also when $x\leq\mu$.
+In particular, the following two sets are non-empty:
+$B_{0}=\big{\{}\beta\in I_{\mu}\,:\,\forall x\in\mathcal{X}\cap[0,\mu)\,,\,\forall Q\in\mathcal{H}_{\mu}^{1}\,,\,\mathbb{E}_{Q}[E(x,X)]\leq 1+\beta(x-\mu)\big{\}}\,;$
+$B_{1}=\big{\{}\beta\in I_{\mu}\,:\,\forall x\in\mathcal{X}\cap(\mu,1]\,,\,\forall Q\in\mathcal{H}_{\mu}^{1}\,,\,\mathbb{E}_{Q}[E(x,X)]\leq 1+\beta(x-\mu)\big{\}}\,.$
+$B_{0}$ and $B_{1}$ are intersections of closed intervals (one per each allowed $x$ and $Q$), and as such they must be closed intervals.
+We can find $\beta_{0}$ and $\beta_{1}$ such that $B_{0}=[(\mu-1)^{-1},\beta_{0}]$ and $B_{1}=[\beta_{1},\mu^{-1}]$.
+We will show that $\beta_{0}\geq\beta_{1}$, and hence that $B_{0}\cap B_{1}\neq\varnothing$.
+The argument is essentially the same one that was depicted in Figure 1.
+Assume that $\beta_{0}<\beta_{1}$, and let $\beta^{\star}\in(\beta_{0},\beta_{1})$.
+As $\beta^{\star}\notin B_{0}$, there is $Q_{0}\in\mathcal{H}^{1}_{\mu}$ and $u_{0}<\mu$ in $\mathcal{X}$ such that $\mathbb{E}_{Q_{0}}[E(u_{0},X)]>1+\beta^{\star}(u_{0}-\mu)$.
+Similarly, there must be $Q_{1}\in\mathcal{H}_{\mu}^{1}$ and $u_{1}>\mu$ in $\mathcal{X}$, such that $\mathbb{E}_{Q_{1}}[E(u_{1},X)]>1+\beta^{\star}(u_{1}-\mu)$, since $\beta^{\star}\notin B_{1}$.
+Let $V=\frac{u_{1}-\mu}{u_{1}-u_{0}}\delta_{u_{0}}\otimes Q_{0}+\frac{\mu-u_{0}}{u_{1}-u_{0}}\delta_{u_{1}}\otimes Q_{1}$.
+Then, $V\in\mathcal{H}^{2}_{\mu}$.
+Moreover,
+$\mathbb{E}_{V}[E]>\tfrac{u_{1}-\mu}{u_{1}-u_{0}}\big{(}1+\beta^{\star}(u_{0}-\mu)\big{)}+\tfrac{\mu-u_{0}}{u_{1}-u_{0}}\big{(}1+\beta^{\star}(u_{1}-\mu)\big{)}=1\,,$
+which is a contradiction since $E\in\mathcal{E}^{2}_{\mu}$.
+We have thus established that $B_{0}\cap B_{1}$ is non-empty, and in particular there is $\lambda_{1}\in I_{\mu}$ such that, for every $x\in\mathcal{X}$ and every $Q\in\mathcal{H}_{\mu}^{1}$, $\mathbb{E}_{Q}[E(x,X)]\leq 1+\lambda_{1}(x-\mu)$.
+Now, fix any $x\in\mathcal{X}$ such that $1+\lambda_{1}(x-\mu)\neq 0$ and define the function $E^{x}:\mathcal{X}\to\mathbb{R}$ as
+$E^{x}:y\mapsto\frac{E(x,y)}{1+\lambda_{1}(x-\mu)}\,.$
+Clearly $E^{x}$ is non-negative, and what we have shown above implies that $\mathbb{E}_{Q}[E^{x}]\leq 1$ for every $Q\in\mathcal{H}^{1}_{\mu}$.
+So, $E^{x}\in\mathcal{E}_{\mu}^{1}$, and there is $\hat{\lambda}_{2}^{x}\in I_{\mu}$ such that
+$E(x,y)\leq\big{(}1+\lambda_{1}(x-\mu)\big{)}\big{(}1+\hat{\lambda}_{2}^{x}(y-\mu)\big{)}$
+for every $y\in\mathcal{X}$.
+On the other hand, if $x\in\mathcal{X}$ is such that $1+\lambda_{1}(x-\mu)=0$, we have that, for every $Q\in\mathcal{H}_{\mu}^{1}$, $\mathbb{E}_{Q}[E(x,X)]=0$.
+Clearly, this implies that $E(x,y)=0$ for every $y\in\mathcal{X}$.
+Hence, we have that, for every $(x,y)\in\mathcal{X}^{2}$,
+$E(x,y)\leq\big{(}1+\lambda_{1}(x-\mu)\big{)}\big{(}1+\hat{\lambda}_{2}(x)(y-\mu)\big{)}\,,$
+where we defined $\hat{\lambda}_{2}:\mathcal{X}\to I_{\mu}$ as $\hat{\lambda}_{2}(x)=\hat{\lambda}_{2}^{x}$ if $1+\lambda_{1}(x-\mu)\neq 0$, and $\hat{\lambda}_{2}(x)=0$ otherwise.
+[…]
+$u(x)$ $=\sup_{y\in\mathcal{X}\cap(\mu,1]}\frac{1}{y-\mu}\left(\frac{E(x,y)}{1+\lambda_{1}(x-\mu)}-1\right)\,;$
+$l(x)$ $=\inf_{y\in\mathcal{X}\cap[0,\mu)}\frac{1}{\mu-y}\left(1-\frac{E(x,y)}{1+\lambda_{1}(x-\mu)}\right)\,.$
+It is straightforward to check that, for any $x\in\mathcal{S}$,
+$(\mu-1)^{-1}\leq u(x)\leq\hat{\lambda}_{2}(x)\leq l(x)\leq\mu^{-1}\,.$
+Now, $u$ is the supremum on $y$ of the Borel mapping $(x,y)\mapsto\frac{1}{y-\mu}\big{(}\frac{E(x,y)}{1+\lambda_{1}(x-\mu)}-1\big{)}$.
+Hence, it is upper semi-analytic by Lemma 6.
+Similarly, $l$ is lower semi-analytic as an infimum.
+In particular, by Proposition 4, there is a Borel mapping $\lambda_{2}:\mathcal{S}\to I_{\mu}$ such that
+$u(x)\leq\lambda_{2}(x)\leq l(x)$
+for all $x\in\mathcal{S}$.
+We can extend $\lambda_{2}$ to the whole $\mathcal{X}$ by letting $\lambda_{2}(x)=0$ if $x\in\mathcal{X}\setminus\mathcal{S}$.
+This extension is still a Borel function since $\mathcal{S}$ is Borel.
+It is now straightforward to check that, for any $(x,y)\in\mathcal{X}^{2}$ we have
+$E(x,y)\leq\big{(}1+\lambda_{1}(x-\mu)\big{)}\big{(}1+\lambda_{2}(x)(y-\mu)\big{)}=E_{\lambda^{2}}(x,y)\,,$
+which concludes the proof for the case $T=2$$\qed$
+A.4 Proof of Theorem 3
+The proof of Theorem 3 involves a few technicalities, mainly in order to solve issues linked to Borel measurability.
+To handle this we will consider “simplified” versions of $\mathcal{H}_{\mu}$, $\mathcal{H}_{\mu}^{T}$, and $\mathcal{H}_{\mu}^{\infty}$.
+A.4.1 Coarsening of the hypothesis
+First, we define a coarser version of $\mathcal{H}_{\mu}$.
+More precisely, we let ¯$\bar{\mathcal{H}}_{\mu}$ denote the set of probability measures in $\mathcal{H}_{\mu}$ whose support has at most two elements.
+We define as $\mathcal{D}$ the set
+$\mathcal{D}=\{(a,a^{\prime})\in\mathcal{X}^{2}\,:\,\mu\in[a,a^{\prime}]\text{ or }\mu\in[a^{\prime},a]\}\,.$
+Define the non-negative function $W:\mathcal{D}\to[0,1]$ as
+$W(a_{1},a_{2})=\frac{a_{2}-\mu}{a_{2}-a_{1}}\,,$ (3)
+here adopting the convention that 0/0=10/0=10 / 0 = 1.
+For $d=(a_{1},a_{2})\in\mathcal{D}$, there is exactly one measure $Q_{d}\in\bar{\mathcal{H}}_{\mu}$ that has support in $\{a_{1},a_{2}\}$.
+This is
+$Q_{d}=W(a_{1},a_{2})\delta_{a_{1}}+(1-W(a_{1},a_{2}))\delta_{a_{2}}\,.$
+Clearly, the mapping $d\mapsto Q_{d}$ is surjective on ¯$\bar{\mathcal{H}}_{\mu}$.
+Moreover, for any Borel $f:\mathcal{X}\to\mathbb{R}$, it is straighforward to check that the mapping $d\mapsto\mathbb{E}_{Q_{d}}[f]$ is Borel from $\mathcal{D}$ to $\mathbb{R}$.
+We can extend these ideas to sequential hypotheses.
+We define ¯$\bar{\mathcal{H}}_{\mu}^{T}$ as a subset of $\mathcal{H}_{\mu}^{T}$ consisting of measures that are built by iterating over $T$ time steps the 2-point construction that we used for ¯$\bar{\mathcal{H}}_{\mu}$.
+More precisely, $Q\in\mathcal{H}_{\mu}$ is in ¯$\bar{\mathcal{H}}_{\mu}^{T}$ if for $X^{T}\sim Q$ the marginal $X_{1}$ is in ¯$\bar{\mathcal{H}}_{\mu}$, and, for every $t\in[2:T]$, the conditional distribution of $X_{t}$ under $Q$, given $X^{t-1}$, belongs to ¯$\bar{\mathcal{H}}_{\mu}$ for $Q$-almost every $X^{t-1}$.
+In particular, the support of $Q$ lies on at most $2^{T}$ trajectories in $\mathcal{X}^{T}$, and each branching step preserves the conditional mean constraint $\mathbb{E}_{Q}[X_{t}|X^{t-1}]=\mu$.
+This defines a subset of $\mathcal{H}_{\mu}^{T}$ where every conditional law is supported on at most two points.
+We can of course extend the whole construction to sequences, and again consider the set ¯$\bar{\mathcal{H}}_{\mu}^{\infty}\subseteq\mathcal{H}_{\mu}^{\infty}$ of the sequences such that $X_{t}$, conditioned on $X^{t-1}$, has mean $\mu$ and support on at most two elements.
+To make things more explicit, consider the case $T=3$.
+Fix $(a_{1},a_{2})\in\mathcal{D}$, $b^{4}=(b_{1},b_{2},b_{3},b_{4})\in\mathcal{D}^{2}$ (namely, $(b_{1},b_{2})\in\mathcal{D}$ and $(b_{3},b_{4})\in\mathcal{D}$), and $c^{8}\in\mathcal{D}^{4}$.
+Then, given this tuple $d=(a^{2},b^{4},c^{8})\in\mathcal{D}^{7}$, we can define a measure $Q_{d}\in\bar{\mathcal{H}}_{\mu}^{3}$ supported on the eight leaves of the tree shown in Figure 2.
+Each level of the tree corresponds to a time step.
+At the root, we begin by choosing among $a_{1}$ and $a_{2}$, by determining the value of $X_{1}$ (whose support under $Q_{d}$ is in $\{a_{1},a_{2}\}$).
+Then, depending on whether we got $a_{1}$ or $a_{2}$, we branch using $(b_{1},b_{2})$ or $(b_{3},b_{4})$, which gives $X_{2}$.
+Finally, again depending on the previous choices, $Q_{d}$ gives different options for $X_{3}$, encoded in the branches leading to the leaves.
+The weights that $Q_{d}$ gives to each leaf is univocally determined by the constraint that the conditional means have to be equal to $\mu$.
+For instance, the mass that $Q_{d}$ assigns to the point $(a_{1},b_{2},c_{3})$ is given by
+$Q_{d}(\{a_{1},b_{2},c_{3}\})=W(a_{1},a_{2})\times(1-W(b_{1},b_{2}))\times W(c_{3},c_{4})\,.$
+Proceeding in this way, we can build a surjection from $\mathcal{D}^{7}$ to ¯$\bar{\mathcal{H}}_{\mu}^{3}$.
+[…]
+Figure 2: Tree representation associated to $Q_{d}$ for $T=3$, where $d=(a^{2},b^{4},c^{8})$.
+More generally, we can proceed analogously and see that every $d\in\mathcal{D}^{2^{T}-1}$ defines a unique $Q_{d}\in\bar{\mathcal{H}}_{\mu}^{T}$.
+For convenience, we henceforth denote $\mathcal{D}^{2^{T}-1}$ as $\mathcal{D}_{T}$.
+We have hence constructed a surjection from $\mathcal{D}_{T}$ to ¯$\bar{\mathcal{H}}^{T}_{\mu}$, mapping $d$ to $Q_{d}$.
+It is clear from its definition that $\mathcal{D}_{T}$ is a Borel subset of $\mathcal{X}^{2(2^{T}-1)}$.
+Moreover, it is easy to check that, for any fixed Borel function $f:\mathcal{X}^{T}\to\mathbb{R}$, $d\mapsto\mathbb{E}_{Q_{d}}[f]$ is a Borel measurable map from $\mathcal{D}_{T}$ to $\mathbb{R}$, which directly follows from the Borel measurability of $W$.
+We remark that, by definition, ¯$\bar{\mathcal{H}}_{\mu}\subseteq\mathcal{H}_{\mu}$.
+In particular, all the e-variables for $\mathcal{H}_{\mu}$ are also e-variables for ¯$\bar{\mathcal{H}}_{\mu}$.
+An equivalent conclusion holds for the e-variables for $\mathcal{H}_{\mu}^{T}$ and the e-processes for $\mathcal{H}_{\mu}^{\infty}$.
+Thus, if we show that $\mathcal{E}^{\mathrm{cb},\infty}_{\mu}$ is a majorising e-process class for ¯$\bar{\mathcal{H}}_{\mu}^{\infty}$, this will automatically imply that it is a majorising e-process class for $\mathcal{H}_{\mu}^{\infty}$.
+This is indeed the strategy that we will follow in the proof of Theorem 3.
+The technical reason that required us to introduce this coarsening of the hypothesis, is that to deal with Borel measurability we will follow an approach similar to what done in the proof of Theorem 2 for $T=2$.
+In particular, we will apply Proposition 4, which tells us that we can always find a Borel function sandwiched between an upper and a lower semi-analytic ones (see A.5).
+To follow this route, we will need the following technical result, whose proof is deferred to A.4.3.
+Lemma 4
+Let $f=(f_{s})_{s\geq 1}$ denote a sequence of bounded Borel functions, with $f_{s}:\mathcal{X}^{s}\to\mathbb{R}$.
+Fix $t\geq 1$ and, for any $x^{t}\in\mathcal{X}^{t}$ and $s\geq 0$, let $f_{s}^{x^{t}}:\mathcal{X}^{s}\to\mathbb{R}$ be defined via $f_{s}^{x^{t}}(y^{s})=f_{t+s}(x^{t},y^{s})$.
+Define the functions $u$ and $l$, from $\mathcal{X}^{t}\to\mathbb{R}$, as
+$u(x^{t})=\sup_{Q\in\bar{\mathcal{H}}^{\infty}_{\mu}}\sup_{\tau\in\mathcal{T}}\mathbb{E}_{Q}[f_{\tau}^{x^{t}}]\quad\text{ and }\quad l(x^{t})=\inf_{Q\in\bar{\mathcal{H}}^{\infty}_{\mu}}\inf_{\tau\in\mathcal{T}}\mathbb{E}_{Q}[f_{\tau}^{x^{t}}]\,.$
+Then, $u$ is upper semi-analytic and $l$ is lower semi-analytic.
+A.4.2 Proof of the theorem
+We start by a preliminary lemma, whose proof follows closely the argument we used to prove Theorem 1.
+Lemma 5
+Let $\mu\in(0,1)$ and $E$ be an e-process for ¯$\bar{\mathcal{H}}^{\infty}_{\mu}$.
+Then there is $\lambda_{1}\in I_{\mu}$ such that, for any $\tau\in\mathcal{T}$, any $Q\in\bar{\mathcal{H}}_{\mu}^{\infty}$, and any $x\in\mathcal{X}$, we have that
+$\mathbb{E}_{Q}[E^{x}_{\tau}]\leq E_{\lambda_{1}}(x)=1+\lambda_{1}(x-\mu)\,,$
+where $E^{x}=(E^{x}_{t})_{t\geq 0}$ is the sequence of Borel functions $E_{t}^{x}:\mathcal{X}^{t}\to I_{\mu}$, defined via $E_{t}^{x}(y^{t})=E_{t+1}(x,y^{t})$, for $t\geq 1$ and $y^{t}\in\mathcal{X}^{t}$, and $E_{0}^{x}=E_{1}(x)$.
+Proof 9
+Define the sets
+$B_{0}=\big{\{}\beta\in I_{\mu}\,:\,\forall x\in\mathcal{X}\cap[0,\mu),\forall Q\in\bar{\mathcal{H}}^{\infty}_{\mu},\forall\tau\in\mathcal{T},\,\mathbb{E}_{Q}[E^{x}_{\tau}]\leq 1+\beta(x-\mu)\big{\}}\,;$
+$B_{1}=\big{\{}\beta\in I_{\mu}\,:\,\forall x\in\mathcal{X}\cap(\mu,1],\forall Q\in\bar{\mathcal{H}}^{\infty}_{\mu},\forall\tau\in\mathcal{T},\,\mathbb{E}_{Q}[E^{x}_{\tau}]\leq 1+\beta(x-\mu)\big{\}}\,.$
+We claim that $B_{0}$ and $B_{1}$ are non-empty.
+Indeed, we now show that $\mu^{-1}\in B_{1}$.
+Analogously one can prove that $(\mu-1)^{-1}\in B_{0}$.
+Fix $x\in\mathcal{X}\cap[\mu,1]$, $Q\in\bar{\mathcal{H}}_{\mu}^{\infty}$, and $\tau\in\mathcal{T}$.
+Define $\tau_{x}:\mathcal{X}^{\infty}\to\mathbb{N}$ as $\tau_{x}(x^{\infty})=1+\tau(x^{2:\infty})$, if $x_{1}=x$, otherwise $\tau_{x}(x^{\infty})=1$.
+It is easily checked that $\tau\in\mathcal{T}$.
+Also, we let $P_{x}=\frac{\mu}{x}\delta_{x}+(1-\frac{\mu}{x})\delta_{0}$, and we define $V_{x,Q}=P_{x}\otimes Q$ [^19] Here we used that, clearly, one can write $\mathcal{X}^{\infty}=\mathcal{X}\times\mathcal{X}^{\infty}$.
+Clearly, $V_{x,Q}\in\bar{\mathcal{H}}_{\mu}^{\infty}$.
+Since $E$ is an e-process,
+$1\geq\mathbb{E}_{V_{x,Q}}[E_{\tau_{x}}]=(1-\tfrac{\mu}{x})E_{1}(0)+\tfrac{\mu}{x}\mathbb{E}_{Q}[E^{x}_{\tau}]\geq\tfrac{\mu}{x}\mathbb{E}_{Q}[E^{x}_{\tau}]\,.$
+So, $E_{Q}[E^{x}_{\tau}]\leq\frac{x}{\mu}=1+\mu^{-1}(x-\mu)$. Thus, $\mu^{-1}\in B_{1}$.
+Once more, our next argument follows closely the one depicted in Figure 1.
+Since both $B_{0}$ and $B_{1}$ can be written as intersections of non-empty closed intervals, we can find $\beta_{0}$ and $\beta_{1}$ such that $B_{0}=[(\mu-1)^{-1},\beta_{0}]$ and $B_{1}=[\beta_{1},\mu^{-1}]$.
+As usual, we want to show that $B_{0}\cap B_{1}$ is non-empty, or equivalently that $\beta_{1}\leq\beta_{0}$.
+Assume that this was not the case and there is $\beta^{\star}\in(\beta_{0},\beta_{1})$.
+Since $\beta^{\star}\notin B_{0}$, there must be $u_{0}<\mu$ in $\mathcal{X}$, $Q_{0}\in\bar{\mathcal{H}}_{\mu}^{\infty}$, and $\tau_{0}\in\mathcal{T}$, such that $\mathbb{E}_{Q_{0}}[E^{u_{0}}_{\tau_{0}}]>1+\beta^{\star}(u_{0}-\mu)$.
+Similarly, as $\beta^{\star}\notin B_{1}$, one can find $u_{1}\in\mathcal{X}\cap(\mu,1]$, $Q_{1}\in\bar{\mathcal{H}}_{\mu}^{\infty}$, and $\tau_{1}\in\mathcal{T}$, such that $\mathbb{E}_{Q_{1}}[E^{u_{1}}_{\tau_{1}}]>1+\beta^{\star}(u_{1}-\mu)$.
+Define $\tau:\mathcal{X}^{\infty}\to\mathbb{N}$ via $\tau(x^{\infty})=1+\tau_{0}(x^{2:\infty})$ if $x_{1}<\mu$, $\tau(x^{\infty})=1+\tau_{1}(x^{2:\infty})$ if $x_{1}>\mu$, and $\tau(x^{\infty})=1$ if $x_{1}=\mu$.
+Also, let $V=\frac{u_{1}-\mu}{u_{1}-u_{0}}\delta_{u_{0}}\otimes Q_{0}+\frac{\mu-u_{0}}{u_{1}-u_{0}}\delta_{u_{1}}\otimes Q_{1}$.
+Then, we have that
+$\mathbb{E}_{V}[E_{\tau}]=\frac{u_{1}-\mu}{u_{1}-u_{0}}\,\mathbb{E}_{Q_{0}}[E^{u_{0}}_{\tau_{0}}]+\frac{\mu-u_{0}}{u_{1}-u_{0}}\,\mathbb{E}_{Q_{1}}[E^{u_{1}}_{\tau_{1}}]>1\,.$
+However, this is a contradiction since $\tau\in\mathcal{T}$, $V\in\bar{\mathcal{H}}_{\mu}^{\infty}$, and $E$ is an e-process.
+We can thus conclude that $B_{0}\cap B_{1}\neq\varnothing$ and, in particular, there must be $\lambda_{1}\in I_{\mu}$ such that, for every $x\in\mathcal{X}\setminus\{\mu\}$, every $Q\in\bar{\mathcal{H}}_{\mu}^{\infty}$, and every $\tau\in\mathcal{T}$, we have $\mathbb{E}_{Q}[E^{x}_{\tau}]\leq 1+\lambda_{1}(x-\mu)$.
+To conclude, we only need to check the case $x=\mu$.
+However, the argument that we used at the beginning of this proof to show that $\mu^{-1}\in B_{1}$ holds for $x=\mu$, showing that $\mathbb{E}_{Q}[E_{\tau}^{\mu}]\leq 1$, for any $Q\in\bar{\mathcal{H}}_{\mu}^{\infty}$ and any $\tau\in\mathcal{T}$.
+So, we conclude.$\qed$
+We can now prove Theorem 3. The main idea is a generalisation of the approach that we used to deal with the case $T=2$ in the proof of Theorem 2.
+Proof 10 (of Theorem 3)
+First, let us show that all the e-processes in $\mathcal{E}^{\mathrm{cb},\infty}_{\mu}$ are maximal (for $\mathcal{H}_{\mu}^{\infty}$).
+Assume that this was not the case.
+Then there is a $\lambda^{\infty}\in\Lambda_{\mu}^{\infty}$ and an e-process (for $\mathcal{H}_{\mu}^{\infty}$) $E\succeq E^{\lambda^{\infty}}$ such that, for some $t\geq t$, $E_{t}\succ E_{\lambda^{t}}$ (clearly this cannot happen for $t=0$, as we must have $E_{0}\leq 1$ for every e-process).
+So, there is $\hat{x}^{t}\in\mathcal{X}^{t}$ such that $E_{t}(\hat{x}^{t})>E_{\lambda^{t}}(\hat{x}^{t})$.
+It is not hard to see that there is a $Q\in\mathcal{H}_{\mu}^{\infty}$ that puts non-zero mass on the set $\{x^{\infty}\in\mathcal{X}^{\infty}\,:\,x^{t}=\hat{x}^{t}\}$.
+We can consider the constant stopping time $\tau=t$.
+Then, we have that $0<E_{t}(\hat{x}^{t})-E_{\lambda^{t}}(\hat{x}^{t})\leq\mathbb{E}_{Q}[E_{t}-E_{\lambda^{t}}]=\mathbb{E}_{Q}[E_{\tau}]-1\leq 0$, a contradiction.
+Hence, we are left with showing that $\mathcal{E}^{\mathrm{cb},\infty}_{\mu}$ is a majorising e-process class for $\mathcal{H}^{\infty}_{\mu}$.
+Since every e-process in for $\mathcal{H}^{\infty}_{\mu}$ is also an e-process for ¯$\bar{\mathcal{H}}^{\infty}_{\mu}$, it is sufficient to show that $\mathcal{E}^{\mathrm{cb},\infty}_{\mu}$ is a majorising e-process class for ¯$\bar{\mathcal{H}}^{\infty}_{\mu}$.
+Fix an e-process $E$ (for ¯$\bar{\mathcal{H}}^{\infty}_{\mu}$).
+We introduce the following notation.
+For $t\geq 1$ and $x^{t}\in\mathcal{X}^{t}$, we denote as $E^{x^{t}}=(E^{x^{t}}_{s})_{s\geq 0}$ the sequence of non-negative functions $E^{x^{t}}_{s}:\mathcal{X}^{s-1}\to I_{\mu}$, defined via $E_{s}^{x^{t}}(y^{s})=E_{t+s}(x^{t},y^{s})$, for any $y^{s}\in\mathcal{X}^{s}$ and $s\geq 1$, and $E_{0}^{x^{t}}=E_{t}(x^{t})$.
+In what follows, we say that a $T$-tuple $\lambda^{T}$ of Borel functions $\lambda_{i}:\mathcal{X}^{i-1}\mapsto I_{\mu}$, dominates $E$ at level $T$ if, for all $Q\in\bar{\mathcal{H}}^{\infty}_{\mu}$, for any $t\in[1:T]$, for all $\tau\in\mathcal{T}$ and $x^{t}\in\mathcal{X}^{t}$, we have that
+$\mathbb{E}_{Q}[E^{x^{t}}_{\tau}]\leq E_{\lambda^{t}}(x^{t})\,.$
+We will show that there is a $\lambda^{\infty}\in\Lambda_{\mu}^{\infty}$ such that, for any $T\geq 1$, $\lambda^{T}$ dominates $E$.
+We construct this sequence progressively.
+By Lemma 5, we know that there is $\lambda_{1}\in I_{\mu}$ that dominates $E$ at level 111.
+Now, say that, for some $T\geq 2$, there is a $\lambda^{T-1}\in\Lambda_{\mu}^{T-1}$ that dominates $E$ at level $T-1$.
+Let us show that this imply the existence of a Borel $\lambda_{T}:\mathcal{X}^{T-1}\to I_{\mu}$, such that $\lambda^{T}=(\lambda^{T-1},\lambda_{T})$ dominates $E$ at level $T$.
+Define $\mathcal{S}=\{x^{T-1}\in\mathcal{X}^{T-1}\,:\,E_{\lambda^{T-1}}(x^{T-1})\neq 0\}$.
+For any $x^{T-1}\in\mathcal{S}$, define the sequence of Borel functions $\tilde{E}^{x^{T-1}}=(\tilde{E}^{x^{T-1}}_{t})_{t\geq 0}$ as follows.
+$\tilde{E}_{0}^{x^{T-1}}=1$, and $\tilde{E}_{t}^{x^{T-1}}(y^{t})=E_{T-1+t}(x^{T-1},y^{t})/E_{\lambda^{T-1}}(x^{T-1})$ for $t\geq 1$ and $y^{t}\in\mathcal{X}^{t}$.
+By construction, $\tilde{E}^{x^{T-1}}$ is an e-process.
+In particular, by Lemma 5, there must be $\tilde{\lambda}_{1}^{x^{T-1}}\in I_{\mu}$ (which of course might depend on $x^{T-1}$ in a non-Borel way) dominating $\tilde{E}^{x^{T-1}}$ at level 111.
+So, for any $x^{T-1}\in\mathcal{S}$ and $y\in\mathcal{X}$,
+$\sup_{Q\in\bar{\mathcal{H}}^{\infty}_{\mu}}\sup_{\tau\in\mathcal{T}}\mathbb{E}_{Q}[E_{\tau}^{(x^{T-1},y)}]\leq E_{\lambda^{T-1}}(x^{T-1})\big{(}1+\tilde{\lambda}_{1}^{x^{T-1}}(y-\mu)\big{)}\,.$
+We now define the mappings $u$ and $l$ on $\mathcal{S}$ as
+$u(x^{T-1})$ $=\sup_{\tau\in\mathcal{T}}\sup_{Q\in\bar{\mathcal{H}}^{\infty}_{\mu}}\sup_{y\in\mathcal{X}\cap(\mu,1]}\frac{1}{y-\mu}\left(\frac{\mathbb{E}_{Q}[E^{(x^{T-1},y)}_{\tau}]}{E_{\lambda^{T-1}}(x^{T-1})}-1\right)\,;$
+$l(x^{T-1})$ $=\inf_{\tau\in\mathcal{T}}\inf_{Q\in\bar{\mathcal{H}}^{\infty}_{\mu}}\inf_{y\in\mathcal{X}\cap[0,\mu)}\frac{1}{\mu-y}\left(\frac{\mathbb{E}_{Q}[E^{(x^{T-1},y)}_{\tau}]}{E_{\lambda^{T-1}}(x^{T-1})}-1\right)\,.$
+It follows from Lemma 4 that $(x^{T-1},y)\mapsto\sup_{\tau\in\mathcal{T}}\sup_{Q\in\bar{\mathcal{H}}^{\infty}_{\mu}}\frac{1}{y-\mu}\big{(}\frac{\mathbb{E}_{Q}[E^{(x^{T-1},y)}_{\tau}]}{E_{\lambda^{T-1}}(x^{T-1})}-1\big{)}$ is upper semi-analytic (note that $\mathcal{S}$ is a Borel set, so the domain restriction does not cause any issue).
+In particular, $u$ is also upper semi-analytic on $\mathcal{S}$ (see Lemma 6 in A.5).
+By an analogous argument, $l$ is lower semi-analytic on $\mathcal{S}$.
+Moreover,
+$(\mu-1)^{-1}\leq u(x^{T-1})\leq\tilde{\lambda}_{1}^{x^{T-1}}\leq l(x^{T-1})\leq\mu^{-1}$
+for all $x^{T-1}\in\mathcal{S}$, and so in particular by Proposition 4 (see A.5) there is a Borel function $\lambda_{T}:\mathcal{S}\to I_{\mu}$ such that $l(x^{T-1})\geq\lambda_{T}(x^{T-1})\geq u(x^{T-1})$ for all $x^{T-1}\in\mathcal{S}$.
+We can extend $\lambda_{T}$ to the whole $\mathcal{X}^{T-1}$, by setting $\lambda_{T}(x^{T-1})=0$, if $x^{T-1}\in\mathcal{X}^{T-1}\setminus\mathcal{S}$.
+Noting that $x^{T-1}\in\mathcal{X}^{T-1}\setminus\mathcal{S}$ implies that $E_{T-1+t}(x^{T-1},y^{t})=0$ for any $t\geq 1$ and $y^{t}\in\mathcal{X}^{t}$, [^20] Indeed, for every $t$ the non-negativity of $E_{T-1+t}$ implies that this must be true for $Q$-almost every $y^{t}$, for every $Q\in\bar{\mathcal{H}}^{\infty}_{\mu}$, and so for every $y^{t}\in\mathcal{X}^{t}$.
+one can easily verify that $\lambda^{T}$ dominates $E$ at level $T$.
+So, we can construct iteratively a sequence $\lambda^{\infty}\in\Lambda_{\mu}^{\infty}$ such that, for each $T\geq 1$, $\lambda^{T}$ dominates $E$ at level $T$.
+It follows immediately that $E^{\lambda^{\infty}}\succeq E$, and so $\mathcal{E}^{\mathrm{cb},\infty}_{\mu}$ is a majorising e-process class for ¯$\bar{\mathcal{H}}_{\mu}^{\infty}$.
+$\qed$
+A.4.3 Proof of Lemma 4
+[…]
+(a) The pruned subtree generated by the stopping time $\tau^{\star}$ (see main text) acting on the tree in Figure 2 (for $d=(a^{2},b^{4},c^{8})$)
+[…]
+(b) The mask that can be applied to the tree in Figure 2 to obtain the same pruned tree as generated by $\tau^{\star}$ (Figure 3(a)).
+Figure 3: Pruned tree and mask representing the action on $d\in\mathcal{D}_{T}$ of a stopping time $\tau^{\star}$ bounded by $T$.
+We recall that for any time horizon $T\geq 1$, $\mathcal{D}_{T}$ is the set of admissible tuples of coefficients that generate elements of ¯$\bar{\mathcal{H}}_{\mu}^{T}$ via the construction that we outlined in A.4.1.
+We now generalise the idea behind Figure 2 and associate to $d\in\mathcal{D}_{T}$ a binary tree structure, with one root and $T$ additional levels.
+This will be particularly useful to deal with e-processes and stopping time.
+For convenience, in what follows we denote as $\mathcal{T}_{T}$ the set of all stopping times bounded by $T$.
+We can notice that a stopping time $\tau\in\mathcal{T}_{T}$ defines a subset of the tree.
+More precisely, it defines a pruned tree, namely a fully connected subtree that contains the root.
+Note that the converse is also true.
+Given $d\in\mathcal{D}_{T}$ and a pruned tree, there is a $\tau\in\mathcal{T}_{T}$ that induces this subtree.
+To make this more explicit, consider the case $T=3$ again.
+We might consider a stopping time $\tau^{\star}$ that stops at 111 if $a_{1}$ is observed.
+Conversely, if $a_{2}$ is present, it if $b_{4}$ is observed stops at 222, otherwise at 333.
+In such case, the observable states are the leaves of the tree in Figure 3(a).
+Conversely, such pruned tree is induced by any stopping time $\tau$ that behaves like $\tau^{\star}$ on $d$.
+[…]
+$\begin{split}\mathbb{E}_{Q}[f_{\tau}]&=W(a_{1},a_{2})f_{1}(a_{1})+(1-W(a_{1},a_{2}))(1-W(b_{3},b_{4}))f_{2}(a_{2},b_{4})\\ &+(1-W(a_{1},a_{2}))W(b_{3},b_{4})\big{(}W(c_{5},c_{6})f_{3}(a_{2},b_{4},c_{5})+(1-W(c_{5},c_{6}))f_{3}(a_{2},b_{4},c_{5})\big{)}\,,\end{split}$ (4)
+no matter the specific $Q$ and $\tau$ involved.
+From this observation, we see that we can well define a mapping $h$ that, given $d\in\mathcal{D}_{T}$, $M\in\mathcal{M}_{T}$, and a vector $f=(f_{t})_{t\in[0:T]}$ of Borel functions, returns the value
+$h(d,M,f)=\mathbb{E}_{Q}[f_{\tau}]\,,$
+[…]
+After all these preliminaries, we are finally ready to prove Lemma 4.
+Proof 11 (of Lemma 4)
+We show that the claim holds for $u$, the proof for $l$ being analogous.
+Let $f$ be a sequence of bounded Borel functions as in the statement.
+Fix $t\geq 1$.
+For $x^{t}\in\mathcal{X}^{t}$ recall that, for all $s\geq 1$, we let $f^{x^{t}}_{s}:y^{s}\mapsto f_{t+s}(x^{t},y^{s})$.
+Fix $T\geq 1$.
+Let $u_{T}:\mathcal{X}^{t}\to\mathbb{R}$ be defined as
+$u_{T}(x^{t})=\sup_{\tau\in\mathcal{T}_{T}}\sup_{Q\in\bar{\mathcal{H}}_{\mu}^{T}}\mathbb{E}_{Q}[f^{x^{t}}_{\tau}]\,.$
+Following the discussion above, we can also write
+$u_{T}(x^{t})=\max_{M\in\mathcal{M}_{T}}\sup_{d\in\mathcal{D}_{T}}h(d,M,f^{x^{t}})\,,$
+where we can take the maximum as $\mathcal{M}_{T}$ has finite cardinality.
+Now, note that for each $M\in\mathcal{M}_{T}$, the mapping $(d,x^{t})\mapsto h(d,M,f^{x^{t}})$ is Borel.
+This follows from the fact that $W$, defined in (3), and all the $f_{t+s}$ are Borel.
+[^21] See (4) (of course replacing $f$ with $f^{x^{t}}$) to get a more concrete idea of how this mapping looks like when $T=3$, with $M$ the mask in Figure 3(b) and $d$ expressed as $(a^{2},b^{4},c^{8})$.
+In particular, by Lemma 6, for each $M\in\mathcal{M}_{T}$ the mapping $x^{t}\mapsto\sup_{d\in\mathcal{D}_{T}}h(d,M,f^{x^{t}})$ is upper semi-analytic.
+Since the maximum of finitely many upper semi-analytic functions is upper semi-analytic, we conclude that $u_{T}$ is upper semi-analytic.
+We can then notice that $u=\sup_{T\geq 1}u_{T}$.
+Since this is the countable supremum of upper semi-analytic functions, it is upper semi-analytic.
+$\qed$
+A.5 A functional variant of Lusin’s separation theorem
+First, let us recall a few standard definitions and results from descriptive set theory. We refer to Karoui and Tan (2013) or to the monograph Kechris (1995) for more details.
+Definition 7
+A set $A\subseteq\mathbb{R}^{d}$ is called analytic if there exists a Polish space $\mathcal{Y}$, and a Borel set $B\subseteq\mathbb{R}^{d}\times\mathcal{Y}$.
+such that $A=\pi(B)$, where $\pi$ is the projection $(x^{d},y)\mapsto x^{d}$, from $\mathbb{R}^{d}\times\mathcal{Y}$ to $\mathbb{R}^{d}$.
+A set $C\subseteq\mathbb{R}^{d}$ is called co-analytic if it is the complement of an analytic set.
+Theorem 4 (Lusin’s separation theorem)
+Let $A$ be an analytic set in $\mathbb{R}^{d}$ and $C$ a co-analytic set in $\mathbb{R}^{d}$. If $A\subseteq C$, then there exists a Borel set $B$ such that $A\subseteq B\subseteq C$.
+Definition 8
+Let $\mathcal{Z}\subseteq\mathbb{R}^{d}$ be a Borel set and $f:\mathcal{Z}\to\overline{\mathbb{R}}$.
+$f$ is upper semi-analytic if its superlevel sets are analytic (namely for every $r\in\mathbb{R}$ the sets $\{f\geq r\}$ and $\{f>r\}$ are analytic).
+$f$ is lower semi-analytic if its sublevel sets are analytic (namely for every $r\in\mathbb{R}$ the sets $\{f\leq r\}$ and $\{f<r\}$ are analytic).
+Clearly, any Borel function is both upper and lower semi-analytic.
+Lemma 6
+Let $\mathcal{Z}\subseteq\mathbb{R}^{d}$ and $\mathcal{Z}^{\prime}\subseteq\mathbb{R}^{d^{\prime}}$ be Borel sets.
+Let $f:\mathcal{Z}\times\mathcal{Z}^{\prime}\to\overline{\mathbb{R}}$ be an upper semi-analytic function.
+Then $u:\mathcal{Z}\to\overline{\mathbb{R}}$ defined as $u(z)=\sup_{z^{\prime}\in\mathcal{Z}^{\prime}}f(z,z^{\prime})$ is upper semi-analytic.
+Similarly, let $l:\mathcal{Z}\to\overline{\mathbb{R}}$ be given by $l(z)=\inf_{z^{\prime}\in\mathcal{Z}^{\prime}}f(z,z^{\prime})$.
+Then, $l$ is lower semi-analytic.
+We now prove a consequence (Proposition 4) of Lusin’s separation theorem that, although might be already known, we could not find in the literature.
+In short, we want to prove that if a lower semi-analytic function dominates an upper semi-analytic function, then there is a Borel function that separates them.
+Lemma 7
+Let $\mathcal{Z}\subseteq\mathbb{R}^{d}$ be a Borel set and $u:\mathcal{Z}\to\mathbb{R}$ be an upper semi-analytic function bounded from below.
+Then there exists a non-decreasing sequence of simple (namely taking finitely many values) upper semi-analytic functions $(u_{n})_{n\geq 1}$ such that $u_{n}\to u$ point-wise.
+Similarly, if $l:\mathcal{Z}\to\mathbb{R}$ is a lower semi-analytic function bounded from above, there exists a non-increasing sequence of simple lower semi-analytic functions $(l_{n})_{n\geq 1}$ such that $l_{n}\to l$ point-wise.
+Proof 12
+First, assume that $u$ is bounded.
+Without loss of generality we can assume that $u$ takes values in [0,1][0,1][ 0 , 1 ].
+For each $n$, for $t=0,\dots,2^{n}$, let $A_{t}^{n}=\{u\geq t2^{-n}\}$.
+Each $A_{t}^{n}$ is an analytic set.
+Define $u_{n}$ as follows.
+For any $x\in A_{2^{n}}^{n}$, $u_{n}(x)=1$.
+For any $t=0,\dots 2^{n}-1$, for $x\in A_{t}^{n}\setminus A_{t+1}^{n}$, $u_{n}(x)=t2^{-n}$.
+Then, it is easily checked that $u_{n}$ is upper semi-analytic.
+Moreover, $u_{n}$ takes finitely many values, and by construction it is non-decreasing and converges uniformly to $u$.
+Now, let $u$ be only bounded from below.
+Without loss of generality we can assume that $u$ is non-negative.
+Then, for each integer $t\geq 1$ we can define $v_{t}=\min(u,t)$.
+Each $v_{t}$ is a bounded upper semi-analytic function, and in particular, we have that there is a non-decreasing sequence $(v_{t,n})_{n\geq 1}$ of simple upper semi-analytic functions that converges uniformly to $v_{t}$.
+Now, for each $t\geq 1$, there is $n_{t}\geq 1$ such that $\sup_{x\in\mathcal{Z}}|v_{t,n_{t}}(x)-v_{t}(x)|\leq 1/t$.
+We define $u_{t}=\max_{s\leq t}v_{s,n_{s}}$.
+Each $u_{t}$ is an upper semi-analytic simple function, as the maximum of finitely many upper semi-analytic simple functions.
+By construction, the sequence $(u_{t})_{t\geq 1}$ is non-decreasing and $u_{t}\to u$ point-wise, since $v_{t}\to u$.
+[…]
+Proposition 4
+Let $\mathcal{Z}\subseteq\mathbb{R}^{d}$ be a Borel set.
+Let $u:\mathcal{Z}\to\mathbb{R}$ be an upper semi-analytic function bounded from below and $l:\mathcal{Z}\to\mathbb{R}$ a lower semi-analytic function bounded from above.
+If $u\preceq l$, there exists a Borel function $b:\mathcal{Z}\to\mathbb{R}$ such that $u\preceq b\preceq l$.
+Proof 13
+We start by considering the case of simple functions, namely we assume that there is a finite set $\Phi=\{\phi_{1},\dots,\phi_{N}\}$ where $u$ and $l$ are valued.
+Without loss of generality we can assume that $\Phi$ is ordered increasingly (namely, $\phi_{i+1}>\phi_{i}$).
+For any $i$, we have that the set $U_{i}=\{u\geq\phi_{i}\}$ is analytic, while $L_{i}=\{l\geq\phi_{i}\}$ is co-analytic.
+The condition $u\preceq l$ implies that $U_{i}\subseteq L_{i}$.
+In particular, by Lusin’s separation theorem (Theorem 4), there is a Borel set $B_{i}$ such that $U_{i}\subseteq B_{i}\subseteq L_{i}$.
+Let $D_{N}=B_{N}$.
+For $1\leq i<N$, define $D_{i}=B_{i}\setminus B_{i+1}$.
+Then, all these sets are Borel, and it is easy to check that the function
+$f=\sum_{i=1}^{N}\phi_{i}\mathbf{1}_{D_{i}}$
+is Borel and satisfies $u\preceq f\preceq l$.
+Now that we have proved the desired claim for the case where $u$ and $l$ are simple, let us consider the generic case.
+Since $u$ is upper semi-analytic and bounded from below, by Lemma 7 there is a non-decreasing sequence $(u_{n})_{n\geq 1}$, of simple upper semi-analytic functions, that converges point-wise to $u$.
+Similarly, there is a non-increasing sequence $(l_{n})_{n\geq 1}$, of simple lower semi-analytic functions, that converges to $l$.
+For each $n$ we have $u_{n}\preceq u\preceq l\preceq l_{n}$, so there is a sequence $(f_{n})_{n\geq 1}$ of Borel functions such that $u_{n}\preceq f_{n}\preceq l_{n}$ for all $n$.
+Let $f=\limsup_{n\to\infty}f_{n}$.
+$f$ is Borel and $u_{n}\preceq f\preceq l_{n}$ for all $n$.
+In particular, $u\preceq f\preceq l$, as desired.$\qed$